### v1

# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

Abstract
Despite considerable efforts on making them robust, real-world ML-based systems remain vulnerable to decision based attacks, as definitive proofs of their operational robustness have so far proven intractable. The canonical approach in robustness evaluation calls for adaptive attacks, that is with complete knowledge of the defense and tailored to bypass it. In this study, we introduce a more expansive notion of being adaptive and show how attacks but also defenses can benefit by it and by learning from each other through interaction.
We propose and evaluate a framework for adaptively optimizing black-box attacks and defenses against each other through the competitive game they form. To reliably measure robustness, it is important to evaluate against realistic and worst-case attacks. We thus augment both attacks and the evasive arsenal at their disposal through adaptive control, and observe that the same can be done for defenses, before we evaluate them first apart and then jointly under a multi-agent perspective. We demonstrate that active defenses, which control how the system responds, are a necessary complement to model hardening when facing decision-based attacks; then how these defenses can be circumvented by adaptive attacks, only to finally elicit active and adaptive defenses.
We validate our observations through a wide theoretical and empirical investigation to confirm that AI-enabled adversaries pose a considerable threat to black-box ML-based systems, rekindling the proverbial arms race where defenses have to be AI-enabled too. Succinctly, we address the challenges posed by adaptive adversaries and develop adaptive defenses, thereby laying out effective strategies in ensuring the robustness of ML-based systems deployed in the real-world.
I Introduction
AI models are predominantly trained, validated, and deployed with little regard to their correct functioning under adversarial activity, often leaving safety, ethical, and broader societal impact considerations as an afterthought. Adversarial contexts further aggravate the typical generalization challenges that these models face with threats beyond model evasion (extraction, inversion, poisoning [27]) while the systems they enable often expose interfaces that can be queried and used as adversarial “instructors”, like in constructing adversarial malware against existing ML-based malware detection [4, 20]. Scoping on model evasion, the most reliable mitigation to date is adversarial training [35, 53], an approach not without limitations as these models often remain irreducibly vulnerable at deployment, particularly against black-box, decision-based attacks [9, 15, 55]. Nevertheless, all such attacks exhibit a behavior at-the-interface that can be described as adversarial itself, a generalization that subsumes adversarial examples and opens a path towards novel defenses and mitigations.
Adversarial behavior is a temporal extension of adversarial examples, perhaps not malicious or harmful in isolation, yet part of an attack as it unfolds over time; it is also the canonical description of adversarial examples in domains like dynamic malware analysis and adversarial RL [51, 24]. Aside from making the underlying models more robust, this behavior can be countered as such rather than relying on hardened models exclusively. As models cannot update their decision boundary in an online manner and in response to adversarial activity on their interface, there has to be a complement to model hardening: for instance active defenses such as rejection or misdirection [7, 44, 16].
In this study we address a crucial gap, as in adversarial machine learning (AML) evaluating the robustness of defenses against oblivious, non-adaptive, and thus suboptimal attackers is inherently problematic [50, 19]. We expand the conventional notion of adaptive, from adapted attacks that have an empirical configuration to bypass the defense, to include the capability to self-adapt, where attacks adaptively control their parameters and evasive actions together in response to how the model under attack and its defenses respond [5]. We demonstrate theoretically and empirically how self-adaptive adversaries can modify their own policies through reinforcement learning (RL) to become both optimal and evade active detection. Notably, we find that this can be performed in a gradient-based manner even in fully black-box contexts [III.1], a capability that properly reflects the level of adversarial threat and does not overestimate the empirical robustness; attackers might compute gradients after all.
This capability, however, enables active defenses in turn, as through rigorous threat modeling and by simulating self-adaptive attackers their full potential is uncovered and effective counter-policies can be learned. To express the call for adaptive evaluations in AML differently, a defense can only be as good as its adversary. This mutual interdependence generates the necessity for both attacks and defenses being self-adaptive, as well as the competitive, zero-sum game they jointly form. In summary, our research engages from two related perspectives: a) countering decision-based attacks against ML models, and b) the self-adaptation of such attacks, resulting in the following contributions:
  • We demonstrate that active defenses against decision-based attacks are a necessary complement to model hardening but not sufficient. Active defenses are inevitably bypassed by self-adaptive attackers; and in turn this necessitates self-adaptive defenses too.
  • To facilitate reasoning on self-adaptive attacks and defenses, we introduce a unified framework called “Adversarial Markov Games” (AMG). We demonstrate how attacks can optimize their policy and evade active detection at the same time; as a counter we propose a novel active defense and employ RL agents to adapt and optimize both.
  • In an extensive empirical evaluation on image classification and across various adversarial settings, we validate our theoretical analysis and show that self-adaptation through RL outperforms baseline attacks, evasion techniques, and model hardening defenses like adversarial training.
In the domain of black-box, decision-based AML, our work indicates that robust evaluations cannot consider attacks or defenses in isolation, and that each should have the capability to modify their operation through interaction and in direct response to other agency in the environment.
II Related Work & Problem Statement
In this work, we focus on the category of adversarial attacks known as decision-based, a subset of query-based attacks that operate solely on the hard-label outputs of the model and are regarded as a highly realistic and pervasive threat in AI-based cybersecurity environments. Despite the lack of the closed-form expression of the model under attack, given enough queries the effectiveness of such black-box attacks can match and even surpass that of white-box techniques like C&W [14].
II-A Attacks & Mitigations
While adversarial attacks have been extensively researched in both white and black-box contexts, defenses have predominantly focused on the white-box context [35, 53]. As the black-box setting discloses considerably less information, a seemingly intuitive conclusion is that white-box defenses should suffice for the black-box case too. Yet black-box attacks like  [9, 15] have shown to be highly effective against a wide range of defenses like gradient masking [6], preprocessing [40, 13], and adversarial training [35]. The vast majority of adversarial defenses provide either limited robustness or are eventually evaded by adapted attacks [50]. Characteristically, preprocessing defenses are often bypassed by expending queries for reconnaissance [46].
The partial exception to this rule is adversarial training [35]. Given dataset $D={(x_{i},y_{i})}^{n}_{i=1}$ with classes $C$ where $x_{i}\in\mathbb{R}^{d}$ is a clean example and $y_{i}\in{1,...,C}$ is the associated label, the objective of adversarial training is to solve the following min-max optimization problem:
$\underset{\phi}{\operatorname{min}}\mathbb{E}_{i\sim D}\underset{\|\delta_{i}\|_{L_{p}}\leq\epsilon}{\operatorname{max}}\;\mathcal{L}(h_{\phi}(x_{i}+\delta_{i}),y_{i})$ (1)
where $x_{i}+\delta_{i}$ is an adversarial example of $x_{i}$, $h_{\phi}:\mathbb{R^{d}}\rightarrow\mathbb{R^{C}}$ is a hypothesis function and $\mathcal{L}(h_{\phi}(x_{i}+\delta_{i}),y_{i})$ is the loss function for the adversarial example $x_{i}+\delta_{i}$. The inner maximization loop finds an adversarial example of $x_{i}$ with label $y_{i}$ for a given $L_{p}$-norm (with $L_{p}\in\{0,1,2,\inf\}$), such that $\|\delta_{i}\|_{l}\leq\epsilon$ and $h_{\phi}(x_{i}+\delta_{i})\neq y_{i}$. The outer loop is the standard minimization task typically solved with stochastic gradient descent. While the convergence and robustness properties of adversarial training have been investigated through the computation of the inner maximization step and by interleaving normal and adversarial training [53], the min-max principle is conspicuous: minimize the possible loss for a worst case (max) scenario.
II-B Stateful Defenses
All decision-based attacks share properties that can be useful in devising defenses against them, beyond adversarial training. Such a property is their inherent sequentiality: as they follow a query generation policy towards the optimal adversarial example, the generated samples are correlated. This might not be the case for the queries themselves however, as the adversary might adaptively employ transformations the model is invariant to, like the query blinding strategy in [16]. This work is also the first to employ a stateful defense against query-based attacks. Another one is PRADA [30], a defense built against model extraction but which also works against model evasion. The efficacy of these approaches however rests on the assumption that queries can be consistently linked to uniquely identifiable actors – that also show limited to no collaboration – so that a buffer of queries can be built for each one of them.
This limitation was recently addressed, together with the ensuing scalability issues in the Blacklight defense, by resourcefully employing hashing and quantization [32]. It remains a similarity/fingerprinting defense however, thus vulnerable to circumvention if an adversary can find a query generation policy to preserve the attack functionality while evading detection; this was recently achieved by adapting existing attacks through the rejection signal that Blacklight returns [21]. Ultimately, any stateful defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be representative only if the adversary has exhausted their adaptive offensive capabilities.
From the mere interaction it is enabled from, stateful detection already points towards adversarial examples being something broader that includes the process of generating them, otherwise defined as adversarial behavior. Furthermore, as AML encompasses domains where visual similarity is of less significance [45], the ensuing question is where does the boundary lie exactly between benign and adversarial examples? In AI-based environments where adversarial activity is also present, we posit that adversarial behavior is a more general formulation, appropriate for describing and countering threats at the interface of interaction. It is also more intuitive in domains where behavior – as information unfolding over time – is the canonical description, e.g in malware or bot detection, where AI can imitate benign programs or human behavior respectively.
II-C On Being Adaptive
The correct way to evaluate any proposed defense is against adaptive attacks: that is, attacks with explicit knowledge of the inner mechanisms of a defense [50]. This has been established in computer security as the stipulation that security through obscurity does not work: every defense has a potential workaround. If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, an active defense like stateful detection is the counterpart to decision-based attacks, and as we will demonstrate further in the paper, also the necessary complement to model hardening.
At the same time, the level of threat that attacks pose is often unclear or not thoroughly evaluated. As demonstrated in the work of Croce et al. [19], it is very common that the parameters of an attack are suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness. This underestimation can be further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have. The practical effectiveness of attacks rests on the ability to adapt the policies that govern both their operation and their evasive capabilities in tandem.
In AML, “adaptive” refers conventionally to attacks that have full knowledge of the inner workings of a defense; a term we will henceforth refer to as adapted. In our work, we expand the term to include adaptive control, defined as the ability of a system to self-adapt: automatically reconfigure itself in response to changes in the dynamics of the environment in order to achieve optimal behavior [5].
Typically, what is to be controlled is known in advance and well-defined. The moment however we consider adaptive evaluations, new controls are immediately implied: in a newly proposed defense, for instance rejection-based, these are controls that can bypass it, like rejection sampling. To flesh out the twofold meaning of adaptive, one has to both imagine new knobs [29], and discover their correct configuration that would lead to the optimal result. We conceptualize this expanded definition of adaptive, essential for having accurate evaluations in AML research, in Figure 1.
Figure 1: In AML adaptive means to invent new knobs that can bypass a defense (lower open set); in control theory it means the precise tuning of all known knobs. In this work, we reformulate adaptive to signify both. In HSJA for example, radius, steps, and jumps are parameters of the attack, while rotate and translate are evasive transformations.
II-D Research Gap
Prior work has focused on adapted attacks by incorporating general knowledge of any defenses and then empirically configured to evade it [14, 15, 9]. Defenses also follow the same adapted paradigm, with parameters empirically defined and then fixed for deployment [16, 32]. Our observation here is that neither attacks nor defenses are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and against other adaptive agents in it, something that comes with considerable limitations in security applications where the latter is a given. To bridge this gap, we provide theoretical treatment and empirical study of existing and novel methodologies adapting through direct interaction with their environment, and denote them as self-adaptive.
Our work builds on a long line of prior research that focuses on different aspects of the competition between adversaries and defenses. Carlini and Wagner [14] show that evaluating existing attacks out-of-the-box is insufficient and that adapted white-box attackers can break defensive distillation. Bose et al[8] propose Adversarial Examples Games (AEG), a zero-sum game between a white-box attacker and a local surrogate of the target model family. At the equilibrium the attacker can generate adversarial examples that have a high success rate against models from the same family, constituting a zero-query, non-interactive approach for generating transferable adversarial examples. Pal et al[37] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium. Ren et al[42] explain adversarial robustness through low and high-order interactions between features. Yasodharan et al[56] study adversarial classification by formulating it as non-zero-sum hypothesis testing game.
As the most relevant and representative setting for threats against deployed AI-based systems, we scope on decision-based, interactive attacks and defenses; we contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive (and thus also the defensive) potential is properly assessed. While “self-adaptive” represents the synthesis of adapted and adaptive control, for the remainder of the paper we employ the term “adaptive” interchangeably, and we use “adapted” to indicate what is conventionally known as adaptive evaluations in AML.
III Theoretical Framework
In this work we engage from two perspectives that are interrelated: a) thwarting decision-based attacks, and b) adapting such attacks together with any evasive actions, as these are in a natural trade-off. Evaluating non-adaptive – in the expansive sense we outlined – attacks or defenses renders any results unreliable and incomplete [50]. When offensive or defensive methodologies become adaptive, the environments that they reside in become non-stationary [28], putting further pressure on the IID foundations that ML builds on. We can approach this multi-agent interaction more generally as a competitive and sequential zero-sum game [34, 26, 8]. To understand the implications of attacks and defenses turning adaptive, before the practical implementation we perform a theoretical investigation of their interactions around an ML-based system, and we initiate our analysis from adversaries as the agency generates all the subsequent reasoning. In the following sections, the notation and terms we introduce are highlighted in red.
III-A Attacks
The most compelling threat for deployed ML-based systems are hard-label, decision-based black-box attacks where no access is assumed to the model or its parameters, only the capacity to submit queries and receive hard labels. One of the first decision-based attacks was Boundary Attack [9]. A large number of others followed – each inventive in its own way – that manage to improve the overall performance, typically measured as the minimum perturbation achieved for the minimal amount of queries submitted. Prominent examples of such attacks are HSJA [15], Guessing Smart (BAGS) [12], Sign-Opt [17], Policy-driven (PDA) [55], QEBA [33], and SurFree [36].
White-box methodologies for constructing adversarial examples like C&W [14] cannot function in black-box environments where there is no closed-form description of the inference pipeline. To facilitate optimization, decision-based attacks commonly reverse the adversarial task by initiating from a sample belonging to the target class, as it can be considered an adversarial example with an unacceptably large perturbation. The task then regains the ability to be solved continuously as it allows for the minimization of the perturbation while staying on the adversarial side of the boundary. Decision-based attacks exhibit further commonalities in their functioning so that one can abstract through their particularities: given starting and original samples $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{g}$ and $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{c}$ respectively, the goal is to iteratively find adversarial candidates $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{t}$, until the distance $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\delta=d(x_{t},x_{c})$ is minimized. This process can follow different algorithmic approaches representing different geometrical intuitions, however to describe them generally we can use a candidate generation policy:
$\pi_{\theta}^{\mathcal{A}}=P\>(x_{t}|x_{g},x_{c},p^{\mathcal{A}},s^{\mathcal{A}})$ (2)
that is conditional on $x_{g}$ and $x_{c}$, with $p^{\mathcal{A}}$ the parameters and $s^{\mathcal{A}}$ the state of the attack. Such parameters are the source and orthogonal steps for BAGS and the number of queries and radius for the gradient approximation in HSJA. Assuming the model answers always and truthfully, this process can be construed as an Markov Decision Process (MDP) to be solved by a policy that minimizes $d$ in the least amount of queries.
Given a multinomial image classifier $\mathcal{M}$, the model under attack is based on a discriminant function $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ that for each input $x\in[0,1]^{d}$ generates an output $y:=\{y\in[0,1]^{m}|\sum_{c=1}^{m}y_{c}=1\}$ – a probability distribution over the $m$ classes. By definition, black-box environments provide no access to these probabilities; instead one can only observe the result of the classifier $C$ that returns the class with maximum probability:
$C(x):=\operatorname*{arg\,max}_{c\in[m]}F_{c}(x)=D(F_{c}(x))$ (3)
with $D$ being $\operatorname*{arg\,max}$, the decision function. The goal in targeted attacks is to change the decision $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}c_{g}\in[m]$ for a correctly classified example $x$, to a predefined target class $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}c_{o}\neq c_{g}$. This process can be facilitated through a function $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\psi$ which given a perturbed example $x_{t}$ at step $t$, it returns a binary indicator of success:
$\psi(x_{t})=\begin{cases}1&\text{if}\quad C(x_{t})=c_{o}\\ -1&\text{if}\quad C(x_{t})\neq c_{o}\end{cases}$ (4)
As long as the model responds, $\psi$ can always be evaluated, it thus constitutes the essential mechanism upon which decision-based attacks build. The adversarial goal can then be described as the following constrained optimization problem:
$\operatorname*{min}_{x_{t}}\>d(x_{t},x_{c})\quad\text{s.t.}\quad\psi(x_{t})=1,$ (5)
where the distance metric $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}d$ is an $\ell_{p}$-norm, with $p\in\{0,1,2,\inf\}$. As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this optimization task highlights the imprecise nature of adversarial examples, something that is further aggravated in domains where visual affinity is of little importance. Customarily, successful or unsuccessful adversarial examples are delimited by an threshold $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\epsilon$ on perturbation, where $d(x,x_{t})\leq\epsilon$. While this constraint might not intuitively translate to non-visual domains, we nonetheless contend that minimal perturbation still remains a central property of adversarial examples.
The most relevant attacks in cybersecurity contexts are black-box and decision-based, and they are becoming increasingly effective. In HSJA for example, its optimization is guaranteed to converge to a stationary point of Eq. (5), and given typical $\epsilon$ values on perturbation imperceptibility this translates to near-perfect attack success rates, even against adversarially trained models. The limitations of adversarial training against decision-based attacks can be attributed to the fundamentally out-of-distribution (OOD) nature of adversarial examples, as that makes the saddle point optimization of Eq. (1) intractable to solve exhaustively. Additionally, it is challenging to incorporate decision-based attacks during stochastic gradient descent: as approaches that navigate the decision boundary, the further the latter is from convergence, the less effective the attack is. Scalability is also an issue as typically in adversarial training a few steps (1-50) of a white-box attack – FGSM or PGD – are required, while decision-based attacks can take orders of magnitude more steps (queries) to produce an adversarial example.
Decision-based attacks typically search for the optimal parameters $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\theta$ of generation policy (2), those that given $x^{i}_{c}$, with $i$ denoting the i-th adversarial episode, minimize Eq. (5) in expectation:
$\operatorname*{arg\,min}_{\theta}\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})],\quad\text{s.t.}\quad\psi(x^{i}_{b})=1,$ (6)
where $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x^{i}_{b}$ is the best adversarial example generated by policy $\pi_{\theta}^{\mathcal{A}}$ during episode $i$. Given its dimensionality, it can be intractable to learn such a policy that modifies the input/problem space directly [39]; CIFAR-10, for instance, has more than 3K features to perturb. In AI enabled systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can assume that $F_{0}=F_{1}=...=F_{t},\forall t$. While this is representative of real-world settings, it is also what enables adversaries to discover adversarial examples that were not identified beforehand. The existence of adversaries which follow a candidate generation policy introduces however a behavior which can be observed and utilized by a defensive methodology. Consequently, while model-hardening approaches like adversarial training are necessary, they can be insufficient in defending against decision-based attacks.
Proposition III.1.
Let $F_{c}$ be the discriminant function of the adversarially trained model $\mathcal{M}$. Then given an HSJA attack, in order for $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]\geq\epsilon$, two capabilities are necessary: a) a decision function $D^{\prime}\neq\operatorname*{arg\,max}$, and b) additional context $\tau$ s.t for adversarial query $x_{t}$, $C(x_{t})=D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$.
Intuitively, HSJA operates in 3 stages which repeat: a binary search that puts $x_{t}$ on the decision boundary, a gradient estimation step, and projection step along the estimated gradient. If the model always responds truthfully, the adversary will be able to accurately perform all these steps and converge to the optimal adversarial; without loss of generality, we can extend this intuition to decision-based attacks which navigate the boundary. Secondly, the model should be able to differentiate between two, otherwise identical, queries when one is part of an attack and the other is not, and this is possible through a stateful representation; see Appendix A for the proof.
III-B Defenses
Proposition III.1 shows that alternative classification policies are useful in the presence of decision-based attacks, for instance classification with rejection or intentional misdirection. In related work, rejection has been realized in the form of conformal prediction where model predictions are sets of classes including the empty one, or learning with rejection [7, 18]; while misdirection has emerged as a technique in adversarial RL and cybersecurity domains [24, 44]. While adversarially training the discriminant function $F$ empirically shows the capacity to resist decision-based attacks, the manner in which the model responds has a complementary potential. This gap between the empirical and theoretically possible robustness to decision-based attacks is the locus where a distinct from model hardening, active and adaptive defense can emerge. Active defenses have direct implications on attacks themselves however. Let us assume an agent carrying out an active defense policy:
$\pi_{\phi}^{\mathcal{D}}=P(\alpha|x_{t},s^{\mathcal{D}}),\>a\in\{0,1\}$ (7)
with $x_{t}$ the query, $s^{\mathcal{D}}$ the state for the defense as created by past queries, and $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\alpha$ the binary decision: when the query is deemed adversarial, it is rejected by returning $\alpha=1$. If this policy is stationary the environment dynamics become in turn stationary and so, next to the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved. In two-player, zero-sum games, the moment an agent follows a stationary policy, it becomes exploitable through the reward obtained by an adversary [49]. Active defenses, as a consequence of decision-based attacks, entail adaptive adversaries.
Proposition III.2.
[…]
A proof for BAGS and HSJA is included in Appendix A. An adversary can reason, as a corollary to Proposition III.1, that such defenses have to be in place as it is suboptimal not too. However, there is a second reason to consider adaptive attacks even in the absence of active defenses, as attack policies are often not optimal out-of-the-box. Adapting attack policies can be seen as performing hyperparameter optimization and as an approach has proven very effective in other black-box or expensive-to-evaluate contexts, like Neural Architecture Search and Data Augmentation [57, 38]. The empirical results in Section V further indicate this correspondence between adaptive and self-optimizing, where the adaptive versions of attacks outperform the non-adaptive, more so against active defenses.
Consider now an active defense that is based on a similarity or a conformal metric. In the twofold meaning of adaptive we introduced in Section II, inventing control implies the capability to bypass a similarity based defense; adaptive control implies strategy instead, the active control of the available tools to evade the defense [1]. Then the adversary’s task is to find the optimal adversarial policy that also evades rejection. The straightforward way to achieve this is by adapting policy (2) itself. Notably, and despite the discrete and black-box context, this optimization can be fully gradient-based [48]. We now demonstrate how adaptively controlling decision-based attacks recovers the gradient-based solvability of the black-box optimization task despite neither the active defense nor the model itself being accessible in closed-form.
Theorem III.1 (Adversarial Policy Gradient).
Given model $\mathcal{M}$ with an active and stationary defense $\pi_{\phi}^{\mathcal{D}}$ (7), adaptive candidate generation policy $\pi_{\theta}^{\mathcal{A}}$ (2) that generates episodes $\tau$ of queries $x_{t}$, and a reward $r(\tau)=\sum_{x_{t}\in\tau}\neg\alpha$, the optimal evasive policy $\mathcal{E}$ is obtained via the gradient of the policy’s expected reward $\mathbb{E}_{\pi^{\mathcal{A}}_{\theta}}[r(\tau)]$.
The proof is included in Appendix A. So far we have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) by mere observation of discrete model decisions, adaptive attacks can become optimal in terms of both evasion and efficiency. The last piece of the puzzle is turning active defenses also adaptive.
Proposition III.3.
An active defense $\pi_{\phi}^{\mathcal{D}}$ (7) achieves its optimal, i.e. maximizes the expectation $\mathbb{E}[\sum_{x_{t}\in\tau}P(\alpha|x_{t},s)]$, by adapting its policy against a stationary evasive policy $\mathcal{E}$.
Proof.
As offensive and defensive policies are strictly competitive, we can define the reward $\rho$ of the defensive policy as $\rho(\tau)=\sum_{x_{t}\in\tau}\alpha$, then by making $\pi_{\theta}^{\mathcal{A}}$ stationary and $\pi_{\phi}^{\mathcal{D}}$ adaptive in Theorem III.1, we can reason that the optimal defensive policy is determined via the gradient of the expected reward $\mathbb{E}_{\pi^{\mathcal{D}}_{\phi}}[\rho(\tau)]$. ∎
Figure 2: Schematic model of an AMG environment. Due to the inherent uncertainty of behavior at either side of the interface, it is a partially observable RL environment mirrored for each agent where one’s decisions become the other’s observations.
III-C Adversarial Markov Games
By reasoning on both offensive and defensive capabilities, we highlighted why we cannot consider them independently from each other. As adaptive decision-based attacks and defenses are logical consequences of each other, by composing them we can form a turn-taking competitive game. A precise game-theoretic formulation requires the exact analytical description of the whole environment: the model, the players and their utility functions, as well as the permitted interactions and the transition dynamics, something exceedingly intractable in this context as well as most cybersecurity environments. Model-free methods however can learn optimal offensive and defensive responses directly through interaction with the environment [44, 43], obviating the need to learn a model of it or to find exact solutions the to bi-level optimization task like Eq. (1) that is inherently NP-hard to solve [11].
To that end, Turn-Taking Partially-Observable Markov Games (TT-POMGs) introduced by Greenwald et al. [25] is a generalization of Extensive-Form Games (EFGs), widely used representations for non-cooperative, sequential decision-making games of imperfect and/or incomplete information; an apt formalism for decision-based attacks and defenses. Another nice property of TT-POMGs is that they can be transformed to equivalent belief state MDPs, significantly simplifying their solution.
The competition underlying adversarial example generation has been explored in no-box and white-box settings [8, 22]. We instead focus on decision-based, interactive environments which are assumed to have unknown but stationary dynamics: any agency present is considered part of the environment and therefore fixed in its behavior. By folding the strategies of other agents into the transition probabilities and the initial probability distribution of the game, an optimal policy computed in the resulting MDP will correspond to the best-response strategy in the original TT-POMG. The congruence between TT-POMGs and MDPs is useful not only from a theoretical perspective, but also for its practical implications in the security of ML-based systems: provided that adversarial agents and their capabilities can be identified through rigorous threat modeling, computing the best response strategy in the simulated environment will correspond to the optimal defense.
This environment that encompasses adversarial attacks, adversarial defenses, and benign queries, can be construed as an Adversarial Markov Game (AMG) – a special case of TT-POMG – and is depicted in Figure 2. Formally, we represent AMG as a tuple $\langle i,S,O,A,\tau,r,\gamma\rangle$
  • $i=\{\mathcal{D},\mathcal{A}\}$ are the players, where $\mathcal{D}$ denotes the defender and $\mathcal{A}$ denotes the adversary. In our model, benign queries are modeled as moves by nature.
  • $S$ is the full state space of the game, while $O=\{O^{\mathcal{D}},O^{\mathcal{A}}\}$ are partial observations of the full state for each player.
  • $A=\{A^{\mathcal{D}},A^{\mathcal{A}}\}$ denotes the action set of each player.
  • $\tau(s,a^{i},s^{\prime})$ represents the transition probability to state $s^{\prime}\in S$ after player $i$ chooses action $a^{i}$.
  • $r=\{r^{\mathcal{D}},r^{\mathcal{A}}\}:O^{i}\times A^{i}\rightarrow\mathbb{R}$ is the reward function where $r^{i}(s,a^{i})$ is the reward of player $i$ if in state $s$ action $a^{i}$ is chosen.
  • $\gamma^{i}\in[0,1)$ is the discount factor for player $i$.
The goal of each player $i$ is to determine a policy $\pi^{i}(A^{i}|O^{i})$ that, given the policy of the other(s), maximizes their expected reward. When a player employs a stationary policy, the AMG reduces to a belief-state MDP where the other interacts with a fixed environment. The game is sequential and turn-taking, so each player $i$ chooses an action $a$ from their set of actions $A^{i}$ which subsequently influences the observations of others.
We have shown that an adaptive defense policy $\pi^{\mathcal{D}}_{\phi}$ is necessary to deter decision-based attacks, and as a consequence the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$ has to be adaptive in turn. As without implausible assumptions one cannot assume access to the exact state of the other agent, the states $O^{\mathcal{D}}$, $O^{\mathcal{A}}$ are partial observations of the complete state $\mathcal{S}$ of the full game. When human agents compete by holding beliefs about each other, they engage in recursive reasoning that in theory of mind is encountered as [I believe that [my opponent believes [that I believe…]]]. In the study of opponent modeling, considering other agent policies as stationary and part of the environment is equivalent to 0th level recursive reasoning: the agent models how the opponent behaves based on the observed history, but not how the opponent would behave based on how the agent behaves [2, 54]. However, as AMGs can be solved by single-agent RL algorithms, we consider more involved recursive reasoning out of scope and perform the empirical evaluation without building explicit models of opponent behavior.
IV Environment Specification
The empirical study we conduct in Section V constitutes diverse instantiations of the general theoretical framework introduced in Section III. Naturally, when working forward from the general to the particular concrete design choices have to be made when specifying the latter, and these choices can have considerable influence on the results. In this section we thus outline the components of the environment: adversaries, defenders, possible interactions and various mechanisms that underscore them.
While extensible to more players, our AMG framework describes a two-player competitive game, so we assume that at a given moment only one attack takes place. From the defensive perspective, incoming queries can be either benign (clean) or part of an attack. A frequently made assumption that influences the effectiveness of stateful detection approaches, however, is the ability to attribute queries to UIDs, e.g., an IP address or a user account. Adversaries can create multiple accounts, use VPNs, collude between them, or in fact accounts or IP addresses might not even be necessary to query the model. To address that, we treat queries irrespective to their source, a strictly more challenging setting for stateful defenses. The environments we build operate solely on the content of queries and not on any other metadata, similar to [32]. Like this work, we reframe the task of detecting and banning UIDs of adversaries with two notable differences. Instead of rejecting queries, a decision that in itself can assist the adversary in optimizing, we misdirect by returning another decision, e.g. the second highest class. Secondly, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution so that is not actually trivial to tell them apart from adversarial noise.
Similarity. As decision-based attacks are governed by a generation policy, successive queries exhibit patterns and degrees of similarity that can be quantified by an appropriate $l_{p}$ norm. However such a norm should not be computed on the original inputs as an adapted adversary can ”invent control” (see Fig. 1) by employing evasive transformations the model is invariant to and bypass the similarity detection. To account for such adversarial strategies, we train a Siamese network with contrastive loss in order to learn an latent space $\mathcal{L}(\cdot)$ where similar inputs are mapped close together, unaffected by added noise or transformations on the inputs. To statefully characterize queries we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign/undetected queries.
[…]
Figure 3: Intuition of misdirection in a two-dimensional decision boundary. The adaptive defense controls a single parameter: the radius around $c_{0}$ (the last known adversarial) where queries $x_{t}$ side are intercepted and the model responds with a non-adversarial decision. $x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
Adaptivity. No evaluation in AML is complete without considering adaptive adversaries: that is with the instruments to bypass the defense, but not necessarily with their optimal configuration. As all stateful defenses so far are similarity based, to bypass them our intuition points towards input transformations the model is invariant to. Formally, for a given query $x_{t}$ we want to compute a transformation $x_{t}^{\prime}=T(x_{t})$ so that $||x_{t}^{\prime}-x_{t}||_{2}\gg||x_{t}-x_{t-1}||_{2}$ while $F(T(x_{t}))=F(x_{t})$. Depending on magnitude and composition of transformations $T$, the identity $F(T(x_{t}))=F(x_{t})$ might not always hold. As we also demonstrate in Section V, $T$ interferes with the perturbations the adversarial policy generates itself: the performance and evasiveness of an attack are thus in a natural trade-off.
So what is the correct composition of transformations $T$ to apply? When shall $T$ be applied, and how does it affect the attack fundamentals? The possible transformations can be considered as set of additional controls, and like attack parameters they themselves can be underperforming out-of-the-box [19]. Thus the combined control of attack and evasion parameters is a prerequisite to properly assess the strength of a defense: their empirical configuration is often suboptimal against a particular defense and a fine balance is often required between evasion and preserving attack performance. This trade-off illustrates why the twofold definition of adaptive we trace in this work is necessary in AML evaluations: first to impart controllability to the task through the definition of what can be controlled, and then to find the precise optimal configuration and strategy of the attack.
Agents & Environments. Unlike the vast majority of competitive games, in AMGs the two players have different action and state sets. AMGs are also asymmetric in the playing cadence: while the defender plays every round, it takes one to several rounds for the adversary to play; HSJA for example is controlled on the iteration rather on the query level. The learning task is complicated further given that the experience upon which each agent learns arrives only after the opponent moves. We address these complications by developing custom learning environments where agents can handle asymmetries and the delayed experience collection by training asynchronously. For the environments and agents respectively, we build on the OpenAI Gym and Stable Baselines 3 libraries  [10, 41]. The exact definitions of the RL states and rewards we experimented with are included in Appendix B.
States & Actions. For the definition of the states we used and the rationale behind them, we point the reader to Appendix B. For actions, we control BAGS through 4 parameters: orthogonal step size, source step size, mask bias, and Perlin bias. HSJA is controlled by 3: the gradient estimation radius, the number of estimation queries, and the jump step size. All evaluations start from controlling these fundamental attack parameters only. If the active defense proves impossible to defeat, we introduce additional knobs corresponding to the evasive actions that control the magnitude and probability of transformations on the input, with the goal to evade detection while preserving semantic content and hence the correct classification. The range of transformations we experimented with as well as their magnitude and probability are listed in Table V. Finally, in both BAGS and HSJA the active defense corresponds to an 1-dimensional continuous action that controls the radius of confinement, as depicted in Fig 3.
Rewards. The success of any RL task relies heavily on how it is rewarded. Engineering an effective reward function is often non-trivial and can contain many intricacies, as reward hacking and specification gaming are common phenomena and the learned behavior can vary significantly [3]. For adversaries, the rewards we experimented with are variations on minimizing the distance to the original example, while defenders are rewarded and penalized for intercepting adversarial and benign queries respectively. More specifically, BAGS terms include tn the distance gains towards the original example (also relative to the initial gap) and on their frequency. For HSJA, additional terms include penalties for the amount of gradient estimation queries, and rewards on the precision of gradient estimation and jump steps. Defensive rewards for BAGS include keeping the best adversarial or the average query close to the starting point, minimizing the size of the step, and correctly intercepting queries. For HSJA, additional terms include penalties for not intercepting queries at the binary search or jump steps. The reward functions in closed-form are shown in Appendix B.
V Evaluation
As the empirical counterpart of our theoretical investigation, we perform an extensive evaluation by defining a range of scenarios that reflect all possible and realistic combinations between adversarial attacks and defenses. Concretely, the research questions corresponding to the theoretical insights and we would like to validate are: 1) Are active defenses a necessary complement to model hardening and to what extent? 2) Are attacks more threatening when adaptive, i.e., do they outperform their vanilla versions and can they evade active detection? 3) If yes, can active defenses recoup their performance by also turning adaptive?
Metrics. The first metric we use is the $\ell_{2}$ norm between perturbed and original samples, over ASR (Attack Success Rate), for several reasons: 1) it is significantly more fine grained than a binary indicator of success, thus a better choice for comparing baseline attacks and defenses, their adaptive versions, and to each other; 2) it is not based on an arbitrary perceptual threshold, thus it is more informative of the full range of behavior between perfect attack and perfect defense, 3) a threshold moved around can yield widely varying results, perturbation size is an unaffected global metric, and 4) our RL agents learn optimal policies (also) by using this distance as a signal, where the adversary is rewarded by minimizing it and the defender by maximizing it. The budgets we evaluate over are 1K, 2K and 5K queries. As robustness and classification accuracy are in a natural trade-off, the second metric of interest is the accuracy on benign samples (Clean Acc.) that the original model and the active defense achieve together.
V-A Evaluation Setup
The explicit goal for the agents is – by querying the model and being queried – to learn offensive or defensive policies that are general: they transfer to any other evasion task. Thus after training and validating the agents, the final performance is reported on a fixed hold-out set of 100 adversarial episodes where the starting and original samples are selected at random. As is best practice in AML, candidate samples are only those that are correctly classified by the model. For each scenario we perform a limited hyperparameter and reward function exploration (max 30 trials), with the intention to root out poor combinations rather than exhaust the search space. These are described in more detail in Appendix C.
The black-box attacks we render adaptive and evaluate are BAGS and HSJA, as they represent two fundamentally different approaches, are highly effective, and show the highest evasion potential [32]. BAGS is a search-based methodology and operates in a stochastic manner where every query submitted is a new and potentially better adversarial example. Instead, HSJA is deterministic, composed of 3 different stages, and the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means to approximate the gradient at the decision boundary.
In training and evaluation, the adversarial game is played as follows: the adversary starts by submitting a query, then the defender responds either truthfully (the true model prediction) or by misdirecting (the second highest probability class). Then the environment decides with chance $p$ if the adversary moves next, otherwise a benign query is drawn. In either case, it is the defender’s turn; in testing they are also oblivious to the nature of the query and know only the content. For the majority of experiments we fix $p=0.5$; we also evaluate adaptive defenses when no attack is present ($p=0$) in subsection V-D.
The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two different datasets – MNIST and CIFAR10 – and over two models with the same architecture but with different training regimes: one with adversarial training and one without. The transition from single-agent to multi-agent RL hides challenges however. In every scenario, we approach the AMG as a belief-state MDP (the requirement of knowing the exact opponent policies is relaxed) and use PPO [43] agents to discover optimal policies that will also constitute best responses for the full AMG [54]. However, learning independently of other agency breaks the theoretical guarantees of convergence [52] – this can be observed in the sole scenario where both agents learn simultaneously and we thus report them separately: 7 & 8 AA-AD. The full list of scenarios is:
TABLE I: Mean $l_{2}$ perturbation for 1K, 2K, and 5K queries for CIFAR-10, for normally and adversarially trained models. Clean Acc. records the accuracy of the base model and any defenses present on benign queries; in the first two scenarios (no active defense) the baseline clean accuracy is recorded. Yellow scenarios denote the baseline attack performance, green denote active and/or adaptive defenses, and red denote adaptive attacks. The green-to-red heatmap is a renormalised range of 0 to 100, where 0 is perfect defense (the gap value) and 100 is perfect attack (lowest value for 5K queries). The asterisk denotes scenarios where input transformations were used for evasion.
Adv. Trained Scenario CIFAR-10 Gap: 20.01
BAGS HSJA Clean Acc.
1K 2K 5K 1K 2K 5K BAGS HSJA
0: VA-ND 8.15 7.64 7.22 3.40 1.45 0.42 91.69 91.69
1: AA-ND 1.26 0.71 0.49 3.12 1.37 0.41 91.69 91.69
2: VA-VD 15.24 15.20 15.17 9.52 9.32 9.21 91.67 91.68
3: AA-VD 4.92 3.86 2.00 5.68 3.61 2.12 91.69 91.69
4: VA-AD 20.01 20.01 20.00 16.57 15.77 15.37 91.55 91.59
*5: AA-TD 3.26 2.07 1.18 13.19 11.82 10.69 91.56 91.46
*6: TA-AD 18.44 17.61 16.07 16.48 16.13 15.69 91.38 91.62
*7: AA-AD 6.93 6.39 5.73 9.77 8.74 7.64 91.68 91.58
*8: AA-AD 19.29 18.63 18.06 16.00 15.74 14.56 91.61 91.39
0: VA-ND 8.67 8.37 8.09 3.68 1.79 0.76 87.76 87.76
1: AA-ND 1.80 1.14 0.77 3.73 1.83 0.74 87.76 87.76
2: VA-VD 15.33 15.33 15.28 11.65 11.42 11.18 87.72 87.73
3: AA-VD 6.30 4.76 2.63 5.66 3.36 1.94 87.74 87.74
4: VA-AD 20.01 20.01 20.00 16.79 16.18 15.77 87.61 87.66
*5: AA-TD 4.53 3.31 2.16 13.59 12.65 11.39 87.58 87.52
*6: TA-AD 18.32 17.66 16.27 16.60 16.26 15.99 87.50 87.68
*7: AA-AD 7.12 6.69 6.00 10.67 9.89 8.92 87.74 87.62
*8: AA-AD 19.41 18.70 18.11 16.75 16.54 15.54 87.66 87.42
  1. 0.
    VA-ND – Vanilla Attack / No Defense: Baseline performance of attacks (BAGS & HSJA) out-of-the-box, without any active defense.
  2. 1.
    AA-ND – Adaptive Attack / No Defense: The degree at which each attack methodology can be optimized by being adaptive compared to the baseline.
  3. 2.
    VA-VD – Vanilla Attack / Vanilla Defense: The performance of our active defense, but the non-adaptive version operating with an empirically defined detection threshold.
  4. 3.
    AA-VD – Adaptive Attack / Vanilla Defense: Similar to scenario (2), but now the attack is adaptive.
  5. 4.
    VA-AD – Vanilla Attack / Adaptive Defense: The first scenario where the active defense is also adaptive, against the baseline adversary.
  6. 5.
    AA-TD – Adaptive Attack / Trained Defense: After the adaptive defense is optimized, its policy is fixed and an adaptive attack is trained against it.
  7. 6.
    TA-AD – Trained Attack / Adaptive Defense: The best policy found in the previous scenario is fixed and an adaptive defense is trained against it.
  8. 7.
    AA-AD – Adaptive Attack / Adaptive Defense: The first scenario where both agents learn simultaneously, making the environment non-stationary. In practice, the convergence will vary and depend on the chosen hyperparameters and rewards. Here we report the best-case for the attack.
  9. 8.
    AA-AD – Adaptive Attack / Adaptive Defense: Same scenario as before, but it reports the best-case for the defense instead.
For each successive scenario we evaluate with the most successful policy found, as this is best practice in Markov Games: the worst case opponent policy is fixed and then a counter to it is learned [34, 49]. Fixing other policies when computing a best response is representative of learning in cybersecurity environments that can contain a range of agents, with the additional benefit of converting the problem to single-agent that, as detailed in Section III-C, one can solve with RL.
V-B Results
For consistency and comparability between evaluations, the listed results are from the same 100 test episodes. The gap value denotes the $\ell_{2}$ perturbation that initially separates the starting and best adversarial example, averaged over the 100 episodes. A perfect defense would therefore acquire a value very close to the gap, while a perfect attack would never be able to reduce the perturbation to 00; it can be evaluated, however, relative to the baseline or other attacks. We tested the trained agents on budgets higher than 5K and we discovered that the depicted trends hold; however, to make the agent training tractable and the evaluation wider, we limit the maximum query budget per adversarial episode to 5K. Tables I & II report the results for CIFAR10 and MNIST respectively. The color gradient ranges from the gap value to the minimum value that can be achieved @ 5K queries. Overall, the empirical results help us extract and highlight several important insights, practical observations, general implications, and crucial takeaways for the broader AML field:
TABLE II: Mean $l_{2}$ perturbation for 1K, 2K, and 5K queries and accuracy on clean data for MNIST. The evaluation scenarios are identical to Table I.
Adv. Trained Scenario MNIST Gap = 10.62
BAGS HSJA Clean Acc.
1K 2K 5K 1K 2K 5K BAGS HSJA
0: VA-ND 5.28 5.26 5.24 3.63 3.06 2.61 99.37 99.37
1: AA-ND 2.75 2.57 2.47 3.61 3.04 2.58 99.37 99.37
2: VA-VD 8.74 8.28 7.94 5.74 5.73 5.73 99.35 99.20
3: AA-VD 3.39 3.29 3.19 3.58 3.12 2.74 99.33 99.31
4: VA-AD 10.45 10.45 10.45 9.87 9.87 9.87 99.31 99.30
5: AA-TD 3.54 3.27 3.10 4.19 3.54 2.80 99.32 99.20
6: TA-AD 9.85 9.77 9.76 9.99 9.99 9.99 99.28 99.34
7: AA-AD 4.95 4.92 4.92 7.07 6.97 6.76 99.31 99.36
8: AA-AD 10.62 10.62 10.62 8.78 8.71 8.66 99.32 99.36
0: VA-ND 5.28 5.27 5.25 4.65 4.02 3.40 99.15 99.15
1: AA-ND 3.28 3.07 2.94 4.60 3.97 3.35 99.15 99.15
2: VA-VD 8.67 8.44 8.08 5.71 5.70 5.70 99.14 98.89
3 AA-VD 4.14 4.04 3.81 4.69 4.35 3.87 99.13 99.15
4: VA-AD 10.39 10.39 10.39 10.07 10.07 10.07 99.09 99.12
5: AA-TD 4.01 3.80 3.61 4.94 4.44 3.89 99.11 99.12
6: TA-AD 10.05 9.97 9.95 10.20 10.20 10.20 99.07 99.06
7: AA-AD 5.51 5.50 5.50 5.47 5.16 4.99 99.01 99.13
8: AA-AD 10.62 10.60 10.60 10.12 10.12 10.12 99.10 99.01
  • A pervasive observation when comparing the upper and lower halves of each table is that adversarial training adds a small amount of robustness; otherwise, the practical effect of adversarial training is a tax on the attacker, forcing them to expend more queries for the same perturbation or having higher perturbation for the same query budget.
  • Evasive transformations interfere with the attack operation, and only when adaptively controlled together does the attack reach its full potential, as demonstrated by the difference between BAGS and HSJA in CIFAR scenarios 5-8.
  • Adaptive defenses outperform non-adaptive or model hardening, as even against the strongest attacker and the worst case for the defense – the 7th scenario – the perturbation for 5K queries is considerably higher.
  • The initial performance of an attack can be misleading: at first glance HSJA appears to be the better one but it is often outperformed by adaptive BAGS, especially in CIFAR-10 and against active defenses.
  • Not surprisingly, adaptive versions of the vanilla attacks perform better (Scenarios 1 & 0); especially BAGS, as is not well optimized out-of-the-box.
  • The advantage of adaptive attacks however is much more pronounced against active defenses, where they significantly outperform non-adaptive versions.
  • The performance of both attacks deteriorates considerably against active defenses, however they reach their full potential only when they are also adaptive.
  • Different attack fundamentals respond differently to active defenses; the gradient-based HSJA is naturally disadvantaged against similarity detection, especially on higher-dimensional data like CIFAR-10.
  • We observe that between consecutive scenarios, where an agent learns against the best opponent policy as previously discovered, the $l_{2}$ values oscillate: following a fixed policy enables the learning of an optimal counter to it.
  • When devising an adaptive defense for HSJA, it proved nearly impossible to engineer a state the agent can learn on by leveraging our knowledge of the attack and its geometric properties. What did prove effective, however, was pure computation [^1] This reminds us of Sutton’s Bitter Lesson [47]: the observation that progress is often driven by gains in computation rather than problem-specific expert knowledge.: to learn a state embedding from raw input, based on Contrastive Learning. These state embeddings also transfer exceedingly well to different attacks like BAGS.
  • The first time an active defense withstands adaptive attacks is in the 5th CIFAR-10 scenario, for BAGS and HSJA both. This is also the first scenario we had to employ evasive transformations to evade the active defense; subsequently, scenarios 5-8 in CIFAR-10 are the only with evasive transformations.
  • Evasive transformations combine differently with different attacks. All the offensive scenarios in BAGS recoup their performance almost completely. This is not the case for HSJA as the latter requires the preservation of the correct gradient information, something that had largely unsatisfactory solutions even with full white-box access to the model and to the true gradient.
V-C Empirical vs. Adaptive Evasion
So far our evaluations have focused on optimizing attack performance and evasiveness jointly. We are also interested in highlighting and measuring the potential of adaptively controlling the evasive actions without controlling the attack parameters themselves, against our proposed defensive agents. To compare how effective are adaptive evasive actions to the empirically defined ones, we evaluate them along with the approach of Chen et al. [16], configured as described in their work. we keep everything else exactly the same and the attacks themselves use their default, non-adaptive parameters: we want to isolate the effect of evasive actions alone. Table III compares the results between the empirical (yellow) and adaptive control (red) settings, and as we can observe our approach consistently outperforms the empirical definition of evasive transformations. Additionally, we note that the evasive agents were trained only in the upper half of the table, that is against non adversarially trained models: notably, adaptively controlling evasive actions transfers to other models too.
TABLE III: Performance comparison (in $l_{2}$ perturbation) between empirical input transformations (yellow scenarios) and the adaptive control of the same transformations (red scenarios) on CIFAR-10. Bottom half of the table is against adversarially trained models.
Scenario CIFAR-10 Gap: 20.01
BAGS HSJA
1K 2K 5K 1K 2K 5K
BA-VD 8.66 8.44 8.24 10.55 9.93 9.29
CA-VD 8.18 7.92 7.63 8.14 6.76 5.04
BA-TD 8.96 8.70 8.44 11.62 10.85 10.09
CA-TD 8.62 8.26 7.95 11.06 10.46 9.78
BA-AD 16.78 16.07 14.87 10.58 9.88 9.16
CA-AD 9.72 9.30 8.91 10.22 9.73 9.10
BA-VD 8.47 8.26 8.05 10.21 9.76 9.33
CA-VD 7.53 7.11 6.72 9.87 9.10 7.81
BA-TD 8.40 8.21 7.95 11.58 10.88 10.13
CA-TD 7.44 6.78 6.14 11.62 10.70 10.02
BA-AD 16.97 16.13 15.02 10.67 10.00 9.43
CA-AD 9.34 8.66 8.34 10.08 9.62 9.07
V-D Base Rate of Attacks
In all the evaluations so far we use a fixed probability $P(adv)=0.5$ that an incoming query is adversarial. To assess how the active and adaptive defenses (Scenarios 4 & 6) perform in the complete absence of attacks, we evaluate them with $P(adv)=0$ without any retraining; the results are shown in Table IV. We observe a small reduction in the accuracy on clean samples that can be attributed to the considerably different base rate of adversarial and benign queries. Note however that as the probability of each is an intrinsic property of each environment, if the base rate of attacks changes any defensive agent can be retrained to adjust to it.
TABLE IV: Clean accuracy on CIFAR-10 for scenarios 4 & 6, with standard and zero probability of attack. Yellow denotes normally trained models and red adversarially trained.
P(adv) BAGS4 BAGS6 HSJA4 HSJA6
0.5 91.55 91.38 91.59 91.62
0.0 90.91 90.95 90.48 90.86
0.5 87.61 87.50 87.58 87.68
0.0 87.02 87.11 86.70 86.98
VI Discussion
With AI-enabled decision-making becoming more pervasive and accessible, in domains like governance, finance, employment, and of course cybersecurity, more and more decisions are delegated to AI which becomes increasingly accountable for upholding safety and ethical constraints. As trustworthy AI is vital for the healthy functioning of whole ecosystems, this study highlights several security risks but also potential mitigations in these inherently black-box environments.
Our work has several implications for performing robust inference in the wild. While adversarial training remains the most reliable defense, the amount of robustness it imparts will vary and even be insufficient. We demonstrated how AI-enabled systems are susceptible to adaptive adversaries that devise new evasive techniques and control them jointly with other attack parameters. This has been achieved in the fully black-box case and against active defenses. Even more concerning is the level of threat that such systems face from AI-enabled adversaries, as it is straightforward to generalize Theorem III.1 to any other domain or modality. This rekindles the proverbial arms race, where as a consequence defenses should also be backed by equally capable AI.
Finally, our study highlights some aspects of the nature of adversarial examples. In black-box interactive environments they can be generalized by the notion of adversarial behavior, something that is more evident in domains where visual affinity is irrelevant, e.g., malware detection. Adversarial examples (and behavior) bring essence into question, what makes something, something; thus they can be approached from various epistemic perspectives. Computational approaches concern themselves with how to effectively acquire them; while their inherently competitive nature is already reflected in seminal works like shortcut learning [23] and invariant risk minimization [1]. Even in phenomenology – the philosophical study of the structures of experience and consciousness – the eidetic variation technique involves imagining and varying the features of an object under investigation to discover the (in)essential ones [31]. What these disparate approaches bring to light is that towards more general and robust AI researchers will have to reflect on the causal structure of perceptions and preconceptions, as well as on ways to automate this process.
Limitations. We note that while adversarial training has limitations, the robustness it provides is against all kinds of adversarial examples (under the same $\ell_{p}$-norm), while active defenses protect primarily against attacks they were trained with. As complementary approaches however, they can be used jointly. To keep the amount of evaluations practical, we narrowed the scope to targeted attacks and to $\ell_{2}$ as the more suitable norm for visual similarity. Targeted attacks are strictly more difficult to perform than untargeted, while in binary classification targeted and untargeted coincide; the latter is also the prevalent mode in cybersecurity contexts. Our framework, however, can accommodate any adversarial goal or metric.
Another simplifying assumption we make is that only one attack can take place at a time; however, the queuing methodology we use for incoming queries is readily extensible to handle concurrent attacks. Other possibilities to consider is how well an active defense handles attack methodologies it was not trained on, and to extend its training to multiple ones.
Finally, in the evaluation we focus on the breadth of scenarios and environments where agents learn concurrently, thus the number of datasets we experiment with is limited. However, we back our empirical study with an extensive theoretical analysis that gives support to our findings generalizing to any dataset or application domain.
Future Work. The AMG framework we introduce is general by design and can accommodate the learning of optimal offensive and defensive policies in any domain of interest beyond image classification. A promising path for future research is the extension of adaptive attacks and defenses to other domains and modalities, e.g. tasks like malware, bot, and network intrusion detection, precisely because our approach circumvents the obstacle of computing and mapping gradients to feasible perturbations and operates instead directly on the problem space [39]. Another compelling but formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing instruments to bypass defenses and thus imparting controllability to adversarial tasks. Finally, in our work we considered opponent agency as part of the environment; other domains, like malware analysis, might benefit from explicit opponent modeling.
VII Conclusion
With decision-based attacks becoming more pervasive in multiple domains, every AI-based decision-making process that exposes a queryable interface is inherently vulnerable. To aggravate matters, this vulnerability cannot be mitigated by employing model hardening approaches like adversarial training alone. To fully defend in the presence of such attacks, active and adaptive defenses are necessary, and we demonstrate how optimal defensive policies can be learned. However, the existence of such defenses elicits in turn adaptive attacks which are able to recover part of their original performance.
We perform a theoretical and empirical investigation of decision-based attacks and defenses under a unified framework we name “Adversarial Markov Games” (AMG). In self-adaptive, we introduce a novel twofold definition of adaptive: both devising new methods of outmaneuvering opponents and adapting one’s operating policy with respect to other agency in the environment. Furthermore, our adversarial policy gradient theorem indicates that any combination of adversarial goals, be it performance, stealthiness, disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any modality. As new attacks and defenses appear (and get broken) regularly, our evaluation methodology is applicable to any within the outlined scope by transforming the existing arms-race through their self-adaptive versions, thus ensuring accurate and robust assessment of their performance.
The AMG framework we introduce helps us reason on and assess the vulnerabilities of AI-based systems by disentangling the inherently complex and non-stationary task of learning in the presence of competing agency. We can model it as part of the environment, and thus simplify the task by computing a best response against the observed behavior. This is an important outcome for cybersecurity domains: as long as proper threat modeling is carried out, one can readily employ RL algorithms in order to devise optimal defenses; but only after they devised optimal attacks too. w
Appendix A Proofs
For a more intuitive understanding of the proofs, we first provide a high-level description of the attack fundamentals. BAGS [12] performs a random walk along the boundary between the adversarial and the non-adversarial regions, by first taking a random step orthogonal to the original image direction, then a source step towards it. The randomness in the directions searched is reduced by utilizing Perlin noise and masks computed on the difference between starting and original samples. HSJA [15] operates in 3 stages: a binary search that places the current best adversarial on the decision boundary, an estimation step that computes the gradient at that point of the boundary, and projection step along the estimated gradient. These steps repeat until convergence.
[Proposition III.1]
Proof.
The proof is constructed in two parts: first that an alternative decision function $D$ is required, and subsequently that a stateful representation of a query $x_{t}$ is also required. Let us denote by $x_{c},x_{g},x_{t}$ the original (unperturbed), the starting, and the current sample at step $t$ respectively. Given a target class $c_{0}\in m$ we can define a function:
$S_{x_{c}}(x_{t})=F_{c_{0}}(x_{t})-\underset{c\neq c_{0}}{\max}(F_{c}(x_{t}))$ (8)
HSJA operates in 3 stages that alternate until convergence or when the maximum query budget is reached: 1) A binary search between $x_{g}$ and $x_{c}$ that places $x_{t}$ on the decision boundary between the classes. 2) A gradient estimation stage that approximates $\nabla S(x_{t})$. 3) A “jump” step along the direction of the gradient $\nabla S(x_{t})$. For practicality we reiterate Eq. 9 of [15] which denotes that the gradient direction is approximated via the following Monte Carlo estimate:
$\widetilde{\nabla S_{x_{c}}}(x_{t},\delta)=\frac{1}{B}\sum_{b=1}^{B}\phi_{x_{c}}(x_{t}+\delta u_{b})u_{b}$ (9)
where $\{u_{b}\}_{b=1}^{B}$ are i.i.d. draws from the uniform distribution over the $d$-dimensional sphere, $\delta$ is a small positive parameter, and $\phi_{x_{c}}$ is the Boolean-valued function that all steps of HSJA rely on:
$\phi_{x_{c}}(x_{t})=\text{sign}(S_{x_{c}}(x_{t}))=\begin{cases}1&\text{if}\quad S(x_{t})>0,\\ -1&\text{if}\quad S(x_{t})\leq 0.\end{cases}$ (10)
Given an initial sample $x_{c}$, in search of adversarial examples HSJA applies the following update function that sequentially updates $x_{c}$:
$x_{t+1}=a_{t}x_{c}+(1-a_{t})\Biggl{\{}x_{t}+\xi_{t}\frac{\nabla S_{x_{c}}(x_{t})}{\|\nabla S_{x_{c}}(x_{t})\|_{2}}\Biggl{\}}$ (11)
where $\xi_{t}$ is a positive step size and $a_{t}$ is a line search parameter in $[0,1]$ s.t. $S(x_{t+1})=0$, i.e. the next query lies on the boundary. Now let us assume that the decision function $D$ is $\operatorname*{arg\,max}$, i.e. $D:\mathbb{R}^{m}\mapsto\mathbb{N}^{m},\>C(x)=D(F_{c}(x))=\operatorname*{arg\,max}F_{c}(x)$, then from Eq. 3 and Eq. 8 we have:
$S_{x_{c}}(x_{t})>0$ $\iff C(x_{t})=c_{0}$ (12)
$S_{x_{c}}(x_{t})<0$ $\iff C(x_{t})\neq c_{0}$
$S_{x_{c}}(x_{t})=0$ $\iff C(x_{t})=\{c_{0},a\},\>a\neq c_{0}$
$\implies S_{x_{c}}(x_{t})\leq 0$ $\iff C(x_{t})\neq c_{0}$
Let us define the function $\mathcal{I}$ of two variables:
$\mathcal{I}(a,b):=\begin{cases}1&\text{if}\quad a=b,\\ -1&\text{if}\quad a\neq b.\end{cases}$ (13)
We can rewrite Eq. 10 through and Eq. 13 and inequalities 12 as follows:
$\phi(x_{t})=\mathcal{I}(C(x_{t}),c_{0})$ (14)
Provided that the gradient estimation happens at the decision boundary where $S(x_{t})=0$, Theorem 2 of [15] guarantees that the estimation of stage 2 is an asymptotically unbiased direction of the true gradient:
$\widetilde{\nabla S_{x_{c}}}(x_{t},\delta)\approx\nabla S_{x_{c}}(x_{t}),\delta\rightarrow 0$ (15)
For $b_{t}=1-a_{t}$ and by plugging Eqs 14 & 15 in Eq. 9, and the result in 11, we get:
$x_{t+1}=a_{t}x_{c}+b_{t}\Biggl{\{}x_{t}+\xi_{t}\frac{\nabla S_{x_{c}}(x_{t})}{\|\nabla S_{x_{c}}(x_{t})\|_{2}}\Biggl{\}}$
$=a_{t}x_{c}+b_{t}\Biggl{\{}x_{t}+\xi_{t}\frac{\frac{1}{B}\sum_{b=1}^{B}\phi_{x_{c}}(x_{t}+\delta u_{b})u_{b}}{\|\frac{1}{B}\sum_{b=1}^{B}\phi_{x_{c}}(x_{t}+\delta u_{b})u_{b}\|_{2}}\Biggl{\}}$ (16)
$=a_{t}x_{c}+b_{t}\Biggl{\{}x_{t}+\xi_{t}\frac{\frac{1}{B}\sum_{b=1}^{B}\mathcal{I}(C(x_{t}+\delta u_{b}),c_{0})u_{b}}{\|\frac{1}{B}\sum_{b=1}^{B}\mathcal{I}(C(x_{t}+\delta u_{b}),c_{0})u_{b}\|_{2}}\Biggl{\}}$
Recall that we assumed $C(x)=\operatorname*{arg\,max}F_{c}(x)$, then the updates in Eq. (16) are guaranteed by Theorem 1 of HSJA [15] to converge to a stationary point $x_{b}$ of Eq. (5). Given a standard threshold $\epsilon$ on imperceptibility and over $N$ adversarial episodes, this implies that $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$. As $C$ is the only term in the expression (16) that is affected by the model, we reach a contradiction and that an alternative classifier $C^{\prime}$ is required, with $C(x)=D(F_{c}(x))$. As the discriminant function $F_{c}$ cannot change without retraining, it follows that $D^{\prime}\neq\operatorname*{arg\,max}$, so that for the adversarial example $x_{t}$ misclassified as $c_{0}$, Eq. 4 can return -1:
$C(x_{t})=c_{0}\Rightarrow\psi(x_{t})=-1$ (17)
$\therefore C(x_{t})=\hat{c},\>\hat{c}=\{c_{0},m\setminus c_{0},\varnothing\}$
where the empty decision $\varnothing$ denotes rejection and $\{m\setminus c_{0}\}$ denotes misdirection, i.e. intentional misclassification.
Decision-based attacks initiate from examples $x_{t}$ that belong to the target class $c_{0}$; while for $t=0$, $x_{t}$ is not adversarial yet, it still is part of an ongoing adversarial attack. To deter the attack, a perfect defender would have to misclassify/reject this example; yet if an identical but benign example $x_{n}$ was submitted, classifier $C$ should preserve its capacity to classify it correctly. For the identical examples $x_{n}=x_{t}$ we want $C(x_{n})\neq C(x_{t})$, something possible only with context. This context $\tau$, based on the sequence of incoming queries $x\in\{x_{t},t=0,1,2,...,N\}\cup x_{n}$, can be provided as an additional argument to the decision function $D(\tau,F_{c}(x))$ s.t. while $x_{n}=x_{t}\implies D(\tau,F_{c}(x_{n}))\neq D(\tau,F_{c}(x_{t}))$. ∎
[Theorem III.1]
Proof.
Let $H$ be a distance-based detector that based on a stateful representation $\mathcal{S}(\tau,x_{t})$ decides if a query $x_{t}$ is adversarial or not, and we assume the candidate generation policy $\pi_{\theta}^{\mathcal{A}}$ can influence this representation through intrinsic parameters like source step size, or applying transformations the underlying model is invariant to (V). For a predefined threshold $\delta$, if $\mathcal{S}(\tau,x_{t})<\delta$ then $H(x_{t})=1$ indicating that $x_{t}$ is part of an adversarial attack, otherwise $H(x_{t})=0$. Now for a number $N$ of attack episodes of fixed length $L$ and a reward function $r(\tau)=\sum_{t=1}^{L}(1-H(x_{t}))$, the objective to be maximized is the expected reward of the adversarial policy:
$\mathcal{J}(\theta)=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)]=\frac{1}{N}\sum_{i=1}^{N}r(\tau)=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-H(x_{t}))$ (18)
As it is not possible to compute the derivative of an expectation, by using the Policy Gradient Theorem [48] we can transform it to the expectation of the product of the reward and the gradient of the logarithm of the policy:
$\nabla\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)]$ $=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)\nabla\log\pi_{\theta}(\tau)]$ (19)
$=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[\sum_{t=1}^{L}(1-H(x_{t}))\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)]$
$=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-H(x_{t}))\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$
To maximize this expectation empirically we sample $e$ episodes from the environment, compute the gradient of the policy $\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$ and then update the policy with a learning step $a$, $\theta^{\prime}=\theta+\alpha\cdot\sum_{0}^{e}r(\tau)\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$. After this process has converged through gradient ascent, and as the maximum in Eq. 18 is attained for $\sum_{i=1}^{N}\sum_{t=1}^{L}H(x_{t})=0$, the obtained policy $\pi_{\theta}^{\mathcal{A}}$ will correspond to the optimally evasive one $\mathcal{E}$. ∎
[…]
Proof.
BAGS. This attack is in effect a gradual interpolation from $x_{g}$ towards $x_{c}$, by first taking orthogonal steps $x_{s}$ on the hypersphere around $x_{c}$ and then source steps towards $x_{c}$ in order to minimize $d(x_{c}-x_{b})$, where $x_{b}$ is the best adversarial example found so far. The source step parameter $\epsilon=(1.3-\min(\lambda_{n},1))\cdot c$ — with $\lambda_{n}$ the ratio of the $n$ last queries $x_{t}$ that are adversarial and $c$ a positive constant — controls the projection towards $x_{c}$:
$x_{t}=x_{s}+\epsilon\cdot(x_{c}-x_{s})$ (20)
[…]
HSJA. We denote the queries during gradient estimation as $x_{n}=x_{t}+\delta u$, $u\sim Uniform_{Sphere}(d)$, the ratio of those $x_{n}$ detected as adversarial by the active defense as $\eta\in[0,1]$, and the estimate $\widetilde{\nabla S_{x_{c}}}(x_{t},\delta)$ as $u_{t}$. We investigate the behavior of active defenses as the ratio of detections $\eta$ goes to 1.
For $\eta=1\implies\mathbb{E}[\phi_{x_{c}}(x_{n})]=-1$, and as $u_{b}$ are uniformly distributed, from Eq. (9) we get:
$\lim_{\eta\to 1}u_{t}$ $=\lim_{\eta\to 1}\frac{1}{B}\sum_{b=1}^{B}\phi_{x_{c}}(x_{t}+\delta u_{b})u_{b}=\frac{1}{B}\sum_{b=1}^{B}-u_{b}$ (21)
At the limit of detection we observe that the gradient estimate $u_{t}$ behaves like a uniformly drawn vector around $x_{t}$ of shrinking size. By the Law of Large Numbers, as $B$ increases the average direction of $u_{t}$ will align with the expected value: that is a random direction on the unit hypersphere. However, due to the $\frac{1}{B}$ term, the size of $u_{t}$ goes to 0. From Eq. (21) then we get: $\lim_{\eta\to 1}u_{t}=0$. The gradient estimation step is followed by the “jump” step that computes $x_{t+1}$ as follows:
$x_{t+1}=x_{t}+\xi{u_{t}}$ (22)
[…]
Appendix B On States & Rewards
[States]. In POMDPs a single observation cannot constitute a Markovian state as typically there is some of dependence on the state history. Partial observability is handled by either using recurrent architectures for the policy and value networks, or by engineering a state that includes past information: in this work we opt for the later.
For BAGS, the adversary uses an 8-dimensional state representation with the following information normalized in the range $[0,1]$ : current amount of queries $i$, average queries that are adversarial $a$, the initial gap $g$, the current gap $d$, the location $l=\frac{d}{g}$, the slope $s=m-l$ where $m$ is a moving average of the location, the frequency of improvement $f$, and $r$ which is a moving average of the perturbation reduction $n$, In HSJA the state representation is slightly different: $r=\frac{n}{g}$, and $f=\frac{1}{j}$ with $j$ number of jump steps in last iteration.
Regarding the defense, for HSJA (and to a lesser extent BAGS) it has been difficult to engineer a representative state for policies to effectively learn on. The knowledge of the attack internals and fundamentals, geometric properties and distances, model activations and logits, and any combination thereof, did not suffice. Ultimately we turned to pure computation to learn a state representation for the defensive agent. This representation is a 64-dimensional embedding of a CNN trained with triplet loss, on data generated by HSJA and benign queries, where the input is a tensor where the last query is subtracted from the 25 most recent adversarial queries and then stacked together.
[Rewards].
The concrete definitions of the rewards for each type of agent are:
[…]
  • For both BAGS and HSJA defenders, the aforementioned are the rewards when $x_{t}$ is adversarial; when it is benign, the reward is $R=1-h$ if the model responded correctly, otherwise $R=-1$.
TABLE V: Input Transformations.
Input Transformations Magnitude Probability
Brightness & Contrast 0 – 0.5 0 – 1
Random Horizontal Flip 0 – 1
Random Vertical Flip 0 – 1
Sharpness 0.8 – 1.8 0 – 1
Perspective 0.25 – 0.5 0 – 1
Rotation °0 – °180 0 – 1
Uniform Pixel Scale 0.8 – 1.2 0 – 1
Crop & Resize 0.6 – 1 0 – 1
Translation -0.2 – 0.2 0 – 1
Appendix C Models & Hyperparameters
The image classification models we use are ResNet-20 for CIFAR-10 and a standard 2 convolutional / 2 fully-connected layer NN for MNIST. For adversarially training models, we follow the canonical approach as described in [53]: the model is trained for 20 epochs, where the first 10 are trained normally and the last 10 on batches containing additional adversarial examples generated with 40 steps of PGD. For learning the similarity space, that is the metric space where defensive agents control the radius of interception around which a query is adversarial or not, we use a Siamese CNN. This network is trained with contrastive loss, where dissimilar examples are generated by adding Gaussian noise and performing evasive transformations on the input from the list in Table V. For the PPO agents trained for each scenario, we use the open source library Stable Baselines 3  [^2] https://github.com/DLR-RM/stable-baselines3. Policies are parameterized by a two fully-connected layer NN; the hyperparameter search space is shown in Table VI.
TABLE VI: Hyperparameter space for the PPO agent.
Hyperparameter BAGS HSJA
learning rate 3e-3 – 1e-4 3e-3 – 1e-4
episode steps 600 – 3000 1000 – 5000
total steps 1e5 – 4e5 2e4 – 2e5
batch size 32 – 128 32 – 64
buffer 2048 – 2048 256 – 2048
epochs 20 – 20 20 – 20
gamma 0.85 – 0.99 0.9 – 0.99

### v2

@@
 Despite considerable efforts on making them robust, real-world ML-based systems remain vulnerable to decision based attacks, as definitive proofs of their operational robustness have so far proven intractable.
-The canonical approach in robustness evaluation calls for adaptive attacks, that is with complete knowledge of the defense and tailored to bypass it.
+The canonical approach in robustness evaluation calls for adaptive attacks, with complete knowledge of the defense and tailored to bypass it.
 In this study, we introduce a more expansive notion of being adaptive and show how attacks but also defenses can benefit by it and by learning from each other through interaction.
@@
 We validate our observations through a wide theoretical and empirical investigation to confirm that AI-enabled adversaries pose a considerable threat to black-box ML-based systems, rekindling the proverbial arms race where defenses have to be AI-enabled too.
-Succinctly, we address the challenges posed by adaptive adversaries and develop adaptive defenses, thereby laying out effective strategies in ensuring the robustness of ML-based systems deployed in the real-world.
+Our approach outperforms the state-of-the-art black-box attacks and defenses, while bringing them together to render effective insights over the robustness of real-world deployed ML-based systems.
 I Introduction
@@
 As models cannot update their decision boundary in an online manner and in response to adversarial activity on their interface, there has to be a complement to model hardening: for instance active defenses such as rejection or misdirection [11, 12, 13].
-In this study we address a crucial gap, as in adversarial machine learning (AML) evaluating the robustness of defenses against oblivious, non-adaptive, and thus suboptimal attackers is inherently problematic [50, 19].
+In this study we address a crucial gap, as in adversarial machine learning (AML) evaluating the robustness of defenses against oblivious, non-adaptive, and therefore suboptimal attackers is inherently problematic [14, 15].
 We expand the conventional notion of adaptive, from adapted attacks that have an empirical configuration to bypass the defense, to include the capability to self-adapt, where attacks adaptively control their parameters and evasive actions together in response to how the model under attack and its defenses respond [16].
-We demonstrate theoretically and empirically how self-adaptive adversaries can modify their own policies through reinforcement learning (RL) to become both optimal and evade active detection.
-Notably, we find that this can be performed in a gradient-based manner even in fully black-box contexts [III.1], a capability that properly reflects the level of adversarial threat and does not overestimate the empirical robustness; attackers might compute gradients after all.
-This capability, however, enables active defenses in turn, as through rigorous threat modeling and by simulating self-adaptive attackers their full potential is uncovered and effective counter-policies can be learned.
-To express the call for adaptive evaluations in AML differently, a defense can only be as good as its adversary.
-This mutual interdependence generates the necessity for both attacks and defenses being self-adaptive, as well as the competitive, zero-sum game they jointly form.
-In summary, our research engages from two related perspectives: a) countering decision-based attacks against ML models, and b) the self-adaptation of such attacks, resulting in the following contributions:
-  • We demonstrate that active defenses against decision-based attacks are a necessary complement to model hardening but not sufficient.
-Active defenses are inevitably bypassed by self-adaptive attackers; and in turn this necessitates self-adaptive defenses too.
-  • To facilitate reasoning on self-adaptive attacks and defenses, we introduce a unified framework called “Adversarial Markov Games” (AMG).
-We demonstrate how attacks can optimize their policy and evade active detection at the same time; as a counter we propose a novel active defense and employ RL agents to adapt and optimize both.
-  • In an extensive empirical evaluation on image classification and across various adversarial settings, we validate our theoretical analysis and show that self-adaptation through RL outperforms baseline attacks, evasion techniques, and model hardening defenses like adversarial training.
-In the domain of black-box, decision-based AML, our work indicates that robust evaluations cannot consider attacks or defenses in isolation, and that each should have the capability to modify their operation through interaction and in direct response to other agency in the environment.
-II Related Work & Problem Statement
+We demonstrate theoretically and empirically how self-adaptive attacks can use reinforcement learning (RL) to modify their policies to become both optimal and evade active detection.
+Notably, this can be performed in a gradient-based manner even in fully black-box contexts [III.3], and is a capability that properly reflects the level of adversarial threat and in that way does not overestimate the empirical robustness; real attackers might compute gradients after all.
+This capability, however, enables active defenses in turn.
+Through rigorous threat modeling and by simulating self-adaptive attackers, their full potential is uncovered and effective counter-policies can be learned.
+To express the call for adaptive evaluations in AML differently, a defense is trustworthy only insofar the adversary is optimal.
+This mutual interdependence creates the imperative for both attacks and defenses to be self-adaptive, thereby establishing the competitive, zero-sum game that their interaction forms.
+Our research approaches the question of robustness from these two closely related perspectives: first, how to optimize decision-based attacks, and secondly what defensive options exist.
+In this manner, we explore both offensive and defensive strategies in depth, resulting in the following key contributions:
+  • We demonstrate that active defenses against decision-based attacks are a necessary but insufficient complement to model hardening.
+Active defenses are inevitably bypassed by self-adaptive attackers however, and necessitate self-adaptive defenses too.
+  • To facilitate reasoning on adaptive attacks and defenses, we introduce a unified framework called “Adversarial Markov Games” (AMG).
+We demonstrate how adversaries can optimize their policy and evade active detection at the same time; as a counter, we propose a novel active defense and employ RL agents to adapt and optimize both.
+For reproducibility and follow-up work, we open-source our code [^1] https://anonymous.4open.science/r/AMG-AD16.
+  • In an extensive empirical evaluation on image classification and across various adversarial settings, we validate our theoretical analysis and show that self-adaptation through RL outperforms vanilla black-box attacks, model hardening defenses like adversarial training, and notably both the state-of-the-art adaptive attacks and stateful defenses.
+Our work highlights that in the domain of black-box AML, robust evaluations should go a step further than adapting attacks: both attacks and defenses should have the capability to optimize their operation through interaction and in direct response to other agency in their environment.
+The remainder of the paper is structured as follows: Section II provides the necessary background on the domain and reviews the related work.
+Section III introduces and motivates our theoretical analysis of robustness under decision-based attacks.
+Section IV elucidates the threat model and the concrete design choices.
+In Section V we elaborate on our experimentation and analyze our results.
+We conclude with Section VI where we discuss insights, limitations and challenges.
+II Preliminaries
 In this work, we focus on the category of adversarial attacks known as decision-based, a subset of query-based attacks that operate solely on the hard-label outputs of the model and are regarded as a highly realistic and pervasive threat in AI-based cybersecurity environments.
-Despite the lack of the closed-form expression of the model under attack, given enough queries the effectiveness of such black-box attacks can match and even surpass that of white-box techniques like C&W [14].
+Despite the lack of the closed-form expression of the model under attack, given enough queries the effectiveness of such black-box attacks can match and even surpass that of white-box techniques [17, 15].
 II-A Attacks & Mitigations
@@
 Characteristically, preprocessing defenses are often bypassed by expending queries for reconnaissance [21].
-The partial exception to this rule is adversarial training [35].
+The partial exception to this rule is adversarial training.
 Given dataset $D={(x_{i},y_{i})}^{n}_{i=1}$ with classes $C$ where $x_{i}\in\mathbb{R}^{d}$ is a clean example and $y_{i}\in{1,...,C}$ is the associated label, the objective of adversarial training is to solve the following min-max optimization problem:
@@
 II-B Stateful Defenses
-All decision-based attacks share properties that can be useful in devising defenses against them, beyond adversarial training.
-Such a property is their inherent sequentiality: as they follow a query generation policy towards the optimal adversarial example, the generated samples are correlated.
-This might not be the case for the queries themselves however, as the adversary might adaptively employ transformations the model is invariant to, like the query blinding strategy in [16].
+All decision-based attacks share properties that can be useful in devising defenses against them, on top of adversarial training.
+Such a property is their inherent sequentiality: by following an attack policy towards the optimal adversarial example, the generated candidates are correlated.
+This might not be the case for the queries themselves however, as the adversary might employ transformations the model is invariant to, like the query blinding strategy in Chen et al.
+[13].
 This work is also the first to employ a stateful defense against query-based attacks.
-Another one is PRADA [30], a defense built against model extraction but which also works against model evasion.
-The efficacy of these approaches however rests on the assumption that queries can be consistently linked to uniquely identifiable actors – that also show limited to no collaboration – so that a buffer of queries can be built for each one of them.
+Another stateful defense is PRADA [22], built against model extraction but which also works against model evasion.
+The efficacy of these approaches however rests on the assumption that queries can be consistently linked (through metadata like IP or account, cf.
+LABEL:tab:comparison) to uniquely identifiable actors – that also show limited to no collaboration – so that a buffer of queries can be built for each.
 This limitation was recently addressed, together with the ensuing scalability issues in the Blacklight defense, by resourcefully employing hashing and quantization [23].
-It remains a similarity/fingerprinting defense however, thus vulnerable to circumvention if an adversary can find a query generation policy to preserve the attack functionality while evading detection; this was recently achieved by adapting existing attacks through the rejection signal that Blacklight returns [21].
-Ultimately, any stateful defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be representative only if the adversary has exhausted their adaptive offensive capabilities.
-From the mere interaction it is enabled from, stateful detection already points towards adversarial examples being something broader that includes the process of generating them, otherwise defined as adversarial behavior.
-Furthermore, as AML encompasses domains where visual similarity is of less significance [45], the ensuing question is where does the boundary lie exactly between benign and adversarial examples?
-In AI-based environments where adversarial activity is also present, we posit that adversarial behavior is a more general formulation, appropriate for describing and countering threats at the interface of interaction.
-It is also more intuitive in domains where behavior – as information unfolding over time – is the canonical description, e.g in malware or bot detection, where AI can imitate benign programs or human behavior respectively.
+It remains a similarity-based defense however, thus vulnerable to circumvention if an adversary can find a query generation policy that preserves attack functionality while evading detection; the recent OARS work achieved this by adapting existing attacks through the rejection signal Blacklight returns [24].
+Ultimately, any (stateful) defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be representative only if the adversary has exhausted their adaptive offensive capabilities.
+Figure 1: In AML, adaptive attacks are those with the capabilities (knobs) to bypass a defense; adaptive control is rather the precise tuning of all the known knobs.
+Against black-box systems, we can reformulate adaptive so that it signifies both.
+For instance in HSJA [7], radius, steps, and jump are parameters of the attack, while rotate and translate are transformations that can evade a similarity-based defense.
 II-C On Being Adaptive
-The correct way to evaluate any proposed defense is against adaptive attacks: that is, attacks with explicit knowledge of the inner mechanisms of a defense [50].
-This has been established in computer security as the stipulation that security through obscurity does not work: every defense has a potential workaround.
-If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, an active defense like stateful detection is the counterpart to decision-based attacks, and as we will demonstrate further in the paper, also the necessary complement to model hardening.
+The correct way to evaluate any proposed defense is against adaptive attacks, that is with explicit knowledge of the inner mechanisms of a defense [14].
+In computer security this is known as the stipulation that security through obscurity does not work, as the robustness of defenses should not rely on keeping their way of functioning secret.
+If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, active defenses like stateful detection are the counterpart to decision-based attacks, and as we will further demonstrate, also the necessary complement to hardening a model against them.
 At the same time, the level of threat that attacks pose is often unclear or not thoroughly evaluated.
-As demonstrated in the work of Croce et al.
-[19], it is very common that the parameters of an attack are suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness.
-This underestimation can be further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have.
-The practical effectiveness of attacks rests on the ability to adapt the policies that govern both their operation and their evasive capabilities in tandem.
-In AML, “adaptive” refers conventionally to attacks that have full knowledge of the inner workings of a defense; a term we will henceforth refer to as adapted.
+Previous work has demonstrated that loss functions and parameters of attacks are often suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness [15, 25].
+This underestimation is further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have.
+The practical effectiveness of attacks therefore rests on the ability to adapt the policies that govern their operation and their evasive capabilities in tandem.
+In AML, “adaptive” by convention refers to attacks with full knowledge of how a defense works and the tools to bypass it; we denote such attacks as adapted.
 In our work, we expand the term to include adaptive control, defined as the ability of a system to self-adapt: automatically reconfigure itself in response to changes in the dynamics of the environment in order to achieve optimal behavior [16].
-Typically, what is to be controlled is known in advance and well-defined.
-The moment however we consider adaptive evaluations, new controls are immediately implied: in a newly proposed defense, for instance rejection-based, these are controls that can bypass it, like rejection sampling.
-To flesh out the twofold meaning of adaptive, one has to both imagine new knobs [29], and discover their correct configuration that would lead to the optimal result.
-We conceptualize this expanded definition of adaptive, essential for having accurate evaluations in AML research, in Figure 1.
-Figure 1: In AML adaptive means to invent new knobs that can bypass a defense (lower open set); in control theory it means the precise tuning of all known knobs.
-In this work, we reformulate adaptive to signify both.
-In HSJA for example, radius, steps, and jumps are parameters of the attack, while rotate and translate are evasive transformations.
+One can think of adaptive control in the sense “attack optimization” is used by Pintor et al.
+[25], but for black-box systems.
+Typically, what is to be controlled is well-defined and known in advance.
+The moment however we consider adaptive evaluations, new controls are potentially implied: in a similarity-based defense for instance, such controls would be input transformations the model is invariant to.
+To flesh out the twofold meaning of adaptive, one has to both invent new knobs [26] (which is the conventional understanding of adaptive, and still very hard to automate), and dynamically control their correct configuration that would lead to the optimal result (self-adaptive).
+We conceptualize this more general definition of adaptive, essential for having accurate evaluations against decision-based attacks, in Figure 1.
 II-D Research Gap
-Prior work has focused on adapted attacks by incorporating general knowledge of any defenses and then empirically configured to evade it [14, 15, 9].
-Defenses also follow the same adapted paradigm, with parameters empirically defined and then fixed for deployment [16, 32].
-Our observation here is that neither attacks nor defenses are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and against other adaptive agents in it, something that comes with considerable limitations in security applications where the latter is a given.
-To bridge this gap, we provide theoretical treatment and empirical study of existing and novel methodologies adapting through direct interaction with their environment, and denote them as self-adaptive.
-Our work builds on a long line of prior research that focuses on different aspects of the competition between adversaries and defenses.
+Prior work has focused on adapted attacks, which incorporate general knowledge of any defenses and empirically configured to evade it [17, 7, 6].
+Defenses also follow the same adapted paradigm of empirically defined and fixed parameters [13, 23].
+Our observation is that neither of them are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and with respect to other adaptive agents in it, with clear limitations when the latter is a given, e.g.
+in cybersecurity.
+To bridge this gap, we provide a theoretical treatment and empirical study of existing and novel methodologies adapting through direct interaction with their environment, denoting them as self-adaptive.
+Our work builds on a long line of prior research that focuses on both sides of the competition between adversaries and defenses.
 Carlini and Wagner [17] show that evaluating existing attacks out-of-the-box is insufficient and that adapted white-box attackers can break defensive distillation.
-Bose et al[8] propose Adversarial Examples Games (AEG), a zero-sum game between a white-box attacker and a local surrogate of the target model family.
+Bose et al.
+[27] propose Adversarial Examples Games (AEG), a zero-sum game between a white-box attacker and a local surrogate of the target model family.
 At the equilibrium the attacker can generate adversarial examples that have a high success rate against models from the same family, constituting a zero-query, non-interactive approach for generating transferable adversarial examples.
-Pal et al[37] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium.
-Ren et al[42] explain adversarial robustness through low and high-order interactions between features.
-Yasodharan et al[56] study adversarial classification by formulating it as non-zero-sum hypothesis testing game.
-As the most relevant and representative setting for threats against deployed AI-based systems, we scope on decision-based, interactive attacks and defenses; we contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive (and thus also the defensive) potential is properly assessed.
-While “self-adaptive” represents the synthesis of adapted and adaptive control, for the remainder of the paper we employ the term “adaptive” interchangeably, and we use “adapted” to indicate what is conventionally known as adaptive evaluations in AML.
+Pal et al.
+[28] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium.
+Feng et al.
+[24] introduce OARS: adaptive versions of existing attacks that bypass Blacklight [23], the state-of-the-art stateful defense.
+To function, OARS presupposes the rejection signal that a defense like Blacklight returns; a strong assumption that as we show in this work does not have to hold for stateful defenses.
+As we demonstrate in section V and Table III, Blacklight can be bypassed without assuming rejection, while the novel stateful defense we introduce can fully withstand the OARS attack.
+As the most relevant and representative threat against real-world AI systems, in this work we scope on decision-based, interactive attacks and defenses.
+We contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive and thus also the defensive potential is properly assessed.
+In the remainder of the paper the term “adaptive” subsumes adaptive control, and is used interchangeably with “self-adaptive”.
+For what is conventionally known as adaptive evaluations in AML, we use the term “adapted”.
+To facilitate comparison, in LABEL:tab:comparison we highlight the most important aspects of our work as the synthesis of adaptive black-box attacks and defenses in a unified framework, and situate it with respect to other relevant and state-of-the-art works in AML.
+Note the importance for an attack to function without assuming rejection, and respectively for a defense to function without access to query metadata like IP addresses or accounts.
+TABLE I: Prominent decision-based attacks and defenses and their individual aspects. While all works are situated either on the offensive or the defensive side, ours brings these two together.
+Work Offensive Defensive
+Optimized Evasive Adaptive $\neg$¬Rejection Active Adaptive $\neg$¬Metadata Misdirection
+Boundary (2018) [6]
+BAGS (2018) [29]
+HSJA (2020) [7]
+OARS (2023) [24]
+Adv. Training (2017) [4]
+Stateful (2020) [13]
+Blacklight (2022) [23]
+AMG – Our work
 III Theoretical Framework
-In this work we engage from two perspectives that are interrelated: a) thwarting decision-based attacks, and b) adapting such attacks together with any evasive actions, as these are in a natural trade-off.
-Evaluating non-adaptive – in the expansive sense we outlined – attacks or defenses renders any results unreliable and incomplete [50].
-When offensive or defensive methodologies become adaptive, the environments that they reside in become non-stationary [28], putting further pressure on the IID foundations that ML builds on.
-We can approach this multi-agent interaction more generally as a competitive and sequential zero-sum game [34, 26, 8].
-To understand the implications of attacks and defenses turning adaptive, before the practical implementation we perform a theoretical investigation of their interactions around an ML-based system, and we initiate our analysis from adversaries as the agency generates all the subsequent reasoning.
-In the following sections, the notation and terms we introduce are highlighted in red.
+To investigate the robustness to evasion of real-world systems, two closely related perspectives are pertinent: a) thwarting decision-based attacks, and b) adapting attacks and evasive capabilities in tandem.
+For the latter, when attacks (and defenses) are evaluated in a non-adaptive manner, in the expansive sense we outlined in subsection II-C, results are rendered incomplete and untrustworthy [14, 25].
+Nevertheless, when offensive or defensive techniques adapt the environments that they inhabit become non-stationary [30], putting further pressure on the IID foundations that ML builds on.
+To understand the implications of such techniques adapting, we perform a theoretical analysis of the possible interactions around the exposed interface of an ML-based system, interactions that can be more generally considered as sequential zero-sum games [31, 32, 27].
+We initiate our analysis from adversaries, as the original factor and cause.
+In the following sections, introduced terms and notation are highlighted in orange.
 III-A Attacks
-The most compelling threat for deployed ML-based systems are hard-label, decision-based black-box attacks where no access is assumed to the model or its parameters, only the capacity to submit queries and receive hard labels.
+The most compelling threat that deployed ML-based systems face are decision-based black-box attacks, where no access is assumed to the model or its parameters, only the capacity to submit queries and receive hard-label responses.
 One of the first decision-based attacks was Boundary Attack [6].
-A large number of others followed – each inventive in its own way – that manage to improve the overall performance, typically measured as the minimum perturbation achieved for the minimal amount of queries submitted.
-Prominent examples of such attacks are HSJA [15], Guessing Smart (BAGS) [12], Sign-Opt [17], Policy-driven (PDA) [55], QEBA [33], and SurFree [36].
-White-box methodologies for constructing adversarial examples like C&W [14] cannot function in black-box environments where there is no closed-form description of the inference pipeline.
-To facilitate optimization, decision-based attacks commonly reverse the adversarial task by initiating from a sample belonging to the target class, as it can be considered an adversarial example with an unacceptably large perturbation.
-The task then regains the ability to be solved continuously as it allows for the minimization of the perturbation while staying on the adversarial side of the boundary.
-Decision-based attacks exhibit further commonalities in their functioning so that one can abstract through their particularities: given starting and original samples $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{g}$ and $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{c}$ respectively, the goal is to iteratively find adversarial candidates $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x_{t}$, until the distance $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\delta=d(x_{t},x_{c})$ is minimized.
-This process can follow different algorithmic approaches representing different geometrical intuitions, however to describe them generally we can use a candidate generation policy:
+A large number of others followed – each inventive in its own way – that manage to improve the overall performance, typically measured as the lowest perturbation achieved for the minimal amount of queries submitted.
+Prominent examples are HSJA [7], Guessing Smart (BAGS) [29], Sign-Opt [33], Policy-driven (PDA) [8], QEBA [34], and SurFree [35].
+White-box attacks like C&W [17] cannot function in black-box environments as there is no closed-form description of the inference pipeline.
+To facilitate optimization, decision-based attacks commonly initialize from a sample belonging to the target class, as it can be considered an adversarial example with an unacceptably large perturbation.
+This switch allows the task to be solved continuously, by minimizing the perturbation while always staying on the adversarial side of the boundary.
+Decision-based attacks share further common aspects in their functioning, ones that we can abstract through: given starting and original samples $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{g}$ and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{c}$ respectively, the goal is to iteratively propose adversarial candidates $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{t}$, until the distance $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\delta=d(x_{t},x_{c})$ is minimized.
+This process follows different algorithmic approaches that represent different geometrical intuitions; we can describe it more generally by means of a candidate generation policy:
 $\pi_{\theta}^{\mathcal{A}}=P\>(x_{t}|x_{g},x_{c},p^{\mathcal{A}},s^{\mathcal{A}})$ (2)
-that is conditional on $x_{g}$ and $x_{c}$, with $p^{\mathcal{A}}$ the parameters and $s^{\mathcal{A}}$ the state of the attack.
-Such parameters are the source and orthogonal steps for BAGS and the number of queries and radius for the gradient approximation in HSJA.
-Assuming the model answers always and truthfully, this process can be construed as an Markov Decision Process (MDP) to be solved by a policy that minimizes $d$ in the least amount of queries.
-Given a multinomial image classifier $\mathcal{M}$, the model under attack is based on a discriminant function $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ that for each input $x\in[0,1]^{d}$ generates an output $y:=\{y\in[0,1]^{m}|\sum_{c=1}^{m}y_{c}=1\}$ – a probability distribution over the $m$ classes.
-By definition, black-box environments provide no access to these probabilities; instead one can only observe the result of the classifier $C$ that returns the class with maximum probability:
+that given $x_{g}$ and $x_{c}$, with $p^{\mathcal{A}}$ the parameters and $s^{\mathcal{A}}$ the state of the attack, generates a candidate $x_{t}$.
+If we consider that attacks execute over discrete time steps, and assume the model always answers, then the attack procedure can be construed as a Markov Decision Process (MDP) to be solved, by finding the parameters $\theta$ that minimize $\delta$ in the least amount of queries.
+Consider now a multinomial image classification model $\mathcal{M}$ under attack, with a discriminant function $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, that for each input $x\in[0,1]^{d}$ generates an output $y:=\{y\in[0,1]^{m}|\sum_{c=1}^{m}y_{c}=1\}$ – a probability distribution over the $m$ classes.
+By definition, black-box environments preclude access to these probabilities; instead one can only observe the decision of the classifier $C$ that returns the highest probability class:
 $C(x):=\operatorname*{arg\,max}_{c\in[m]}F_{c}(x)=D(F_{c}(x))$ (3)
-with $D$ being $\operatorname*{arg\,max}$, the decision function.
-The goal in targeted attacks is to change the decision $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}c_{g}\in[m]$ for a correctly classified example $x$, to a predefined target class $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}c_{o}\neq c_{g}$.
-This process can be facilitated through a function $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\psi$ which given a perturbed example $x_{t}$ at step $t$, it returns a binary indicator of success:
+with $D$ being the decision function, here $D=\operatorname*{arg\,max}$.
+The goal in targeted attacks is to change the decision $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}c_{g}\in[m]$ for a correctly classified example $x$, to a predefined target class $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}c_{o}\neq c_{g}$.
+This process can be facilitated through a function $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\psi$ which given a perturbed example $x_{t}$ at step $t$, it returns a binary indicator of success:
 $\psi(x_{t})=\begin{cases}1&\text{if}\quad C(x_{t})=c_{o}\\ -1&\text{if}\quad C(x_{t})\neq c_{o}\end{cases}$ (4)
-As long as the model responds, $\psi$ can always be evaluated, it thus constitutes the essential mechanism upon which decision-based attacks build.
+As long as the model responds, $\psi$ can always be evaluated, it thus constitutes an essential mechanism upon which decision-based attacks build.
 The adversarial goal can then be described as the following constrained optimization problem:
 $\operatorname*{min}_{x_{t}}\>d(x_{t},x_{c})\quad\text{s.t.}\quad\psi(x_{t})=1,$ (5)
-where the distance metric $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}d$ is an $\ell_{p}$-norm, with $p\in\{0,1,2,\inf\}$.
-As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this optimization task highlights the imprecise nature of adversarial examples, something that is further aggravated in domains where visual affinity is of little importance.
-Customarily, successful or unsuccessful adversarial examples are delimited by an threshold $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\epsilon$ on perturbation, where $d(x,x_{t})\leq\epsilon$.
-While this constraint might not intuitively translate to non-visual domains, we nonetheless contend that minimal perturbation still remains a central property of adversarial examples.
-The most relevant attacks in cybersecurity contexts are black-box and decision-based, and they are becoming increasingly effective.
-In HSJA for example, its optimization is guaranteed to converge to a stationary point of Eq.
-(5), and given typical $\epsilon$ values on perturbation imperceptibility this translates to near-perfect attack success rates, even against adversarially trained models.
-The limitations of adversarial training against decision-based attacks can be attributed to the fundamentally out-of-distribution (OOD) nature of adversarial examples, as that makes the saddle point optimization of Eq.
-(1) intractable to solve exhaustively.
-Additionally, it is challenging to incorporate decision-based attacks during stochastic gradient descent: as approaches that navigate the decision boundary, the further the latter is from convergence, the less effective the attack is.
-Scalability is also an issue as typically in adversarial training a few steps (1-50) of a white-box attack – FGSM or PGD – are required, while decision-based attacks can take orders of magnitude more steps (queries) to produce an adversarial example.
-Decision-based attacks typically search for the optimal parameters $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\theta$ of generation policy (2), those that given $x^{i}_{c}$, with $i$ denoting the i-th adversarial episode, minimize Eq.
+where the distance metric $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}d$ is an $\ell_{p}$-norm, with $p\in\{0,1,2,\inf\}$.
+As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this optimization task highlights the fuzzy nature of adversarial examples, something that is further aggravated in domains where visual proximity is of little importance.
+Successful or unsuccessful adversarial examples are therefore delimited by an threshold $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon$ on perturbation, where $d(x,x_{t})\leq\epsilon$.
+While this constraint might not intuitively translate to non-visual domains, we nonetheless contend that minimal perturbation still remains a desideratum of adversarial examples.
+The fact that real-world attacks are black-box and decision-based does not make them any less effective.
+For instance, HSJA is guaranteed to converge to a stationary point of Eq.
+(5).
+Given typical $\epsilon$ values for imperceptibility, this results in high attack success rates, even against adversarially trained models.
+The limited effectiveness of adversarial training against decision-based attacks can be attributed to the fundamentally out-of-distribution (OOD) nature of adversarial examples, and the inherent challenges in the saddle point optimization problem in Eq.
+(1) that make it difficult for algorithms to converge to a global solution.
+Additionally, decision-based attacks are challenging to incorporate during stochastic gradient descent (SGD) because these attacks rely on boundary navigation, and the farther the model is from convergence, the less effective the attack is.
+Scalability is another concern: while adversarial training typically involves a few steps (1-50) of white-box attacks like FGSM or PGD, decision-based attacks can take orders of magnitude more steps (queries) to produce an adversarial example.
+Decision-based attacks search for the optimal parameters $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\theta^{*}$ of the generation policy (2), those that given $x^{i}_{c}$, with $i$ denoting the i-th adversarial episode, minimize Eq.
 (5) in expectation:
 $\operatorname*{arg\,min}_{\theta}\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})],\quad\text{s.t.}\quad\psi(x^{i}_{b})=1,$ (6)
-where $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}x^{i}_{b}$ is the best adversarial example generated by policy $\pi_{\theta}^{\mathcal{A}}$ during episode $i$.
-Given its dimensionality, it can be intractable to learn such a policy that modifies the input/problem space directly [39]; CIFAR-10, for instance, has more than 3K features to perturb.
-In AI enabled systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can assume that $F_{0}=F_{1}=...=F_{t},\forall t$.
+where $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x^{i}_{b}$ is the best adversarial example generated by policy $\pi_{\theta}^{\mathcal{A}}$ during episode $i$.
+Given the dimensionality, it can be intractable to learn a policy that modifies the inputs at the feature space directly [36]; CIFAR-10, for instance, has more than 3K features to perturb.
+In AI-enabled systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can therefore assume that $F_{0}=F_{1}=...=F_{t},\forall t$.
 While this is representative of real-world settings, it is also what enables adversaries to discover adversarial examples that were not identified beforehand.
-The existence of adversaries which follow a candidate generation policy introduces however a behavior which can be observed and utilized by a defensive methodology.
-Consequently, while model-hardening approaches like adversarial training are necessary, they can be insufficient in defending against decision-based attacks.
+The fact however that adversaries follow a candidate generation policy, introduces a querying behavior which can be observed and exploited by a defensive methodology.
+Consequently, while model-hardening approaches like adversarial training are necessary, they can be insufficient in defending from decision-based attacks.
 Proposition III.1.
 Let $F_{c}$ be the discriminant function of the adversarially trained model $\mathcal{M}$.
-Then given an HSJA attack, in order for $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]\geq\epsilon$, two capabilities are necessary: a) a decision function $D^{\prime}\neq\operatorname*{arg\,max}$, and b) additional context $\tau$ s.t for adversarial query $x_{t}$, $C(x_{t})=D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$.
+Then in order for $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]\geq\epsilon$ in HSJA, two capabilities are necessary: a) a decision function $D^{\prime}\neq\operatorname*{arg\,max}$, and b) additional context $\tau$ s.t for adversarial query $x_{t}$, $C(x_{t})=D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$.
 Intuitively, HSJA operates in 3 stages which repeat: a binary search that puts $x_{t}$ on the decision boundary, a gradient estimation step, and projection step along the estimated gradient.
-If the model always responds truthfully, the adversary will be able to accurately perform all these steps and converge to the optimal adversarial; without loss of generality, we can extend this intuition to decision-based attacks which navigate the boundary.
+If the model always responds truthfully, the adversary will be able to accurately perform all these steps and converge to the optimal adversarial; without loss of generality, we can extend this intuition to other decision-based attacks which navigate the boundary.
 Secondly, the model should be able to differentiate between two, otherwise identical, queries when one is part of an attack and the other is not, and this is possible through a stateful representation; see Appendix A for the proof.
 III-B Defenses
-Proposition III.1 shows that alternative classification policies are useful in the presence of decision-based attacks, for instance classification with rejection or intentional misdirection.
+Proposition III.1 indicates that alternative classification policies are necessary in the presence of decision-based attacks, for instance classification with rejection or intentional misdirection.
 In related work, rejection has been realized in the form of conformal prediction where model predictions are sets of classes including the empty one, or learning with rejection [11, 37]; while misdirection has emerged as a technique in adversarial RL and cybersecurity domains [10, 12].
@@
 Active defenses have direct implications on attacks themselves however.
-Let us assume an agent carrying out an active defense policy:
-$\pi_{\phi}^{\mathcal{D}}=P(\alpha|x_{t},s^{\mathcal{D}}),\>a\in\{0,1\}$ (7)
-with $x_{t}$ the query, $s^{\mathcal{D}}$ the state for the defense as created by past queries, and $\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\pgfsys@color@rgb@stroke{.75}{0}{.25}\pgfsys@color@rgb@fill{.75}{0}{.25}\alpha$ the binary decision: when the query is deemed adversarial, it is rejected by returning $\alpha=1$.
-If this policy is stationary the environment dynamics become in turn stationary and so, next to the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved.
+Let us now assume an agent carrying out an active defense policy:
+$\pi_{\phi}^{\mathcal{D}}=P(\alpha_{t}|x_{t},s^{\mathcal{D}}_{t}),\>\alpha\in\{0,1\}$ (7)
+with $x_{t}$ the query, $s^{\mathcal{D}}_{t}$ the state for the defense as created by past queries, and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\alpha$ the binary decision: when the query is deemed adversarial, it is rejected by returning $\alpha=1$.
+If this policy is stationary, the environment dynamics in turn become stationary too, and thus together with the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved.
 In two-player, zero-sum games, the moment an agent follows a stationary policy, it becomes exploitable through the reward obtained by an adversary [38].
-Active defenses, as a consequence of decision-based attacks, entail adaptive adversaries.
+Active defenses, a consequence of decision-based attacks, entail therefore adaptive adversaries.
 Proposition III.2.
@@
 An adversary can reason, as a corollary to Proposition III.1, that such defenses have to be in place as it is suboptimal not too.
-However, there is a second reason to consider adaptive attacks even in the absence of active defenses, as attack policies are often not optimal out-of-the-box.
-Adapting attack policies can be seen as performing hyperparameter optimization and as an approach has proven very effective in other black-box or expensive-to-evaluate contexts, like Neural Architecture Search and Data Augmentation [57, 38].
-The empirical results in Section V further indicate this correspondence between adaptive and self-optimizing, where the adaptive versions of attacks outperform the non-adaptive, more so against active defenses.
+However, there is a second reason to consider adaptive attacks even in the absence of active defenses, as attack policies are often suboptimal with their default, empirically defined parameters.
+Adapting attack policies is essentially the optimization of these parameters, and as an approach has proven very effective in other black-box or expensive-to-evaluate contexts, like Neural Architecture Search and Data Augmentation [39, 40, 41].
+Our results in Section V further indicate the correspondence between adaptive and self-optimizing, showing that adaptive attacks consistently outperform non-adaptive, particularly against active defenses.
 Consider now an active defense that is based on a similarity or a conformal metric.
-In the twofold meaning of adaptive we introduced in Section II, inventing control implies the capability to bypass a similarity based defense; adaptive control implies strategy instead, the active control of the available tools to evade the defense [1].
-Then the adversary’s task is to find the optimal adversarial policy that also evades rejection.
-The straightforward way to achieve this is by adapting policy (2) itself.
-Notably, and despite the discrete and black-box context, this optimization can be fully gradient-based [48].
-We now demonstrate how adaptively controlling decision-based attacks recovers the gradient-based solvability of the black-box optimization task despite neither the active defense nor the model itself being accessible in closed-form.
-Theorem III.1 (Adversarial Policy Gradient).
-Given model $\mathcal{M}$ with an active and stationary defense $\pi_{\phi}^{\mathcal{D}}$ (7), adaptive candidate generation policy $\pi_{\theta}^{\mathcal{A}}$ (2) that generates episodes $\tau$ of queries $x_{t}$, and a reward $r(\tau)=\sum_{x_{t}\in\tau}\neg\alpha$, the optimal evasive policy $\mathcal{E}$ is obtained via the gradient of the policy’s expected reward $\mathbb{E}_{\pi^{\mathcal{A}}_{\theta}}[r(\tau)]$.
+In the twofold meaning we introduced in Section II, adaptive attack implies the capability to bypass a similarity based defense; adaptive control implies optimization instead, the active tuning of all the available tools to evade the defense and minimize the perturbation (cf.
+Fig.
+1).
+The revised adversarial objective then is to find the optimal policy that also evades detection, and the way to achieve this is by adapting the candidate generation policy (2) itself.
+Notably, and despite the black-box and discontinuous nature of the task, this optimization can be fully gradient-based.
+Decision-based attacks can recover gradient-based solutions to their objective, despite neither the active defense nor the model itself being accessible in closed-form.
+For model $\mathcal{M}$, adversarial queries $x_{t}$, and active defense $\pi_{\phi}^{\mathcal{D}}$ making decisions $\alpha_{t}$, we can thus formulate the following:
+Proposition III.3 (Adversarial Policy Gradient).
+Given adversarial policy $\pi_{\theta}^{\mathcal{A}}$ (2) that generates episodes $\tau_{i}$ of queries $x_{t}$, and reward function $r(\tau_{i})=\sum_{x_{t}\in\tau_{i}}(1-\alpha_{t})$, the optimal evasive policy ${\pi^{\mathcal{A}}_{\theta^{*}}}$ is obtained by gradient ascent on the policy’s expected reward, $\nabla_{\theta}\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau_{i})]$.
 The proof is included in Appendix A.
-So far we have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) by mere observation of discrete model decisions, adaptive attacks can become optimal in terms of both evasion and efficiency.
-The last piece of the puzzle is turning active defenses also adaptive.
-Proposition III.3.
-An active defense $\pi_{\phi}^{\mathcal{D}}$ (7) achieves its optimal, i.e.
-maximizes the expectation $\mathbb{E}[\sum_{x_{t}\in\tau}P(\alpha|x_{t},s)]$, by adapting its policy against a stationary evasive policy $\mathcal{E}$.
-Proof.
-As offensive and defensive policies are strictly competitive, we can define the reward $\rho$ of the defensive policy as $\rho(\tau)=\sum_{x_{t}\in\tau}\alpha$, then by making $\pi_{\theta}^{\mathcal{A}}$ stationary and $\pi_{\phi}^{\mathcal{D}}$ adaptive in Theorem III.1, we can reason that the optimal defensive policy is determined via the gradient of the expected reward $\mathbb{E}_{\pi^{\mathcal{D}}_{\phi}}[\rho(\tau)]$.
-∎
+So far we have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) adaptive attacks can become optimal in terms of both evasion and efficiency by observing and adapting to the discrete model decisions.
+To complete the puzzle, the last piece is turning active defenses also adaptive.
+Corollary III.4.
+The active defense achieves its optimal $\pi_{\phi^{*}}^{\mathcal{D}}$ (7), i.e.
+maximizing expectation $\mathbb{E}[\sum_{x_{t}\in\tau_{i}}P(\alpha_{t}|x_{t},s^{\mathcal{D}}_{t})]$, by adapting its policy against the optimal evasive policy ${\pi^{\mathcal{A}}_{\theta^{*}}}$.
 Figure 2: Schematic model of an AMG environment.
 Due to the inherent uncertainty of behavior at either side of the interface, it is a partially observable RL environment mirrored for each agent where one’s decisions become the other’s observations.
+(I) denotes an adaptive attacker (cf.
+Fig.
+1), (II) model hardening (passive defense), and (III) an active defense.
+Proof.
+As offensive and defensive policies are strictly competitive, we can define a reward $\rho$ of the defensive policy as $\rho(\tau_{i})=\sum_{x_{t}\in\tau_{i}}\alpha_{t}$.
+Then by fixing the optimal evasive policy $\pi_{\theta^{*}}^{\mathcal{A}}$, and updating $\pi_{\phi}^{\mathcal{D}}$ with expected reward $\mathbb{E}_{\pi^{\mathcal{D}}_{\phi}}[\rho(\tau_{i})]$ gradients, at convergence the resulting ${\pi^{\mathcal{A}}_{\theta^{*}}}$ will be the optimal defensive policy.
+∎
 III-C Adversarial Markov Games
-By reasoning on both offensive and defensive capabilities, we highlighted why we cannot consider them independently from each other.
-As adaptive decision-based attacks and defenses are logical consequences of each other, by composing them we can form a turn-taking competitive game.
-A precise game-theoretic formulation requires the exact analytical description of the whole environment: the model, the players and their utility functions, as well as the permitted interactions and the transition dynamics, something exceedingly intractable in this context as well as most cybersecurity environments.
-Model-free methods however can learn optimal offensive and defensive responses directly through interaction with the environment [44, 43], obviating the need to learn a model of it or to find exact solutions the to bi-level optimization task like Eq.
-(1) that is inherently NP-hard to solve [11].
+By reasoning on both offensive and defensive capabilities, we highlighted why we cannot consider them independently.
+As adaptive attacks and defenses are logical consequences of each other, their composition results in a turn-taking competitive game.
+A precise game-theoretic formulation requires however the exact analytical description of the environment: the model, the players and their utility functions, as well as the permitted interactions and the transition dynamics, something exceedingly intractable in this as well as most other cybersecurity contexts.
+Model-free methods however can learn optimal offensive and defensive responses directly through interaction with their environment [12, 42], obviating the need to learn a model of it or to find exact solutions to the bi-level optimization problem in Eq.
+(1) that is NP-hard [43].
 To that end, Turn-Taking Partially-Observable Markov Games (TT-POMGs) introduced by Greenwald et al.
-[25] is a generalization of Extensive-Form Games (EFGs), widely used representations for non-cooperative, sequential decision-making games of imperfect and/or incomplete information; an apt formalism for decision-based attacks and defenses.
-Another nice property of TT-POMGs is that they can be transformed to equivalent belief state MDPs, significantly simplifying their solution.
+[44] is a generalization of Extensive-Form Games (EFGs), widely used representations for non-cooperative, sequential decision-making games of imperfect and/or incomplete information.
+TT-POMGs are an apt formalism for decision-based attacks and defenses, with the additional useful property that they can be transformed to equivalent belief state MDPs, significantly simplifying their solution.
 The competition underlying adversarial example generation has been explored in no-box and white-box settings [27, 45].
-We instead focus on decision-based, interactive environments which are assumed to have unknown but stationary dynamics: any agency present is considered part of the environment and therefore fixed in its behavior.
+We instead focus on decision-based, interactive environments, and assume that they have unknown but stationary dynamics: any agency present is considered part of the environment and therefore fixed in its behavior.
 By folding the strategies of other agents into the transition probabilities and the initial probability distribution of the game, an optimal policy computed in the resulting MDP will correspond to the best-response strategy in the original TT-POMG.
 The congruence between TT-POMGs and MDPs is useful not only from a theoretical perspective, but also for its practical implications in the security of ML-based systems: provided that adversarial agents and their capabilities can be identified through rigorous threat modeling, computing the best response strategy in the simulated environment will correspond to the optimal defense.
-This environment that encompasses adversarial attacks, adversarial defenses, and benign queries, can be construed as an Adversarial Markov Game (AMG) – a special case of TT-POMG – and is depicted in Figure 2.
+The environment that encompasses adversarial attacks, adversarial defenses, and benign queries, we describe it as an Adversarial Markov Game (AMG) – a special case of TT-POMG – and is depicted in Figure 2.
 Formally, we represent AMG as a tuple $\langle i,S,O,A,\tau,r,\gamma\rangle$
@@
 The game is sequential and turn-taking, so each player $i$ chooses an action $a$ from their set of actions $A^{i}$ which subsequently influences the observations of others.
-We have shown that an adaptive defense policy $\pi^{\mathcal{D}}_{\phi}$ is necessary to deter decision-based attacks, and as a consequence the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$ has to be adaptive in turn.
-As without implausible assumptions one cannot assume access to the exact state of the other agent, the states $O^{\mathcal{D}}$, $O^{\mathcal{A}}$ are partial observations of the complete state $\mathcal{S}$ of the full game.
-When human agents compete by holding beliefs about each other, they engage in recursive reasoning that in theory of mind is encountered as [I believe that [my opponent believes [that I believe…]]].
+We have shown that an adaptive defense policy $\pi^{\mathcal{D}}_{\phi}$ is necessary to deter decision-based attacks, and that consequently the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$ has to be also be adaptive.
+As with plausible assumptions we cannot assume access to the exact state of the other agent, the states $O^{\mathcal{D}}$, $O^{\mathcal{A}}$ are partial observations of the complete state $\mathcal{S}$ of the full game.
+When the competing agents that hold beliefs about each other are human, they engage in recursive reasoning expressed as [I believe that [my opponent believes [that I believe…]]].
 In the study of opponent modeling, considering other agent policies as stationary and part of the environment is equivalent to 0th level recursive reasoning: the agent models how the opponent behaves based on the observed history, but not how the opponent would behave based on how the agent behaves [46, 47].
-However, as AMGs can be solved by single-agent RL algorithms, we consider more involved recursive reasoning out of scope and perform the empirical evaluation without building explicit models of opponent behavior.
+In this work we consider more involved recursive reasoning out of scope, as AMGs can be solved by single-agent RL algorithms, and perform the empirical evaluation without building explicit models of opponent behavior.
 IV Environment Specification
-The empirical study we conduct in Section V constitutes diverse instantiations of the general theoretical framework introduced in Section III.
-Naturally, when working forward from the general to the particular concrete design choices have to be made when specifying the latter, and these choices can have considerable influence on the results.
-In this section we thus outline the components of the environment: adversaries, defenders, possible interactions and various mechanisms that underscore them.
-While extensible to more players, our AMG framework describes a two-player competitive game, so we assume that at a given moment only one attack takes place.
-From the defensive perspective, incoming queries can be either benign (clean) or part of an attack.
-A frequently made assumption that influences the effectiveness of stateful detection approaches, however, is the ability to attribute queries to UIDs, e.g., an IP address or a user account.
-Adversaries can create multiple accounts, use VPNs, collude between them, or in fact accounts or IP addresses might not even be necessary to query the model.
-To address that, we treat queries irrespective to their source, a strictly more challenging setting for stateful defenses.
-The environments we build operate solely on the content of queries and not on any other metadata, similar to [32].
-Like this work, we reframe the task of detecting and banning UIDs of adversaries with two notable differences.
-Instead of rejecting queries, a decision that in itself can assist the adversary in optimizing, we misdirect by returning another decision, e.g.
-the second highest class.
-Secondly, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution so that is not actually trivial to tell them apart from adversarial noise.
-Similarity.
-As decision-based attacks are governed by a generation policy, successive queries exhibit patterns and degrees of similarity that can be quantified by an appropriate $l_{p}$ norm.
-However such a norm should not be computed on the original inputs as an adapted adversary can ”invent control” (see Fig.
-1) by employing evasive transformations the model is invariant to and bypass the similarity detection.
-To account for such adversarial strategies, we train a Siamese network with contrastive loss in order to learn an latent space $\mathcal{L}(\cdot)$ where similar inputs are mapped close together, unaffected by added noise or transformations on the inputs.
-To statefully characterize queries we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign/undetected queries.
+The empirical study we conduct in Section V reflects diverse instantiations of the general theoretical framework introduced in Section III.
+When working forward from the theoretical to the practical, concrete design choices have to be made when specifying the latter, choices that can have considerable influence on the results.
+To elucidate our proposed robustness evaluation methodology, in this section we provide the concrete details on the threat model and the environment.
+[Threat Model].
+Our AMG framework describes a two-player competitive game; while extensible to more players, in this work we assume that at a given moment only one attack takes place.
+From the defensive perspective, incoming queries can be either benign or part of an attack.
+An assumption that influences the effectiveness of stateful detection is that queries can be attributed to UIDs, e.g., an IP address or a user account.
+However, adversaries can collude, create multiple accounts, use VPNs, or in fact accounts and IP addresses might not even be required to query the model.
+To address this, we treat queries irrespective to their source.
+This is a strictly more challenging setting for stateful defenses, where we operate solely on the content of queries and not on any other metadata, similar to [23].
+Unlike Blacklight however, instead of rejecting queries, something that by itself can provide more information to the adversary and thus facilitate evasion (cf.
+OARS [24], LABEL:tab:comparison), we misdirect by returning the second highest probability class.
+Furthermore, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution and so that is not trivial for a defense to tell adversarial noise apart.
+In summary, the black-box threat model we consider is delineated as follows:
+  • Assets: Trained and deployed model $\mathcal{M}$ with corresponding weights ${w}$.
+  • Agents: Adversary / Defender / Benign user.
+  • Adversary Goal: Generate minimal perturbation adversarial examples in as few queries as possible, while evading the defense.
+  • Defender Goal: Stop the adversary from generating adversarial examples, while preserving the correct functionality of the model $\mathcal{M}$ on benign users.
+  • Adversary Knowledge: The model $\mathcal{M}$ is known as the black-box function that transforms inputs $x\in[0,1]^{d}$ to outputs $c\in[m]$, $m$ being the number of classes.
+The weights ${w}$ and the closed-form expression of $\mathcal{M}$ are unknown, as unknown is if an active defense $\pi_{\phi}^{\mathcal{D}}$ exists or not.
+  • Defender Knowledge: The defender observes only the content of incoming queries, without knowing if they come from a benign user or the adversary.
+  • Adversary Capabilities: Adapt the parameters of the attack and optionally any evasive transformations; in essence, optimize the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$.
+  • Defender Capabilities For each query $x$, decide between answering truthfully with the true prediction $C(x)=c_{T}$ or misdirect with the second highest probability class $c_{S}$.
+Figure 3: Misdirection in a two-dimensional decision boundary.
+The adaptive defense controls a single parameter, the hypersphere radius around $k_{0}$ (the last known adversarial); for queries $x_{t}$ that fall within this hypersphere the model responds with a non-adversarial decision.
+$x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
+[Similarity].
+Decision-based attacks typically follow a policy that generates successive queries: these exhibit degrees of similarity which can be quantified by an appropriate $l_{p}$ norm.
+If that norm is computed on the original inputs however, an adversary can adapt (see Figure 1) by employing evasive transformations the model is invariant to and bypass the similarity detection.
+To account for this capability, we train a Siamese network with contrastive loss in order to learn a latent space $\mathcal{L}(\cdot)$ where similar inputs are mapped close together, unaffected by added noise or transformations on the inputs.
+For the stateful characterization of queries, we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign or undetected ones.
 […]
-Figure 3: Intuition of misdirection in a two-dimensional decision boundary.
-The adaptive defense controls a single parameter: the radius around $c_{0}$ (the last known adversarial) where queries $x_{t}$ side are intercepted and the model responds with a non-adversarial decision.
-$x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
-Adaptivity.
-No evaluation in AML is complete without considering adaptive adversaries: that is with the instruments to bypass the defense, but not necessarily with their optimal configuration.
-As all stateful defenses so far are similarity based, to bypass them our intuition points towards input transformations the model is invariant to.
-Formally, for a given query $x_{t}$ we want to compute a transformation $x_{t}^{\prime}=T(x_{t})$ so that $||x_{t}^{\prime}-x_{t}||_{2}\gg||x_{t}-x_{t-1}||_{2}$ while $F(T(x_{t}))=F(x_{t})$.
+[Adaptivity].
+No evaluation in AML is complete without considering adaptive adversaries; a notion we expand in this work, that is with the instruments to bypass the defense and their optimal configuration.
+As stateful defenses are so far similarity based, to bypass them intuition points towards input transformations the model is invariant to.
+For a given query $x_{t}$ we want to compute a transformation $x_{t}^{\prime}=T(x_{t})$ so that $||x_{t}^{\prime}-x_{t}||_{2}\gg||x_{t}-x_{t-1}||_{2}$ while $F(T(x_{t}))\approx F(x_{t})$.
 Depending on magnitude and composition of transformations $T$, the identity $F(T(x_{t}))=F(x_{t})$ might not always hold.
 As we also demonstrate in Section V, $T$ interferes with the perturbations the adversarial policy generates itself: the performance and evasiveness of an attack are thus in a natural trade-off.
-So what is the correct composition of transformations $T$ to apply?
+At this point we should inquire what is the correct composition of transformations $T$ to apply.
 When shall $T$ be applied, and how does it affect the attack fundamentals?
-The possible transformations can be considered as set of additional controls, and like attack parameters they themselves can be underperforming out-of-the-box [19].
-Thus the combined control of attack and evasion parameters is a prerequisite to properly assess the strength of a defense: their empirical configuration is often suboptimal against a particular defense and a fine balance is often required between evasion and preserving attack performance.
-This trade-off illustrates why the twofold definition of adaptive we trace in this work is necessary in AML evaluations: first to impart controllability to the task through the definition of what can be controlled, and then to find the precise optimal configuration and strategy of the attack.
-Agents & Environments.
-Unlike the vast majority of competitive games, in AMGs the two players have different action and state sets.
-AMGs are also asymmetric in the playing cadence: while the defender plays every round, it takes one to several rounds for the adversary to play; HSJA for example is controlled on the iteration rather on the query level.
-The learning task is complicated further given that the experience upon which each agent learns arrives only after the opponent moves.
-We address these complications by developing custom learning environments where agents can handle asymmetries and the delayed experience collection by training asynchronously.
-For the environments and agents respectively, we build on the OpenAI Gym and Stable Baselines 3 libraries  [10, 41].
-The exact definitions of the RL states and rewards we experimented with are included in Appendix B.
-States & Actions.
-For the definition of the states we used and the rationale behind them, we point the reader to Appendix B.
+The transformations $T$ can be considered as a set of additional controls, and like attack parameters they themselves can be suboptimal out-of-the-box [15].
+Thus the combined control of attack and evasion parameters is a prerequisite to properly assess the strength of a defense.
+Their trade-off illustrates why the twofold definition of adaptive is necessary in AML evaluations: first to impart the tools to accomplish to the task through the definition of what can be controlled, and then to find the precise optimal configuration and strategy of the attack.
+[Agents & Environments].
+Unlike common competitive games, in AMGs the two players have different action and state sets.
+AMGs are also asymmetric in the playing cadence: while the defender plays every round, the adversary might wait one to several rounds; HSJA for example is controlled on the iteration rather on the query level.
+Training is complicated further given that the experience upon which each agent learns arrives only after the opponent moves.
+We address these complications by developing custom learning environments for agents with asymmetries, with delayed experience collection, and asynchronous training, build with the OpenAI Gym and Stable-Baselines3 libraries.
+[States & Actions].
+For the definition and the rationale behind the states we use, we point the reader to Appendix B.
 For actions, we control BAGS through 4 parameters: orthogonal step size, source step size, mask bias, and Perlin bias.
 HSJA is controlled by 3: the gradient estimation radius, the number of estimation queries, and the jump step size.
-All evaluations start from controlling these fundamental attack parameters only.
-If the active defense proves impossible to defeat, we introduce additional knobs corresponding to the evasive actions that control the magnitude and probability of transformations on the input, with the goal to evade detection while preserving semantic content and hence the correct classification.
+All evaluations start from controlling these attack parameters only; if the active defense proves impossible to defeat, we introduce additional knobs that control the magnitude and probability of transformations on the input, with the goal to evade detection while preserving semantic content and hence the correct classification.
 The range of transformations we experimented with as well as their magnitude and probability are listed in Table V.
-Finally, in both BAGS and HSJA the active defense corresponds to an 1-dimensional continuous action that controls the radius of confinement, as depicted in Fig 3.
-Rewards.
-The success of any RL task relies heavily on how it is rewarded.
-Engineering an effective reward function is often non-trivial and can contain many intricacies, as reward hacking and specification gaming are common phenomena and the learned behavior can vary significantly [3].
-For adversaries, the rewards we experimented with are variations on minimizing the distance to the original example, while defenders are rewarded and penalized for intercepting adversarial and benign queries respectively.
-More specifically, BAGS terms include tn the distance gains towards the original example (also relative to the initial gap) and on their frequency.
-For HSJA, additional terms include penalties for the amount of gradient estimation queries, and rewards on the precision of gradient estimation and jump steps.
-Defensive rewards for BAGS include keeping the best adversarial or the average query close to the starting point, minimizing the size of the step, and correctly intercepting queries.
-For HSJA, additional terms include penalties for not intercepting queries at the binary search or jump steps.
-The reward functions in closed-form are shown in Appendix B.
+Finally, in both BAGS and HSJA the active defense consists of an 1-dimensional continuous action that controls the radius of confinement $a$, as depicted in Fig 3.
+[Rewards].
+Success in an RL task relies heavily on how it is rewarded.
+Engineering an effective reward function is non-trivial and hides intricacies, as reward hacking and specification gaming are common phenomena and the learned behavior can vary [48].
+For adversaries, the rewards we experimented with are variations on minimizing the distance to the original example – with extra reward shaping based on the fundamental operation of each attack – while defenders are rewarded or penalized for intercepting adversarial or benign queries respectively.
+The rewards are described in Appendix B.
 V Evaluation
-As the empirical counterpart of our theoretical investigation, we perform an extensive evaluation by defining a range of scenarios that reflect all possible and realistic combinations between adversarial attacks and defenses.
-Concretely, the research questions corresponding to the theoretical insights and we would like to validate are: 1) Are active defenses a necessary complement to model hardening and to what extent?
-2) Are attacks more threatening when adaptive, i.e., do they outperform their vanilla versions and can they evade active detection?
+For evaluation, we define a range of scenarios intended to reflect all possible and realistic combinations between adversarial attacks and defenses, and their adaptive versions.
+Concretely, the research questions we want to validate are: 1) Are active defenses a necessary complement to model hardening and to what extent?
+2) Are attacks more threatening when adaptive, i.e., do they outperform their vanilla versions and evade active detection?
 3) If yes, can active defenses recoup their performance by also turning adaptive?
-Metrics.
-The first metric we use is the $\ell_{2}$ norm between perturbed and original samples, over ASR (Attack Success Rate), for several reasons: 1) it is significantly more fine grained than a binary indicator of success, thus a better choice for comparing baseline attacks and defenses, their adaptive versions, and to each other; 2) it is not based on an arbitrary perceptual threshold, thus it is more informative of the full range of behavior between perfect attack and perfect defense, 3) a threshold moved around can yield widely varying results, perturbation size is an unaffected global metric, and 4) our RL agents learn optimal policies (also) by using this distance as a signal, where the adversary is rewarded by minimizing it and the defender by maximizing it.
+[Metrics].
+We employ ASR (Attack Success Rate) and $\mathbf{L_{2}}$ norm of the perturbation.
+For the former we set a fixed threshold of 3 for consistency between experiments, while the latter is a more fine-grained metric well suited for comparing baseline attacks, defenses, and their adaptive versions, as it is not based on an (arbitrary) perceptual threshold that can yield widely varying results when moved.
 The budgets we evaluate over are 1K, 2K and 5K queries.
-As robustness and classification accuracy are in a natural trade-off, the second metric of interest is the accuracy on benign samples (Clean Acc.) that the original model and the active defense achieve together.
+As robustness and classification accuracy are typically in trade-off, the third metric of interest is the benign sample accuracy (Clean Acc.) that the original model and the active defense achieve together.
 V-A Evaluation Setup
-The explicit goal for the agents is – by querying the model and being queried – to learn offensive or defensive policies that are general: they transfer to any other evasion task.
+Our goal is to learn offensive or defensive policies that are general: they transfer to any other evasion task.
 Thus after training and validating the agents, the final performance is reported on a fixed hold-out set of 100 adversarial episodes where the starting and original samples are selected at random.
 As is best practice in AML, candidate samples are only those that are correctly classified by the model.
-For each scenario we perform a limited hyperparameter and reward function exploration (max 30 trials), with the intention to root out poor combinations rather than exhaust the search space.
-These are described in more detail in Appendix C.
-The black-box attacks we render adaptive and evaluate are BAGS and HSJA, as they represent two fundamentally different approaches, are highly effective, and show the highest evasion potential [32].
-BAGS is a search-based methodology and operates in a stochastic manner where every query submitted is a new and potentially better adversarial example.
-Instead, HSJA is deterministic, composed of 3 different stages, and the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means to approximate the gradient at the decision boundary.
+For each scenario we perform a limited hyperparameter and reward function exploration (max 30 trials), with the intention to root out poor combinations rather than exhaust the search space, described in more detail in Appendix D.
+The black-box attacks we render adaptive and evaluate are BAGS and HSJA, as they represent two fundamentally different approaches, are highly effective, and have the highest evasion potential [23].
+BAGS is a stochastic, search-based method where every query submitted is a new and potentially better adversarial example.
+Contrastively, HSJA is deterministic and composed of 3 different stages where the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means to approximate the gradient at the decision boundary.
 In training and evaluation, the adversarial game is played as follows: the adversary starts by submitting a query, then the defender responds either truthfully (the true model prediction) or by misdirecting (the second highest probability class).
@@
 In either case, it is the defender’s turn; in testing they are also oblivious to the nature of the query and know only the content.
-For the majority of experiments we fix $p=0.5$; we also evaluate adaptive defenses when no attack is present ($p=0$) in subsection V-D.
-The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two different datasets – MNIST and CIFAR10 – and over two models with the same architecture but with different training regimes: one with adversarial training and one without.
-The transition from single-agent to multi-agent RL hides challenges however.
-In every scenario, we approach the AMG as a belief-state MDP (the requirement of knowing the exact opponent policies is relaxed) and use PPO [43] agents to discover optimal policies that will also constitute best responses for the full AMG [54].
-However, learning independently of other agency breaks the theoretical guarantees of convergence [52] – this can be observed in the sole scenario where both agents learn simultaneously and we thus report them separately: 7 & 8 AA-AD.
-The full list of scenarios is:
-TABLE I: Mean $l_{2}$ perturbation for 1K, 2K, and 5K queries for CIFAR-10, for normally and adversarially trained models.
+All experiments are performed with $p=0.5$; we also evaluate our trained defense when no attack is present ($p=0$) in Appendix C.
+The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two different datasets – CIFAR10 and MNIST – and over two models with the same architecture but with different training regimes: one with adversarial training and one without.
+The transition from single-agent to multi-agent RL hides challenges: we approach the AMG as a belief-state MDP (the requirement of knowing the exact opponent policies is relaxed) and use PPO [42] agents to discover optimal policies that will also constitute best responses for the full AMG [47].
+However, learning independently of other agency breaks the theoretical guarantees of convergence [49], eg.
+in scenarios 7 & 8 where both agents learn simultaneously.
+The complete list of scenarios is:
+TABLE II: ASR and mean $L_{2}$ perturbation for 1K, 2K, and 5K queries for CIFAR-10, against normally and adversarially trained models.
 Clean Acc.
-records the accuracy of the base model and any defenses present on benign queries; in the first two scenarios (no active defense) the baseline clean accuracy is recorded.
+reports the accuracy on benign queries of the base model plus any defenses present; in the first two scenarios (no active defense) the baseline clean accuracy is reported.
 Yellow scenarios denote the baseline attack performance, green denote active and/or adaptive defenses, and red denote adaptive attacks.
-The green-to-red heatmap is a renormalised range of 0 to 100, where 0 is perfect defense (the gap value) and 100 is perfect attack (lowest value for 5K queries).
-The asterisk denotes scenarios where input transformations were used for evasion.
+The asterisk denotes where input transformations were used for evasion.
 Adv. Trained Scenario CIFAR-10 Gap: 20.01
 BAGS HSJA Clean Acc.
-1K 2K 5K 1K 2K 5K BAGS HSJA
-0: VA-ND 8.15 7.64 7.22 3.40 1.45 0.42 91.69 91.69
-1: AA-ND 1.26 0.71 0.49 3.12 1.37 0.41 91.69 91.69
-2: VA-VD 15.24 15.20 15.17 9.52 9.32 9.21 91.67 91.68
-3: AA-VD 4.92 3.86 2.00 5.68 3.61 2.12 91.69 91.69
-4: VA-AD 20.01 20.01 20.00 16.57 15.77 15.37 91.55 91.59
-*5: AA-TD 3.26 2.07 1.18 13.19 11.82 10.69 91.56 91.46
-*6: TA-AD 18.44 17.61 16.07 16.48 16.13 15.69 91.38 91.62
-*7: AA-AD 6.93 6.39 5.73 9.77 8.74 7.64 91.68 91.58
-*8: AA-AD 19.29 18.63 18.06 16.00 15.74 14.56 91.61 91.39
-0: VA-ND 8.67 8.37 8.09 3.68 1.79 0.76 87.76 87.76
-1: AA-ND 1.80 1.14 0.77 3.73 1.83 0.74 87.76 87.76
-2: VA-VD 15.33 15.33 15.28 11.65 11.42 11.18 87.72 87.73
-3: AA-VD 6.30 4.76 2.63 5.66 3.36 1.94 87.74 87.74
-4: VA-AD 20.01 20.01 20.00 16.79 16.18 15.77 87.61 87.66
-*5: AA-TD 4.53 3.31 2.16 13.59 12.65 11.39 87.58 87.52
-*6: TA-AD 18.32 17.66 16.27 16.60 16.26 15.99 87.50 87.68
-*7: AA-AD 7.12 6.69 6.00 10.67 9.89 8.92 87.74 87.62
-*8: AA-AD 19.41 18.70 18.11 16.75 16.54 15.54 87.66 87.42
+1K 2K 5K ASR 1K 2K 5K ASR BAGS HSJA
+*0: VA-ND 8.27 7.86 7.26 5% 3.42 1.43 0.41 100% 91.69 91.69
+*1: AA-ND 1.26 0.71 0.49 100% 3.14 1.31 0.39 100% 91.69 91.69
+*2: VA-VD 15.27 15.26 15.20 0% 11.14 10.81 10.33 7% 91.68 91.68
+*3: AA-VD 2.63 2.03 1.77 93% 5.68 3.61 2.12 85% 91.69 91.69
+*4: VA-AD 20.01 20.01 20.00 0% 17.17 16.35 15.56 0% 91.60 91.50
+*5: AA-TD 6.28 5.45 4.52 30% 13.19 11.82 10.69 2% 91.52 91.46
+*6: TA-AD 19.52 19.40 18.95 0% 16.48 16.13 15.69 0% 91.38 91.62
+*7: AA-AD 9.95 9.80 9.80 5% 10.30 9.04 7.55 23% 91.66 91.55
+*8: AA-AD 19.85 19.85 19.85 0% 14.46 13.93 13.08 1% 91.69 91.37
+*0: VA-ND 8.72 8.42 7.94 4% 3.73 1.74 0.75 100% 87.76 87.76
+*1: AA-ND 1.74 1.13 0.79 100% 3.64 1.77 0.73 100% 87.76 87.76
+*2: VA-VD 15.42 15.35 15.20 0% 11.10 10.73 10.38 4% 87.72 87.73
+*3: AA-VD 2.82 2.26 2.06 81% 5.66 3.36 1.94 86% 87.74 87.74
+*4: VA-AD 20.01 20.01 20.00 0% 17.06 16.40 15.81 0% 87.66 87.66
+*5: AA-TD 8.48 7.68 6.82 9% 13.59 12.65 11.39 1% 87.58 87.52
+*6: TA-AD 19.58 19.40 18.95 0% 16.60 16.26 15.99 0% 87.50 87.68
+*7: AA-AD 10.43 10.24 10.17 1% 10.21 9.22 7.82 12% 87.73 87.61
+*8: AA-AD 19.86 19.86 19.86 0% 15.71 15.35 14.30 1% 87.67 87.40
   1. 0.
@@
   2. 1.
-    AA-ND – Adaptive Attack / No Defense: The degree at which each attack methodology can be optimized by being adaptive compared to the baseline.
+    AA-ND – Adaptive Attack / No Defense: How more optimal is the adaptive version of an attack compared to the baseline.
   3. 2.
-    VA-VD – Vanilla Attack / Vanilla Defense: The performance of our active defense, but the non-adaptive version operating with an empirically defined detection threshold.
+    VA-VD – Vanilla Attack / Vanilla Defense: The performance of our active defense, the non-adaptive version that has an empirically defined detection threshold.
   4. 3.
@@
   9. 8.
-    AA-AD – Adaptive Attack / Adaptive Defense: Same scenario as before, but it reports the best-case for the defense instead.
+    AA-AD – Adaptive Attack / Adaptive Defense: The exact setup as scenario 7, but the best-case for the defense is reported instead.
 For each successive scenario we evaluate with the most successful policy found, as this is best practice in Markov Games: the worst case opponent policy is fixed and then a counter to it is learned [31, 38].
-Fixing other policies when computing a best response is representative of learning in cybersecurity environments that can contain a range of agents, with the additional benefit of converting the problem to single-agent that, as detailed in Section III-C, one can solve with RL.
+Fixing other policies when computing a best response is representative of learning in environments that can contain multiple agents, with the additional benefit of converting the problem to single-agent that, as detailed in Section III-C, one can solve with RL.
+Comparison to SOTA.
+In Scenarios 0-8 we evaluate all possible combinations between our attack and our defense.
+As a baseline to compare to, we additionally evaluate our approach to the state-of-the-art stateful defenses and adaptive attacks, that is Blacklight [23] and OARS [24] respectively.
+We implement both Blacklight and OARS in our interactive environments by using their publicly available code and parameters.
+As our environments do not return a rejection signal and to make a fair comparison, for OARS rejection coincides with a non-adversarial decision.
+We define 5 further scenarios:
+  1. 9.
+    VA-BD – Vanilla Attack / Blacklight Defense: Baseline performance of the attacks against Blacklight.
+  2. 10.
+    OA-BD – OARS Attack / Blacklight Defense: OARS against Blacklight.
+  3. 11.
+    AA-BD – Adaptive Attack / Blacklight Defense: Our adaptive attack against Blacklight.
+  4. 12.
+    OA-TD – OARS Attack / Trained Defense: OARS against our trained defense from Scenario 6.
+  5. 13.
+    OA-AD – OARS Attack / Adaptive Defense: Our adaptive defense retrained against OARS.
+Our experiments were run on multiple machines, however to give an idea for the time complexity of our defense, on an Intel i7-7700 CPU one forward pass in CIFAR – that is one response to one query – takes $8\pm 1.4$ ms for 700 MFLOPs.
 V-B Results
-For consistency and comparability between evaluations, the listed results are from the same 100 test episodes.
-The gap value denotes the $\ell_{2}$ perturbation that initially separates the starting and best adversarial example, averaged over the 100 episodes.
-A perfect defense would therefore acquire a value very close to the gap, while a perfect attack would never be able to reduce the perturbation to 00; it can be evaluated, however, relative to the baseline or other attacks.
-We tested the trained agents on budgets higher than 5K and we discovered that the depicted trends hold; however, to make the agent training tractable and the evaluation wider, we limit the maximum query budget per adversarial episode to 5K.
-Tables I & II report the results for CIFAR10 and MNIST respectively.
-The color gradient ranges from the gap value to the minimum value that can be achieved @ 5K queries.
-Overall, the empirical results help us extract and highlight several important insights, practical observations, general implications, and crucial takeaways for the broader AML field:
-TABLE II: Mean $l_{2}$ perturbation for 1K, 2K, and 5K queries and accuracy on clean data for MNIST. The evaluation scenarios are identical to Table I.
+For consistency and comparability between evaluations, all results are from the same 100 test episodes.
+The gap value denotes the $L_{2}$ perturbation that initially separates the starting and best adversarial example, averaged over the 100 episodes.
+By testing the trained agents on budgets higher than 5K we discovered that the trend in reducing $L_{2}$ holds; to make the agent training tractable and the evaluation wider however, we limit the maximum query budget per adversarial episode to 5K.
+Tables II & III report the results for CIFAR10, while IV reports MNIST.
+Overall, the empirical results help us extract and highlight several important insights, practical observations, and general implications for the broader AML field:
+TABLE III: ASR and mean $L_{2}$ perturbation for CIFAR-10, comparing our Adaptive Attack (AA) and Adaptive Defense (AD) to Blacklight (BD) and OARS (OA).
+Adv. Trained Scenario CIFAR-10 Gap: 20.01
+BAGS HSJA Clean Acc.
+1K 2K 5K ASR 1K 2K 5K ASR BAGS HSJA
+*9: VA-BD 9.55 9.32 9.17 0% 8.41 8.19 7.80 15% 91.71 91.71
+10: OA-BD 9.46 9.46 9.46 1% 6.54 5.83 4.67 50% 91.71 91.71
+11: AA-BD 2.26 1.39 1.32 98% 4.55 3.08 2.44 78% 91.71 91.71
+12: OA-TD 20.01 20.01 20.01 0% 7.07 6.38 5.53 50% 91.61 91.59
+13: OA-AD 20.01 20.01 20.01 0% 11.03 11.00 10.95 5% 91.61 91.69
+*9: VA-BD 9.75 9.56 9.46 0% 8.67 8.50 8.28 7% 87.76 87.76
+10: OA-BD 9.79 9.79 9.79 1% 5.77 4.53 3.26 72% 87.76 87.76
+11: AA-BD 5.59 4.04 2.55 79% 5.59 4.04 2.55 79% 87.76 87.76
+12: OA-TD 20.01 20.01 20.01 0% 6.44 5.49 4.38 65% 87.66 87.64
+13: OA-AD 20.01 20.01 20.01 0% 11.31 11.12 10.97 7% 87.66 87.74
+  • When comparing the upper and lower halves of each table, we can observe that adversarial training adds a limited amount of robustness; otherwise, the practical effect of adversarial training is a tax on the attacker, forcing them to expend more queries for the same perturbation or having higher perturbation for the same query budget.
+  • Our adaptive defense (AD) outperforms both Blacklight (BD) and non-adaptive stateful (VD), also when transferred (Scen.
+12).
+In HSJA, it reduces ASR by $\sim$90when trained against OA specifically, while it offers similar protection to BD when transferred from another attack.
+  • Even against the strongest attacker and for the worst case for the defense (Scen. 7), our AD achieves an ASR $\leq 23\%$.
+  • Our adaptive attack (AA) outperforms OARS (Scen.
+11) and vanilla attacks (VA) by a wide margin, without access to rejection sampling.
+and irrespective of the defense it faces; the one exception is our adaptive defense (AD), against which it has very limited success.
+  • Evasive transformations interfere with the attack operation, exemplified by the difference between BAGS and HSJA in Scen.
+5; for the attack to reach its full potential, they have to be adaptively controlled together.
+  • The initial performance of an attack can be misleading: out-of-the-box HSJA appears to be the better attack, but it is often outperformed by adaptive BAGS, especially in CIFAR and against active defenses.
+  • In Scen.
+1-8, where agents train against the best opponent policy as previously discovered, ASR oscillates as following a fixed policy enables the learning of an optimal counter to it; and eventually plateaus, indicating an equilibrium.
+  • The advantage of AA is much more pronounced against active defenses, where they significantly outperform non-adaptive versions.
+  • The performance of both attacks deteriorates considerably against active defenses, however the defenses reach their full potential only when also adaptive.
+  • The first time active defenses effectively resist adaptive attacks is in Scen. 5 of CIFAR; we employed evasive transformations from then onward.
+  • Different attack fundamentals respond differently to defenses; the gradient estimation stage of HSJA is naturally disadvantaged against similarity detection, while the jump and binary stages are advantaged.
+  • When devising the adaptive defense for HSJA, it proved nearly impossible to engineer a state the agent can learn on by leveraging our knowledge of the attack and its geometric functioning.
+What did prove effective, however, was pure computation [^2] This reminds us of Sutton’s Bitter Lesson [50], the observation that progress in AI is often driven by gains in computation rather than problem-specific expert knowledge.: we learn an embedding with Contrastive Learning [51] from raw input to use a state, which transfers exceedingly well to other attacks like BAGS.
+TABLE IV: ASR and mean $L_{2}$ perturbation for 1K, 2K, and 5K queries and accuracy on clean data for MNIST. The evaluation scenarios are identical to Tables II and III.
 Adv. Trained Scenario MNIST Gap = 10.62
 BAGS HSJA Clean Acc.
-1K 2K 5K 1K 2K 5K BAGS HSJA
-0: VA-ND 5.28 5.26 5.24 3.63 3.06 2.61 99.37 99.37
-1: AA-ND 2.75 2.57 2.47 3.61 3.04 2.58 99.37 99.37
-2: VA-VD 8.74 8.28 7.94 5.74 5.73 5.73 99.35 99.20
-3: AA-VD 3.39 3.29 3.19 3.58 3.12 2.74 99.33 99.31
-4: VA-AD 10.45 10.45 10.45 9.87 9.87 9.87 99.31 99.30
-5: AA-TD 3.54 3.27 3.10 4.19 3.54 2.80 99.32 99.20
-6: TA-AD 9.85 9.77 9.76 9.99 9.99 9.99 99.28 99.34
-7: AA-AD 4.95 4.92 4.92 7.07 6.97 6.76 99.31 99.36
-8: AA-AD 10.62 10.62 10.62 8.78 8.71 8.66 99.32 99.36
-0: VA-ND 5.28 5.27 5.25 4.65 4.02 3.40 99.15 99.15
-1: AA-ND 3.28 3.07 2.94 4.60 3.97 3.35 99.15 99.15
-2: VA-VD 8.67 8.44 8.08 5.71 5.70 5.70 99.14 98.89
-3 AA-VD 4.14 4.04 3.81 4.69 4.35 3.87 99.13 99.15
-4: VA-AD 10.39 10.39 10.39 10.07 10.07 10.07 99.09 99.12
-5: AA-TD 4.01 3.80 3.61 4.94 4.44 3.89 99.11 99.12
-6: TA-AD 10.05 9.97 9.95 10.20 10.20 10.20 99.07 99.06
-7: AA-AD 5.51 5.50 5.50 5.47 5.16 4.99 99.01 99.13
-8: AA-AD 10.62 10.60 10.60 10.12 10.12 10.12 99.10 99.01
-  • A pervasive observation when comparing the upper and lower halves of each table is that adversarial training adds a small amount of robustness; otherwise, the practical effect of adversarial training is a tax on the attacker, forcing them to expend more queries for the same perturbation or having higher perturbation for the same query budget.
-  • Evasive transformations interfere with the attack operation, and only when adaptively controlled together does the attack reach its full potential, as demonstrated by the difference between BAGS and HSJA in CIFAR scenarios 5-8.
-  • Adaptive defenses outperform non-adaptive or model hardening, as even against the strongest attacker and the worst case for the defense – the 7th scenario – the perturbation for 5K queries is considerably higher.
-  • The initial performance of an attack can be misleading: at first glance HSJA appears to be the better one but it is often outperformed by adaptive BAGS, especially in CIFAR-10 and against active defenses.
-  • Not surprisingly, adaptive versions of the vanilla attacks perform better (Scenarios 1 & 0); especially BAGS, as is not well optimized out-of-the-box.
-  • The advantage of adaptive attacks however is much more pronounced against active defenses, where they significantly outperform non-adaptive versions.
-  • The performance of both attacks deteriorates considerably against active defenses, however they reach their full potential only when they are also adaptive.
-  • Different attack fundamentals respond differently to active defenses; the gradient-based HSJA is naturally disadvantaged against similarity detection, especially on higher-dimensional data like CIFAR-10.
-  • We observe that between consecutive scenarios, where an agent learns against the best opponent policy as previously discovered, the $l_{2}$ values oscillate: following a fixed policy enables the learning of an optimal counter to it.
-  • When devising an adaptive defense for HSJA, it proved nearly impossible to engineer a state the agent can learn on by leveraging our knowledge of the attack and its geometric properties.
-What did prove effective, however, was pure computation [^1] This reminds us of Sutton’s Bitter Lesson [47]: the observation that progress is often driven by gains in computation rather than problem-specific expert knowledge.: to learn a state embedding from raw input, based on Contrastive Learning.
-These state embeddings also transfer exceedingly well to different attacks like BAGS.
-  • The first time an active defense withstands adaptive attacks is in the 5th CIFAR-10 scenario, for BAGS and HSJA both.
-This is also the first scenario we had to employ evasive transformations to evade the active defense; subsequently, scenarios 5-8 in CIFAR-10 are the only with evasive transformations.
-  • Evasive transformations combine differently with different attacks.
-All the offensive scenarios in BAGS recoup their performance almost completely.
-This is not the case for HSJA as the latter requires the preservation of the correct gradient information, something that had largely unsatisfactory solutions even with full white-box access to the model and to the true gradient.
-V-C Empirical vs. Adaptive Evasion
-So far our evaluations have focused on optimizing attack performance and evasiveness jointly.
-We are also interested in highlighting and measuring the potential of adaptively controlling the evasive actions without controlling the attack parameters themselves, against our proposed defensive agents.
-To compare how effective are adaptive evasive actions to the empirically defined ones, we evaluate them along with the approach of Chen et al.
-[16], configured as described in their work.
-we keep everything else exactly the same and the attacks themselves use their default, non-adaptive parameters: we want to isolate the effect of evasive actions alone.
-Table III compares the results between the empirical (yellow) and adaptive control (red) settings, and as we can observe our approach consistently outperforms the empirical definition of evasive transformations.
-Additionally, we note that the evasive agents were trained only in the upper half of the table, that is against non adversarially trained models: notably, adaptively controlling evasive actions transfers to other models too.
-TABLE III: Performance comparison (in $l_{2}$ perturbation) between empirical input transformations (yellow scenarios) and the adaptive control of the same transformations (red scenarios) on CIFAR-10.
-Bottom half of the table is against adversarially trained models.
-Scenario CIFAR-10 Gap: 20.01
-BAGS HSJA
-1K 2K 5K 1K 2K 5K
-BA-VD 8.66 8.44 8.24 10.55 9.93 9.29
-CA-VD 8.18 7.92 7.63 8.14 6.76 5.04
-BA-TD 8.96 8.70 8.44 11.62 10.85 10.09
-CA-TD 8.62 8.26 7.95 11.06 10.46 9.78
-BA-AD 16.78 16.07 14.87 10.58 9.88 9.16
-CA-AD 9.72 9.30 8.91 10.22 9.73 9.10
-BA-VD 8.47 8.26 8.05 10.21 9.76 9.33
-CA-VD 7.53 7.11 6.72 9.87 9.10 7.81
-BA-TD 8.40 8.21 7.95 11.58 10.88 10.13
-CA-TD 7.44 6.78 6.14 11.62 10.70 10.02
-BA-AD 16.97 16.13 15.02 10.67 10.00 9.43
-CA-AD 9.34 8.66 8.34 10.08 9.62 9.07
-V-D Base Rate of Attacks
-In all the evaluations so far we use a fixed probability $P(adv)=0.5$ that an incoming query is adversarial.
-To assess how the active and adaptive defenses (Scenarios 4 & 6) perform in the complete absence of attacks, we evaluate them with $P(adv)=0$ without any retraining; the results are shown in Table IV.
-We observe a small reduction in the accuracy on clean samples that can be attributed to the considerably different base rate of adversarial and benign queries.
-Note however that as the probability of each is an intrinsic property of each environment, if the base rate of attacks changes any defensive agent can be retrained to adjust to it.
-TABLE IV: Clean accuracy on CIFAR-10 for scenarios 4 & 6, with standard and zero probability of attack. Yellow denotes normally trained models and red adversarially trained.
-P(adv) BAGS4 BAGS6 HSJA4 HSJA6
-0.5 91.55 91.38 91.59 91.62
-0.0 90.91 90.95 90.48 90.86
-0.5 87.61 87.50 87.58 87.68
-0.0 87.02 87.11 86.70 86.98
+1K 2K 5K ASR 1K 2K 5K ASR BAGS HSJA
+*0: VA-ND 5.30 5.28 5.26 3% 3.59 3.07 2.61 73% 99.37 99.37
+*1: AA-ND 2.74 2.57 2.47 78% 3.61 3.09 2.60 74% 99.37 99.37
+*2: VA-VD 7.44 6.66 5.63 22% 5.82 5.78 5.73 2% 99.34 99.20
+*3: AA-VD 3.79 3.66 3.44 29% 3.54 3.09 2.77 61% 99.37 99.31
+*4: VA-AD 10.57 10.57 10.57 0% 10.05 10.05 10.05 0% 99.31 99.30
+*5: AA-TD 3.57 3.29 3.14 39% 5.00 3.97 3.38 36% 99.32 98.84
+*6: TA-AD 10.62 10.62 10.62 0% 10.23 10.23 10.18 0% 99.28 99.34
+*7: AA-AD 4.89 4.89 4.86 8% 5.06 4.76 4.38 36% 99.31 99.35
+*8: AA-AD 10.62 10.62 10.62 0% 10.21 10.21 10.21 0% 99.32 99.23
+*9: VA-BD 10.62 10.62 10.62 0% 5.65 5.65 5.65 2% 99.37 99.37
+10: OA-BD 10.62 10.62 10.62 0% 4.53 4.00 3.15 46% 99.37 99.37
+11: AA-BD 3.83 3.69 3.60 17% 4.18 3.66 3.19 52% 99.37 99.37
+12: OA-TD 10.62 10.62 10.62 0% 10.21 10.20 10.20 0% 99.22 99.28
+13: OA-AD 10.62 10.62 10.62 0% 10.21 10.20 10.20 0% 99.32 99.28
+*0: VA-ND 5.26 5.25 5.24 2% 4.61 4.04 3.41 30% 99.15 99.15
+*1: AA-ND 3.28 3.08 2.96 51% 4.59 3.97 3.35 34% 99.15 99.15
+*2: VA-VD 7.70 6.86 5.86 17% 5.81 5.78 5.76 2% 99.14 99.12
+*3 AA-VD 4.18 4.08 3.86 22% 4.63 4.27 3.86 25% 99.13 99.15
+*4: VA-AD 10.55 10.55 10.55 0% 10.02 10.02 10.02 0% 99.09 99.08
+*5: AA-TD 4.04 3.74 3.54 27% 5.82 5.09 4.26 16% 99.11 98.78
+*6: TA-AD 10.62 10.62 10.62 0% 10.20 10.20 10.20 0% 99.06 99.06
+*7: AA-AD 5.59 5.56 5.56 5% 5.47 5.16 4.99 14% 99.09 99.13
+*8: AA-AD 10.62 10.62 10.62 0% 10.12 10.12 10.12 0% 99.10 99.01
+*9: VA-BD 10.62 10.62 10.62 0% 5.65 5.64 5.64 1% 99.15 99.15
+10: OA-BD 10.62 10.62 10.62 0% 5.18 4.80 4.11 17% 99.15 99.15
+11: AA-BD 4.31 4.07 3.96 13% 5.04 4.65 4.20 19% 99.15 99.15
+12: OA-TD 10.62 10.62 10.62 0% 10.26 10.26 10.26 0% 99.00 99.06
+13: OA-AD 10.62 10.62 10.62 0% 10.26 10.26 10.26 0% 99.10 99.06
 VI Discussion
-With AI-enabled decision-making becoming more pervasive and accessible, in domains like governance, finance, employment, and of course cybersecurity, more and more decisions are delegated to AI which becomes increasingly accountable for upholding safety and ethical constraints.
-As trustworthy AI is vital for the healthy functioning of whole ecosystems, this study highlights several security risks but also potential mitigations in these inherently black-box environments.
-Our work has several implications for performing robust inference in the wild.
-While adversarial training remains the most reliable defense, the amount of robustness it imparts will vary and even be insufficient.
+Our work has several implications for performing robust inference in the real-world.
+While adversarial training remains the most reliable defense, the empirical robustness it imparts will vary and even be insufficient.
+We note that this robustness is against all adversarial examples under the same $L_{p}$-norm; active defenses protect only against querying attacks, but as they do transfer between attacks (cf.
+Scen.
+12) they can be used jointly as complementary approaches.
 We demonstrated how AI-enabled systems are susceptible to adaptive adversaries that devise new evasive techniques and control them jointly with other attack parameters.
 This has been achieved in the fully black-box case and against active defenses.
-Even more concerning is the level of threat that such systems face from AI-enabled adversaries, as it is straightforward to generalize Theorem III.1 to any other domain or modality.
-This rekindles the proverbial arms race, where as a consequence defenses should also be backed by equally capable AI.
-Finally, our study highlights some aspects of the nature of adversarial examples.
-In black-box interactive environments they can be generalized by the notion of adversarial behavior, something that is more evident in domains where visual affinity is irrelevant, e.g., malware detection.
-Adversarial examples (and behavior) bring essence into question, what makes something, something; thus they can be approached from various epistemic perspectives.
-Computational approaches concern themselves with how to effectively acquire them; while their inherently competitive nature is already reflected in seminal works like shortcut learning [23] and invariant risk minimization [1].
-Even in phenomenology – the philosophical study of the structures of experience and consciousness – the eidetic variation technique involves imagining and varying the features of an object under investigation to discover the (in)essential ones [31].
-What these disparate approaches bring to light is that towards more general and robust AI researchers will have to reflect on the causal structure of perceptions and preconceptions, as well as on ways to automate this process.
+Notably, the level of threat that adaptive adversaries pose against such systems is considerable, as it is straightforward to generalize Proposition III.3 to any other domain or modality.
+This rekindles the proverbial arms race, where as a consequence defenses should also be equally capable and adaptive.
 Limitations.
-We note that while adversarial training has limitations, the robustness it provides is against all kinds of adversarial examples (under the same $\ell_{p}$-norm), while active defenses protect primarily against attacks they were trained with.
-As complementary approaches however, they can be used jointly.
-To keep the amount of evaluations practical, we narrowed the scope to targeted attacks and to $\ell_{2}$ as the more suitable norm for visual similarity.
-Targeted attacks are strictly more difficult to perform than untargeted, while in binary classification targeted and untargeted coincide; the latter is also the prevalent mode in cybersecurity contexts.
-Our framework, however, can accommodate any adversarial goal or metric.
-Another simplifying assumption we make is that only one attack can take place at a time; however, the queuing methodology we use for incoming queries is readily extensible to handle concurrent attacks.
-Other possibilities to consider is how well an active defense handles attack methodologies it was not trained on, and to extend its training to multiple ones.
-Finally, in the evaluation we focus on the breadth of scenarios and environments where agents learn concurrently, thus the number of datasets we experiment with is limited.
-However, we back our empirical study with an extensive theoretical analysis that gives support to our findings generalizing to any dataset or application domain.
+To keep the amount of evaluations practical, we narrowed the scope to targeted attacks and to $L_{2}$ as the more suitable norm for visual similarity.
+Targeted attacks are strictly more difficult to perform than untargeted, while for binary classification targeted and untargeted coincide; our framework, however, can accommodate any adversarial goal or metric.
+Another simplifying assumption we make is that only one attack can take place at a time; however, the queuing technique we use for incoming queries is readily extensible to handle concurrent attacks.
+While we demonstrate that our active defense does transfer between attacks, another possibility to explore is training the defense on multiple kinds of attacks.
+Finally, in our evaluation we focus on a wide range of adaptive and non-adaptive scenarios where agents learn and adapt interactively, thus limiting the number of datasets we experiment with; we back our empirical study however with an extensive theoretical analysis that supports the generality of our findings independent of context.
 Future Work.
 The AMG framework we introduce is general by design and can accommodate the learning of optimal offensive and defensive policies in any domain of interest beyond image classification.
-A promising path for future research is the extension of adaptive attacks and defenses to other domains and modalities, e.g.
-tasks like malware, bot, and network intrusion detection, precisely because our approach circumvents the obstacle of computing and mapping gradients to feasible perturbations and operates instead directly on the problem space [39].
-Another compelling but formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing instruments to bypass defenses and thus imparting controllability to adversarial tasks.
-Finally, in our work we considered opponent agency as part of the environment; other domains, like malware analysis, might benefit from explicit opponent modeling.
+A promising path for future research is the extension of our adaptive attacks and defenses to other domains and modalities, for instance malware, bot, and network intrusion detection.
+This is specifically because our approach circumvents the main obstacle of mapping gradient-based perturbations to feasible objects (eg.
+binaries) and instead can function directly in the problem space [36].
+Another compelling and formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing instruments to bypass defenses and thus imparting controllability to adversarial tasks.
+Finally, in our work we considered opponent agency as part of the environment; other domains, like malware detection, might benefit from explicit opponent modeling.
 VII Conclusion
-With decision-based attacks becoming more pervasive in multiple domains, every AI-based decision-making process that exposes a queryable interface is inherently vulnerable.
+With adaptive, decision-based attacks becoming more pervasive in multiple domains, every AI-based decision-making process that exposes a queryable interface is inherently vulnerable.
 To aggravate matters, this vulnerability cannot be mitigated by employing model hardening approaches like adversarial training alone.
@@
 However, the existence of such defenses elicits in turn adaptive attacks which are able to recover part of their original performance.
-We perform a theoretical and empirical investigation of decision-based attacks and defenses under a unified framework we name “Adversarial Markov Games” (AMG).
+We perform a theoretical and empirical investigation of decision-based attacks and stateful defenses under a unified framework we name “Adversarial Markov Games” (AMG).
 In self-adaptive, we introduce a novel twofold definition of adaptive: both devising new methods of outmaneuvering opponents and adapting one’s operating policy with respect to other agency in the environment.
-Furthermore, our adversarial policy gradient theorem indicates that any combination of adversarial goals, be it performance, stealthiness, disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any modality.
-As new attacks and defenses appear (and get broken) regularly, our evaluation methodology is applicable to any within the outlined scope by transforming the existing arms-race through their self-adaptive versions, thus ensuring accurate and robust assessment of their performance.
-The AMG framework we introduce helps us reason on and assess the vulnerabilities of AI-based systems by disentangling the inherently complex and non-stationary task of learning in the presence of competing agency.
-We can model it as part of the environment, and thus simplify the task by computing a best response against the observed behavior.
-This is an important outcome for cybersecurity domains: as long as proper threat modeling is carried out, one can readily employ RL algorithms in order to devise optimal defenses; but only after they devised optimal attacks too.
-w
+Furthermore, our theoretical analysis highlights that any combination of adversarial goals, be it performance, stealthiness [52], or disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any modality.
+As new attacks and defenses appear and get broken regularly, our evaluation methodology is generally applicable within the outlined scope by turning any approaches in the existing arms-race self-adaptive and thus ensuring accurate and robust assessment of their performance.
+The AMG framework we introduce helps us reason on and properly assess the vulnerabilities of AI-based systems, disentangling the inherently complex and non-stationary task of learning in the presence of competing agency.
+By modeling the latter as part of the environment, we can simplify this task by computing a best response to the observed behavior.
+This is a significant implication for the security of AI-based systems: as long as proper threat modeling is carried out, one can readily employ RL algorithms in order to devise optimal defenses, but only after they devised optimal attacks too.
 Appendix A Proofs
-For a more intuitive understanding of the proofs, we first provide a high-level description of the attack fundamentals.
+For a more intuitive understanding of the proofs, we provide a high-level description of the attack fundamentals.
 BAGS [29] performs a random walk along the boundary between the adversarial and the non-adversarial regions, by first taking a random step orthogonal to the original image direction, then a source step towards it.
@@
 $=a_{t}x_{c}+b_{t}\Biggl{\{}x_{t}+\xi_{t}\frac{\frac{1}{B}\sum_{b=1}^{B}\mathcal{I}(C(x_{t}+\delta u_{b}),c_{0})u_{b}}{\|\frac{1}{B}\sum_{b=1}^{B}\mathcal{I}(C(x_{t}+\delta u_{b}),c_{0})u_{b}\|_{2}}\Biggl{\}}$
-Recall that we assumed $C(x)=\operatorname*{arg\,max}F_{c}(x)$, then the updates in Eq.
-(16) are guaranteed by Theorem 1 of HSJA [15] to converge to a stationary point $x_{b}$ of Eq.
-(5).
-Given a standard threshold $\epsilon$ on imperceptibility and over $N$ adversarial episodes, this implies that $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$.
-As $C$ is the only term in the expression (16) that is affected by the model, we reach a contradiction and that an alternative classifier $C^{\prime}$ is required, with $C(x)=D(F_{c}(x))$.
-As the discriminant function $F_{c}$ cannot change without retraining, it follows that $D^{\prime}\neq\operatorname*{arg\,max}$, so that for the adversarial example $x_{t}$ misclassified as $c_{0}$, Eq.
+In Eq.
+(16), the iterates $x_{t}$ are guaranteed by Theorem 1 of HSJA [7] to converge to a stationary point $x_{b}$ of Eq.
+(5), that is $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$, for $\epsilon$ a standard imperceptibility threshold and $N$ adversarial episodes.
+We thus reach a contradiction, and as in Eq.
+(16) the only model-dependent term is $C(\cdot)$, an alternative classifier $C^{\prime}$ is required.
+With $C(x)=D(F_{c}(x))$, and the discriminant function $F_{c}$ unable to change without retraining, it follows that $D^{\prime}\neq\operatorname*{arg\,max}$, so for adv.
+example $x_{t}$ misclassified as $c_{0}$, Eq.
 4 can return -1:
@@
 while $x_{n}=x_{t}\implies D(\tau,F_{c}(x_{n}))\neq D(\tau,F_{c}(x_{t}))$.
-∎
-[Theorem III.1]
-Proof.
-Let $H$ be a distance-based detector that based on a stateful representation $\mathcal{S}(\tau,x_{t})$ decides if a query $x_{t}$ is adversarial or not, and we assume the candidate generation policy $\pi_{\theta}^{\mathcal{A}}$ can influence this representation through intrinsic parameters like source step size, or applying transformations the underlying model is invariant to (V).
-For a predefined threshold $\delta$, if $\mathcal{S}(\tau,x_{t})<\delta$ then $H(x_{t})=1$ indicating that $x_{t}$ is part of an adversarial attack, otherwise $H(x_{t})=0$.
-Now for a number $N$ of attack episodes of fixed length $L$ and a reward function $r(\tau)=\sum_{t=1}^{L}(1-H(x_{t}))$, the objective to be maximized is the expected reward of the adversarial policy:
-$\mathcal{J}(\theta)=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)]=\frac{1}{N}\sum_{i=1}^{N}r(\tau)=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-H(x_{t}))$ (18)
-As it is not possible to compute the derivative of an expectation, by using the Policy Gradient Theorem [48] we can transform it to the expectation of the product of the reward and the gradient of the logarithm of the policy:
-$\nabla\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)]$ $=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau)\nabla\log\pi_{\theta}(\tau)]$ (19)
-$=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[\sum_{t=1}^{L}(1-H(x_{t}))\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)]$
-$=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-H(x_{t}))\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$
-To maximize this expectation empirically we sample $e$ episodes from the environment, compute the gradient of the policy $\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$ and then update the policy with a learning step $a$, $\theta^{\prime}=\theta+\alpha\cdot\sum_{0}^{e}r(\tau)\nabla\log\pi_{\theta}^{\mathcal{A}}(\tau)$.
-After this process has converged through gradient ascent, and as the maximum in Eq.
-18 is attained for $\sum_{i=1}^{N}\sum_{t=1}^{L}H(x_{t})=0$, the obtained policy $\pi_{\theta}^{\mathcal{A}}$ will correspond to the optimally evasive one $\mathcal{E}$.
 ∎
@@
 […]
+[Proposition III.3]
+Proof.
+Let $\pi_{\theta}^{\mathcal{A}}$ be the adversarial policy generating queries $x_{t}$, in $N$ episodes $\tau_{i}$ of length $L$.
+The defense $\pi_{\phi}^{\mathcal{D}}$ (7), upon receiving a query $x_{t}$ outputs a decision $\alpha_{t}=\pi_{\phi}^{\mathcal{D}}(x_{t},s_{t}^{\mathcal{D}})\in\{0,1\}$, with $1$ and 00 indicating rejection and acceptance respectively.
+The goal of the adversary is to find the parameters $\theta^{*}$ that maximize the expected reward:
+$\mathcal{J}(\theta)=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau_{i})]=\frac{1}{N}\sum_{i=1}^{N}r(\tau_{i})=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-\alpha_{t})$ (21)
+The gradient of $\mathcal{J}(\theta)$ with respect to the policy parameters $\theta$ is the direction of steepest ascent for Eq.
+(21).
+Through the Policy Gradient Theorem [53] we can express the gradient of the expected reward in terms of the gradient of the log-likelihood of the policy, weighted by the reward:
+$\nabla_{\theta}\mathcal{J}(\theta)$ $=\nabla_{\theta}\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}[r(\tau_{i})]$ (22)
+$=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}\left[r(\tau_{i})\nabla_{\theta}\log\pi_{\theta}(\tau_{i})\right]$
+$=\mathbb{E}_{\pi_{\theta}^{\mathcal{A}}}\left[\sum_{t=1}^{L}(1-a_{t})\nabla_{\theta}\log\pi_{\theta}^{\mathcal{A}}(\tau_{i})\right]$
+$\approx\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{L}(1-a_{t})\nabla_{\theta}\log\pi_{\theta}^{\mathcal{A}}(\tau_{i})$
+The gradient is thus estimated by sampling $N$ episodes from the policy $\pi_{\theta}^{\mathcal{A}}$ to compute Eq.
+(LABEL:eqn:rew_app).
+To maximize the expectation, we iteratively update the policy parameters $\theta$ using gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}(\theta)$.
+With $\eta>0$ the learning rate, this process converges to $\theta^{*}$.
+Recall that Eq.
+(21) attains its maximum for $\sum_{i=1}^{N}\sum_{t=1}^{L}a_{t}=0$, therefore the converged policy $\pi^{\mathcal{A}}_{\theta^{*}}$ will correspond to the optimal evasive policy.
 Appendix B On States & Rewards
 [States].
-In POMDPs a single observation cannot constitute a Markovian state as typically there is some of dependence on the state history.
-Partial observability is handled by either using recurrent architectures for the policy and value networks, or by engineering a state that includes past information: in this work we opt for the later.
+By merit of operating in black-box environments, the optimization process for both offensive and defensive agents can be described as a POMDP.
+To handle this partial observability, the states we engineer incorporate past information.
 For BAGS, the adversary uses an 8-dimensional state representation with the following information normalized in the range $[0,1]$ : current amount of queries $i$, average queries that are adversarial $a$, the initial gap $g$, the current gap $d$, the location $l=\frac{d}{g}$, the slope $s=m-l$ where $m$ is a moving average of the location, the frequency of improvement $f$, and $r$ which is a moving average of the perturbation reduction $n$, In HSJA the state representation is slightly different: $r=\frac{n}{g}$, and $f=\frac{1}{j}$ with $j$ number of jump steps in last iteration.
-Regarding the defense, for HSJA (and to a lesser extent BAGS) it has been difficult to engineer a representative state for policies to effectively learn on.
+For the defense, in HSJA (and to a lesser extent BAGS) it has been difficult to engineer a state for policies to effectively learn on.
 The knowledge of the attack internals and fundamentals, geometric properties and distances, model activations and logits, and any combination thereof, did not suffice.
-Ultimately we turned to pure computation to learn a state representation for the defensive agent.
-This representation is a 64-dimensional embedding of a CNN trained with triplet loss, on data generated by HSJA and benign queries, where the input is a tensor where the last query is subtracted from the 25 most recent adversarial queries and then stacked together.
-[Rewards].
-The concrete definitions of the rewards for each type of agent are:
+Ultimately we decided to learn a representation instead.
+This representation is a 64-dimensional embedding of a CNN trained with triplet loss, on data generated by HSJA and benign queries, with the input being a tensor of the last query subtracted from the 25 most recent adversarial queries and then stacked.
+[Rewards]. The concrete definitions of the rewards for each type of agent are:
 […]
@@
 Translation -0.2 – 0.2 0 – 1
-Appendix C Models & Hyperparameters
+Appendix C Base Rate of Attacks
+In all the evaluations so far we use a fixed probability $P(adv)=0.5$ that an incoming query is adversarial.
+To assess how our adaptive stateful defenses (Scenarios 4 & 6) perform in the complete absence of attacks, we evaluate them with $P(adv)=0$ without retraining; the results are shown in Table VI.
+We observe a small reduction in the accuracy on clean samples that can be attributed to the considerably different base rate of adversarial and benign queries.
+Note however that as the probability of adversarial queries is an intrinsic property of each environment, if the base rate of attacks changes the defensive agents can be retrained to adjust to it.
+TABLE VI: Clean accuracy on CIFAR-10 for scenarios 4 & 6, where $P(adv)$ denotes the probability that a query is part of an attack.
+Adv. Trained $P(adv)$ BAGS4 BAGS6 HSJA4 HSJA6
+0.5 91.55 91.38 91.59 91.62
+0.0 90.91 90.95 90.48 90.86
+0.5 87.61 87.50 87.58 87.68
+0.0 87.02 87.11 86.70 86.98
+Appendix D Models & Hyperparameters
 The image classification models we use are ResNet-20 for CIFAR-10 and a standard 2 convolutional / 2 fully-connected layer NN for MNIST.
@@
 This network is trained with contrastive loss, where dissimilar examples are generated by adding Gaussian noise and performing evasive transformations on the input from the list in Table V.
-For the PPO agents trained for each scenario, we use the open source library Stable Baselines 3  [^2] https://github.com/DLR-RM/stable-baselines3.
-Policies are parameterized by a two fully-connected layer NN; the hyperparameter search space is shown in Table VI.
-TABLE VI: Hyperparameter space for the PPO agent.
+For the PPO agents trained for each scenario, we use the open source library Stable-Baselines3  [^3] https://github.com/DLR-RM/stable-baselines3.
+Policies are parameterized by a two fully-connected layer NN; the hyperparameter search space is shown in Table VII.
+Blacklight & OARS.
+For evaluating Blacklight [23] and OARS [24] we use their default hyperparameters as those are provided in the publicly available implementations.
+In particular, as OARS spends 200 extra queries per episode to adapt the proposal distribution, we add those on top of b the evaluation budget.
+Additionally, as our defense is not rejection based, we replace the rejection decision with a non-adversarial one.
+TABLE VII: Hyperparameter space for the PPO agent.
 Hyperparameter BAGS HSJA
@@
 batch size 32 – 128 32 – 64
-buffer 2048 – 2048 256 – 2048
+buffer 2048 – 2048 64 – 1024
 epochs 20 – 20 20 – 20

### v3

@@
-Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses
+The Adaptive Arms Race: Redefining Robustness in AI Security
 Abstract
-Despite considerable efforts on making them robust, real-world ML-based systems remain vulnerable to decision based attacks, as definitive proofs of their operational robustness have so far proven intractable.
-The canonical approach in robustness evaluation calls for adaptive attacks, with complete knowledge of the defense and tailored to bypass it.
-In this study, we introduce a more expansive notion of being adaptive and show how attacks but also defenses can benefit by it and by learning from each other through interaction.
-We propose and evaluate a framework for adaptively optimizing black-box attacks and defenses against each other through the competitive game they form.
-To reliably measure robustness, it is important to evaluate against realistic and worst-case attacks.
-We thus augment both attacks and the evasive arsenal at their disposal through adaptive control, and observe that the same can be done for defenses, before we evaluate them first apart and then jointly under a multi-agent perspective.
-We demonstrate that active defenses, which control how the system responds, are a necessary complement to model hardening when facing decision-based attacks; then how these defenses can be circumvented by adaptive attacks, only to finally elicit active and adaptive defenses.
-We validate our observations through a wide theoretical and empirical investigation to confirm that AI-enabled adversaries pose a considerable threat to black-box ML-based systems, rekindling the proverbial arms race where defenses have to be AI-enabled too.
-Our approach outperforms the state-of-the-art black-box attacks and defenses, while bringing them together to render effective insights over the robustness of real-world deployed ML-based systems.
+Despite considerable efforts on making them robust, real-world AI-based systems remain vulnerable to decision based attacks, as definitive proofs of their operational robustness have so far proven intractable.
+Canonical robustness evaluation relies on adaptive attacks, which leverage complete knowledge of the defense and are tailored to bypass it.
+This work broadens the notion of adaptivity, which we employ to enhance both attacks and defenses, showing how they can benefit from mutual learning through interaction.
+We introduce a framework for adaptively optimizing black-box attacks and defenses under the competitive game they form.
+To assess robustness reliably, it is essential to evaluate against realistic and worst-case attacks.
+We thus enhance attacks and their evasive arsenal together using reinforcement learning (RL), apply the same principle to defenses, and evaluate them first independently and then jointly under a multi-agent perspective.
+We find that active defenses, those that dynamically control system responses, are an essential complement to model hardening against decision-based attacks; that these defenses can be circumvented by adaptive attacks, something that elicits defenses being adaptive too.
+Our findings, supported by an extensive theoretical and empirical investigation, confirm that adaptive adversaries pose a serious threat to black-box AI-based systems, rekindling the proverbial arms race.
+Notably, our approach outperforms the state-of-the-art black-box attacks and defenses, while bringing them together to render effective insights into the robustness of real-world deployed ML-based systems.
 I Introduction
-AI models are predominantly trained, validated, and deployed with little regard to their correct functioning under adversarial activity, often leaving safety, ethical, and broader societal impact considerations as an afterthought.
-Adversarial contexts further aggravate the typical generalization challenges that these models face with threats beyond model evasion (extraction, inversion, poisoning [1]) while the systems they enable often expose interfaces that can be queried and used as adversarial “instructors”, like in constructing adversarial malware against existing ML-based malware detection [2, 3].
-Scoping on model evasion, the most reliable mitigation to date is adversarial training [4, 5], an approach not without limitations as these models often remain irreducibly vulnerable at deployment, particularly against black-box, decision-based attacks [6, 7, 8].
+AI models are predominantly trained, validated, and deployed with little regard to their correct functioning under adversarial activity, often leaving safety and security considerations as an afterthought.
+Adversarial contexts further aggravate the typical generalization challenges that these models face with threats beyond model evasion (misclassification), like model extraction, model inversion, and model poisoning [1].
+At the same time, the systems these models are components of often expose interfaces that can be queried and used as adversarial “instructors”, like in constructing adversarial malware against existing AI-based malware detection [2, 3].
+Focusing on adversarial examples for model evasion, the most reliable mitigation to date is adversarial training [4, 5], an approach not without limitations as these models often remain irreducibly vulnerable at deployment, particularly against black-box, decision-based attacks [6, 7, 8].
 Nevertheless, all such attacks exhibit a behavior at-the-interface that can be described as adversarial itself, a generalization that subsumes adversarial examples and opens a path towards novel defenses and mitigations.
 Adversarial behavior is a temporal extension of adversarial examples, perhaps not malicious or harmful in isolation, yet part of an attack as it unfolds over time; it is also the canonical description of adversarial examples in domains like dynamic malware analysis and adversarial RL [9, 10].
-Aside from making the underlying models more robust, this behavior can be countered as such rather than relying on hardened models exclusively.
-As models cannot update their decision boundary in an online manner and in response to adversarial activity on their interface, there has to be a complement to model hardening: for instance active defenses such as rejection or misdirection [11, 12, 13].
-In this study we address a crucial gap, as in adversarial machine learning (AML) evaluating the robustness of defenses against oblivious, non-adaptive, and therefore suboptimal attackers is inherently problematic [14, 15].
-We expand the conventional notion of adaptive, from adapted attacks that have an empirical configuration to bypass the defense, to include the capability to self-adapt, where attacks adaptively control their parameters and evasive actions together in response to how the model under attack and its defenses respond [16].
-We demonstrate theoretically and empirically how self-adaptive attacks can use reinforcement learning (RL) to modify their policies to become both optimal and evade active detection.
-Notably, this can be performed in a gradient-based manner even in fully black-box contexts [III.3], and is a capability that properly reflects the level of adversarial threat and in that way does not overestimate the empirical robustness; real attackers might compute gradients after all.
-This capability, however, enables active defenses in turn.
-Through rigorous threat modeling and by simulating self-adaptive attackers, their full potential is uncovered and effective counter-policies can be learned.
-To express the call for adaptive evaluations in AML differently, a defense is trustworthy only insofar the adversary is optimal.
-This mutual interdependence creates the imperative for both attacks and defenses to be self-adaptive, thereby establishing the competitive, zero-sum game that their interaction forms.
-Our research approaches the question of robustness from these two closely related perspectives: first, how to optimize decision-based attacks, and secondly what defensive options exist.
-In this manner, we explore both offensive and defensive strategies in depth, resulting in the following key contributions:
+Aside from making the underlying models more robust, this behavior can be countered as such, rather than relying on hardened models exclusively.
+As AI models cannot update their decision boundary in an online manner and in response to adversarial activity on their interface, there has to be a complement to model hardening: for instance active defenses such as rejection or misdirection [11, 12, 13].
+In this study we identify and address a crucial gap: evaluating the robustness of defenses against oblivious, non-adaptive, and therefore suboptimal attackers, renders any results unreliable [14, 15].
+The key observation we make is that robustness must account for the ability of the adversary to adapt while interacting with the model.
+To that end, we expand the conventional notion of adaptive, from adapted attacks that have an empirical configuration to bypass the defense, to include the capability to self-adapt, where attacks adapt their parameters and evasive actions together, based on how the target model and its defenses respond [16].
+We demonstrate theoretically and empirically how self-adaptive attacks can use RL to modify their policies to become both optimal and evade active detection.
+Notably, this can be performed in a gradient-based manner even in fully black-box contexts [III.3], and is a capability that properly reflects the level of adversarial threat and in that way does not overestimate the empirical robustness; real attackers will compute gradients after all [17].
+Through proper threat modeling and self-adaptation, attacks can reach their full potential, enabling the development of effective defensive policies.
+To frame the need for adaptive evaluations in adversarial machine learning (AML) differently: a defense can be considered trustworthy only if it is evaluated against an optimal adversary.
+This mutual interdependence underscores the necessity for both attacks and defenses to be self-adaptive, thereby establishing the competitive, zero-sum dynamic inherent in their interaction.
+In this work, we examine robustness from both perspectives: first, how to fully optimize decision-based attacks, and second, how to devise reliable countermeasures.
+We explore both offensive and defensive strategies in depth, and make the following key contributions:
+  1. 1.
     We demonstrate that active defenses against decision-based attacks are a necessary but insufficient complement to model hardening.
 Active defenses are inevitably bypassed by self-adaptive attackers however, and necessitate self-adaptive defenses too.
+  2. 2.
     To facilitate reasoning on adaptive attacks and defenses, we introduce a unified framework called “Adversarial Markov Games” (AMG).
-We demonstrate how adversaries can optimize their policy and evade active detection at the same time; as a counter, we propose a novel active defense and employ RL agents to adapt and optimize both.
-For reproducibility and follow-up work, we open-source our code [^1] https://anonymous.4open.science/r/AMG-AD16.
-  • In an extensive empirical evaluation on image classification and across various adversarial settings, we validate our theoretical analysis and show that self-adaptation through RL outperforms vanilla black-box attacks, model hardening defenses like adversarial training, and notably both the state-of-the-art adaptive attacks and stateful defenses.
-Our work highlights that in the domain of black-box AML, robust evaluations should go a step further than adapting attacks: both attacks and defenses should have the capability to optimize their operation through interaction and in direct response to other agency in their environment.
+We demonstrate how adversaries can optimize their attack policy and evade active detection at the same time; as a counter, we develop a novel active defense and employ RL agents to adapt and optimize both.
+  3. 3.
+    In an extensive empirical evaluation on image classification and across a wide set of adversarial scenarios, we validate our theoretical analysis and show that self-adaptation with RL outperforms vanilla black-box attacks, model hardening defenses like adversarial training, and notably both the state-of-the-art adaptive attack (OARS [18]) and stateful defense (Blacklight [19]).
+This supports self-adaptation as an essential component when evaluating robustness to black-box attacks.
+  4. 4.
+    For reproducibility, and to facilitate further research, we open-source our code [^1] https://anonymous.4open.science/r/AMG-AD16.
+Our work highlights that in the domain of black-box AML, robust evaluations should go a step further than adapting attacks: both attacks and defenses should have the capability to optimize their strategies through interaction and in direct response to other agency in their environment.
 The remainder of the paper is structured as follows: Section II provides the necessary background on the domain and reviews the related work.
 Section III introduces and motivates our theoretical analysis of robustness under decision-based attacks.
-Section IV elucidates the threat model and the concrete design choices.
+Section IV explains the threat model and the concrete design choices.
 In Section V we elaborate on our experimentation and analyze our results.
-We conclude with Section VI where we discuss insights, limitations and challenges.
+We conclude with Section VI where we discuss key insights, limitations and challenges.
 II Preliminaries
-In this work, we focus on the category of adversarial attacks known as decision-based, a subset of query-based attacks that operate solely on the hard-label outputs of the model and are regarded as a highly realistic and pervasive threat in AI-based cybersecurity environments.
-Despite the lack of the closed-form expression of the model under attack, given enough queries the effectiveness of such black-box attacks can match and even surpass that of white-box techniques [17, 15].
+In this work, we focus on the category of adversarial attacks known as decision-based, a subset of query-based attacks that operate solely on the hard-label outputs of the model and are a highly realistic and pervasive threat in AI-based cybersecurity environments.
+Despite the lack of the closed-form expression of the model under attack, given enough queries their effectiveness can match the one of white-box techniques [20, 15].
 II-A Attacks & Mitigations
-While adversarial attacks have been extensively researched in both white and black-box contexts, defenses have predominantly focused on the white-box context [4, 5].
+While adversarial attacks have been extensively researched in both white and black-box contexts, defenses have predominantly focused on white-box [4, 5].
 As the black-box setting discloses considerably less information, a seemingly intuitive conclusion is that white-box defenses should suffice for the black-box case too.
@@
 The vast majority of adversarial defenses provide either limited robustness or are eventually evaded by adapted attacks [14].
-Characteristically, preprocessing defenses are often bypassed by expending queries for reconnaissance [21].
+Characteristically, preprocessing defenses are identified and bypassed by expending queries for reconnaissance [24].
 The partial exception to this rule is adversarial training.
@@
 The inner maximization loop finds an adversarial example of $x_{i}$ with label $y_{i}$ for a given $L_{p}$-norm (with $L_{p}\in\{0,1,2,\inf\}$), such that $\|\delta_{i}\|_{l}\leq\epsilon$ and $h_{\phi}(x_{i}+\delta_{i})\neq y_{i}$.
-The outer loop is the standard minimization task typically solved with stochastic gradient descent.
-While the convergence and robustness properties of adversarial training have been investigated through the computation of the inner maximization step and by interleaving normal and adversarial training [5], the min-max principle is conspicuous: minimize the possible loss for a worst case (max) scenario.
+The outer loop is the ordinary minimization task, typically solved with stochastic gradient descent.
+While the convergence and robustness properties of adversarial training have been investigated through the computation of the saddle point and by interleaving normal and adversarial training [5], the min-max principle is conspicuous: minimize the possible loss for a worst-case (max) scenario.
 II-B Stateful Defenses
-All decision-based attacks share properties that can be useful in devising defenses against them, on top of adversarial training.
-Such a property is their inherent sequentiality: by following an attack policy towards the optimal adversarial example, the generated candidates are correlated.
-This might not be the case for the queries themselves however, as the adversary might employ transformations the model is invariant to, like the query blinding strategy in Chen et al.
+Decision-based attacks possess properties that can be valuable for devising defenses against them, in addition to adversarial training.
+One such property is their intrinsic sequentiality: by following a policy toward the optimal adversarial example, the generated candidates are correlated.
+Note that this might not hold for the queries themselves, as the adversary may apply transformations that the model is invariant to, such as the query blinding strategy in Chen et al.
 [13].
-This work is also the first to employ a stateful defense against query-based attacks.
-Another stateful defense is PRADA [22], built against model extraction but which also works against model evasion.
-The efficacy of these approaches however rests on the assumption that queries can be consistently linked (through metadata like IP or account, cf.
-LABEL:tab:comparison) to uniquely identifiable actors – that also show limited to no collaboration – so that a buffer of queries can be built for each.
-This limitation was recently addressed, together with the ensuing scalability issues in the Blacklight defense, by resourcefully employing hashing and quantization [23].
-It remains a similarity-based defense however, thus vulnerable to circumvention if an adversary can find a query generation policy that preserves attack functionality while evading detection; the recent OARS work achieved this by adapting existing attacks through the rejection signal Blacklight returns [24].
-Ultimately, any (stateful) defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be representative only if the adversary has exhausted their adaptive offensive capabilities.
+This work is the first to employ a stateful defense against query-based attacks.
+Another stateful defense is PRADA [25], devised against model extraction but effective against evasion too.
+These approaches assume however that queries can be consistently linked (via metadata like IP or account, cf.
+LABEL:tab:comparison) to uniquely identifiable actors – who also exhibit limited to no collaboration – so that a query buffer can be built for each.
+This limitation, together with the scalability issues, was recently addressed in the Blacklight defense, by employing hashing and quantization [19].
+Blacklight remains a similarity-based defense, thus vulnerable to circumvention if an adversary can find a query generation policy that preserves the attack functionality while evading detection.
+OARS achieved this by adapting existing attacks through the rejection signal Blacklight returns [18].
+Ultimately, any (stateful) defense has to balance the trade-off between robust and clean accuracy; as we demonstrate in this work, this trade-off can be measured reliably only if the attacker is properly adaptive.
+II-C On Being Adaptive
+The correct way to evaluate any proposed defense is against adaptive attacks, that is with explicit knowledge of the concrete mechanisms of a defense [14].
+In computer security this is known as the stipulation that security through obscurity does not work, as the robustness of defenses should not rely on keeping their mechanism secret.
+If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, active defenses like stateful detection are the counterpart to decision-based attacks, and as we will further demonstrate, also the necessary complement to hardening a model against them.
+At the same time, the level of threat that attacks pose is often unclear or not thoroughly evaluated.
+Previous work has demonstrated that the loss functions and parameters of attacks are often suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness [15, 26].
+This underestimation is further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have.
+The true performance of attacks therefore rests on the ability to adapt their operation policy and their evasive capabilities in tandem.
 Figure 1: In AML, adaptive attacks are those with the capabilities (knobs) to bypass a defense; adaptive control is rather the precise tuning of all the known knobs.
@@
 For instance in HSJA [7], radius, steps, and jump are parameters of the attack, while rotate and translate are transformations that can evade a similarity-based defense.
-II-C On Being Adaptive
-The correct way to evaluate any proposed defense is against adaptive attacks, that is with explicit knowledge of the inner mechanisms of a defense [14].
-In computer security this is known as the stipulation that security through obscurity does not work, as the robustness of defenses should not rely on keeping their way of functioning secret.
-If model hardening – for instance by adversarial training – is the defensive counterpart to white-box attacks, active defenses like stateful detection are the counterpart to decision-based attacks, and as we will further demonstrate, also the necessary complement to hardening a model against them.
-At the same time, the level of threat that attacks pose is often unclear or not thoroughly evaluated.
-Previous work has demonstrated that loss functions and parameters of attacks are often suboptimal, leading to underestimating their performance and thus overestimating the claimed degree of robustness [15, 25].
-This underestimation is further aggravated in decision-based contexts, where the attacker is largely oblivious of any preprocessing or active defenses the black-box system might have.
-The practical effectiveness of attacks therefore rests on the ability to adapt the policies that govern their operation and their evasive capabilities in tandem.
 In AML, “adaptive” by convention refers to attacks with full knowledge of how a defense works and the tools to bypass it; we denote such attacks as adapted.
-In our work, we expand the term to include adaptive control, defined as the ability of a system to self-adapt: automatically reconfigure itself in response to changes in the dynamics of the environment in order to achieve optimal behavior [16].
-One can think of adaptive control in the sense “attack optimization” is used by Pintor et al.
-[25], but for black-box systems.
-Typically, what is to be controlled is well-defined and known in advance.
-The moment however we consider adaptive evaluations, new controls are potentially implied: in a similarity-based defense for instance, such controls would be input transformations the model is invariant to.
-To flesh out the twofold meaning of adaptive, one has to both invent new knobs [26] (which is the conventional understanding of adaptive, and still very hard to automate), and dynamically control their correct configuration that would lead to the optimal result (self-adaptive).
+In this work we expand the term to include adaptive control, defined as the ability of a system to self-adapt: automatically reconfigure itself in response to changes in the dynamics of the environment in order to achieve optimal behavior [16].
+We use adaptive control in the sense “attack optimization” is used by Pintor et al.
+[26], but here for black-box systems.
+What is to be controlled is typically known in advance and well-defined.
+However, the moment we consider adaptive evaluations, new controls are directly implied: in a similarity-based defense for instance, such controls would be transformations to the input that the model is invariant to.
+To flesh out the twofold meaning of adaptive, one has to both invent new knobs [27] – the conventional understanding of adaptive, intractable to automate yet – and dynamically control their correct configuration that would lead to the optimal result (self-adaptive).
 We conceptualize this more general definition of adaptive, essential for having accurate evaluations against decision-based attacks, in Figure 1.
-II-D Research Gap
-Prior work has focused on adapted attacks, which incorporate general knowledge of any defenses and empirically configured to evade it [17, 7, 6].
-Defenses also follow the same adapted paradigm of empirically defined and fixed parameters [13, 23].
-Our observation is that neither of them are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and with respect to other adaptive agents in it, with clear limitations when the latter is a given, e.g.
-in cybersecurity.
-To bridge this gap, we provide a theoretical treatment and empirical study of existing and novel methodologies adapting through direct interaction with their environment, denoting them as self-adaptive.
-Our work builds on a long line of prior research that focuses on both sides of the competition between adversaries and defenses.
-Carlini and Wagner [17] show that evaluating existing attacks out-of-the-box is insufficient and that adapted white-box attackers can break defensive distillation.
-Bose et al.
-[27] propose Adversarial Examples Games (AEG), a zero-sum game between a white-box attacker and a local surrogate of the target model family.
-At the equilibrium the attacker can generate adversarial examples that have a high success rate against models from the same family, constituting a zero-query, non-interactive approach for generating transferable adversarial examples.
-Pal et al.
-[28] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium.
-Feng et al.
-[24] introduce OARS: adaptive versions of existing attacks that bypass Blacklight [23], the state-of-the-art stateful defense.
-To function, OARS presupposes the rejection signal that a defense like Blacklight returns; a strong assumption that as we show in this work does not have to hold for stateful defenses.
-As we demonstrate in section V and Table III, Blacklight can be bypassed without assuming rejection, while the novel stateful defense we introduce can fully withstand the OARS attack.
-As the most relevant and representative threat against real-world AI systems, in this work we scope on decision-based, interactive attacks and defenses.
-We contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive and thus also the defensive potential is properly assessed.
-In the remainder of the paper the term “adaptive” subsumes adaptive control, and is used interchangeably with “self-adaptive”.
-For what is conventionally known as adaptive evaluations in AML, we use the term “adapted”.
-To facilitate comparison, in LABEL:tab:comparison we highlight the most important aspects of our work as the synthesis of adaptive black-box attacks and defenses in a unified framework, and situate it with respect to other relevant and state-of-the-art works in AML.
-Note the importance for an attack to function without assuming rejection, and respectively for a defense to function without access to query metadata like IP addresses or accounts.
-TABLE I: Prominent decision-based attacks and defenses and their individual aspects. While all works are situated either on the offensive or the defensive side, ours brings these two together.
+TABLE I: Positioning of our work relative to prominent decision-based attacks and defenses and their individual properties.
+While prior works focus exclusively on offense or defense, ours unifies and reasons from both perspectives.
 Work Offensive Defensive
@@
 Blacklight (2022) [19]
-AMG – Our work
+Our work
+II-D Research Gap
+Prior work has focused on adapted attacks, which incorporate general knowledge of any defenses, then empirically configured to evade it [20, 7, 6].
+Defenses also follow the same adapted paradigm of empirically defined and fixed operation [13, 19].
+Our observation is that neither of them are formalized or performed in a fully adaptive manner, that is in response to how they influence their environment and with respect to other adaptive agents in it, with clear limitations when the latter is a given, e.g.
+in cybersecurity.
+To bridge this gap, we provide a theoretical analysis and an empirical study of existing and novel methodologies adapting through direct interaction with their environment, denoting them as self-adaptive.
+Our work builds on a long line of prior research that focuses on both sides of the competition between adversaries and defenses.
+Carlini and Wagner [20] show that evaluating existing attacks out-of-the-box is insufficient and that adapted white-box attackers can break defensive distillation.
+Bose et al.
+[29] propose Adversarial Examples Games, a zero-sum game between a white-box attacker and a local surrogate of the target model family.
+At the equilibrium the attacker can generate adversarial examples that have a high success rate against models from the same family, constituting a zero-query, non-interactive approach for generating transferable adversarial examples.
+Pal et al.
+[30] propose a game-theoretic framework for studying white-box attacks and defenses that occur in equilibrium.
+Feng et al.
+[18] introduce OARS: adaptive versions of existing attacks that bypass Blacklight [19], the state-of-the-art stateful defense.
+To function, OARS presupposes the rejection signal that a defense like Blacklight returns; a strong assumption that as we show in this work does not have to hold for stateful defenses.
+As we demonstrate in section V and Table III, Blacklight can be bypassed without assuming rejection, while the novel stateful defense we introduce can fully withstand the OARS adaptive attack.
+As the most relevant and representative threat against real-world AI systems, in this work we scope on decision-based, interactive attacks and defenses.
+We contribute a theoretical and practical framework for self-adaptation, under which the full extent of the offensive and thus also the defensive potential is properly assessed.
+In the remainder of the paper the term “adaptive” subsumes adaptive control, and is used interchangeably with “self-adaptive”.
+For what is conventionally known as adaptive evaluations in AML, we use the term “adapted”.
+To facilitate comparison, in LABEL:tab:comparison we highlight the most important aspects of our work as the synthesis of adaptive black-box attacks and defenses in a unified framework, and situate it with respect to prominent and state-of-the-art works in AML.
+Note the importance for an attack to function without access to a rejection signal, and respectively for a defense to function without access to query metadata like IP addresses or accounts.
 III Theoretical Framework
-To investigate the robustness to evasion of real-world systems, two closely related perspectives are pertinent: a) thwarting decision-based attacks, and b) adapting attacks and evasive capabilities in tandem.
-For the latter, when attacks (and defenses) are evaluated in a non-adaptive manner, in the expansive sense we outlined in subsection II-C, results are rendered incomplete and untrustworthy [14, 25].
-Nevertheless, when offensive or defensive techniques adapt the environments that they inhabit become non-stationary [30], putting further pressure on the IID foundations that ML builds on.
-To understand the implications of such techniques adapting, we perform a theoretical analysis of the possible interactions around the exposed interface of an ML-based system, interactions that can be more generally considered as sequential zero-sum games [31, 32, 27].
-We initiate our analysis from adversaries, as the original factor and cause.
+In this section, we abstract through the individual properties of decision-based attacks and defenses to extract more general insights than a purely empirical study would render.
+To investigate how robust real-world systems are to evasion, two related perspectives are crucial: a) resisting decision-based attacks, and b) adapting attacks and evasive capabilities together.
+When attacks (and defenses) are evaluated in a non-adaptive manner, in the expansive sense we outlined in subsection II-C, results are unreliable [14, 26].
+Note that with offensive or defensive methodologies adapting, their environments become non-stationary [31], putting further pressure on the IID foundations that ML builds on.
+To understand the implications of this adaptation, we perform an analysis of the possible interactions on the interface of an ML-based system, interactions that can be more generally considered as sequential zero-sum games [32, 33, 29].
 In the following sections, introduced terms and notation are highlighted in orange.
@@
 The most compelling threat that deployed ML-based systems face are decision-based black-box attacks, where no access is assumed to the model or its parameters, only the capacity to submit queries and receive hard-label responses.
-One of the first decision-based attacks was Boundary Attack [6].
-A large number of others followed – each inventive in its own way – that manage to improve the overall performance, typically measured as the lowest perturbation achieved for the minimal amount of queries submitted.
+One of the first decision-based attacks was Boundary Attack [6], followed by others that improve the overall performance, typically measured as the lowest perturbation achieved for the minimal amount of queries submitted.
 Prominent examples are HSJA [7], Guessing Smart (BAGS) [28], Sign-Opt [34], Policy-driven (PDA) [8], QEBA [35], and SurFree [36].
-White-box attacks like C&W [17] cannot function in black-box environments as there is no closed-form description of the inference pipeline.
+White-box attacks like C&W [20] do not function in black-box environments, as there is no access to the inference pipeline.
 To facilitate optimization, decision-based attacks commonly initialize from a sample belonging to the target class, as it can be considered an adversarial example with an unacceptably large perturbation.
 This switch allows the task to be solved continuously, by minimizing the perturbation while always staying on the adversarial side of the boundary.
-Decision-based attacks share further common aspects in their functioning, ones that we can abstract through: given starting and original samples $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{g}$ and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{c}$ respectively, the goal is to iteratively propose adversarial candidates $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{t}$, until the distance $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\delta=d(x_{t},x_{c})$ is minimized.
+Decision-based attacks share further common aspects in their function that we can abstract through: given starting and original samples $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{g}$ and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{c}$ respectively, the goal is to iteratively propose adversarial candidates $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x_{t}$, until the distance $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\delta=d(x_{t},x_{c})$ is minimized.
 This process follows different algorithmic approaches that represent different geometrical intuitions; we can describe it more generally by means of a candidate generation policy:
-$\pi_{\theta}^{\mathcal{A}}=P\>(x_{t}|x_{g},x_{c},p^{\mathcal{A}},s^{\mathcal{A}})$ (2)
+$\pi_{\theta}^{\mathcal{A}}=P\>(x_{t}|x_{g},x_{c},p^{\mathcal{A}},s^{\mathcal{A}}),$ (2)
 that given $x_{g}$ and $x_{c}$, with $p^{\mathcal{A}}$ the parameters and $s^{\mathcal{A}}$ the state of the attack, generates a candidate $x_{t}$.
-If we consider that attacks execute over discrete time steps, and assume the model always answers, then the attack procedure can be construed as a Markov Decision Process (MDP) to be solved, by finding the parameters $\theta$ that minimize $\delta$ in the least amount of queries.
+As attacks execute over discrete time steps, if we assume that the model always answers the attack procedure can be construed as a Markov Decision Process (MDP) to be solved, by finding the parameters $\theta$ that minimize $\delta$ for a given number of queries.
 Consider now a multinomial image classification model $\mathcal{M}$ under attack, with a discriminant function $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, that for each input $x\in[0,1]^{d}$ generates an output $y:=\{y\in[0,1]^{m}|\sum_{c=1}^{m}y_{c}=1\}$ – a probability distribution over the $m$ classes.
-By definition, black-box environments preclude access to these probabilities; instead one can only observe the decision of the classifier $C$ that returns the highest probability class:
+As black-box environments prevent access to these probabilities, one can only observe the decision of the classifier $C$ that returns the highest probability class:
 $C(x):=\operatorname*{arg\,max}_{c\in[m]}F_{c}(x)=D(F_{c}(x))$ (3)
@@
 This process can be facilitated through a function $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\psi$ which given a perturbed example $x_{t}$ at step $t$, it returns a binary indicator of success:
-$\psi(x_{t})=\begin{cases}1&\text{if}\quad C(x_{t})=c_{o}\\ -1&\text{if}\quad C(x_{t})\neq c_{o}\end{cases}$ (4)
-As long as the model responds, $\psi$ can always be evaluated, it thus constitutes an essential mechanism upon which decision-based attacks build.
+$\psi(x_{t})=\begin{cases}+1&\text{if}\quad C(x_{t})=c_{o}\\ -1&\text{if}\quad C(x_{t})\neq c_{o}\end{cases}$ (4)
+As long as the model responds, $\psi$ can always be evaluated, and constitutes the fundamental mechanism upon which decision-based attacks build.
 The adversarial goal can then be described as the following constrained optimization problem:
@@
 where the distance metric $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}d$ is an $\ell_{p}$-norm, with $p\in\{0,1,2,\inf\}$.
-As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this optimization task highlights the fuzzy nature of adversarial examples, something that is further aggravated in domains where visual proximity is of little importance.
+As the threshold between adversarial and non-adversarial relies on the subjectivity of human perception, this highlights the indefinite nature of adversarial examples, further exemplified in domains where visual proximity is of little importance.
 Successful or unsuccessful adversarial examples are therefore delimited by an threshold $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\epsilon$ on perturbation, where $d(x,x_{t})\leq\epsilon$.
-While this constraint might not intuitively translate to non-visual domains, we nonetheless contend that minimal perturbation still remains a desideratum of adversarial examples.
-The fact that real-world attacks are black-box and decision-based does not make them any less effective.
+Real-world attacks being black-box does not make them less effective.
 For instance, HSJA is guaranteed to converge to a stationary point of Eq.
@@
 Given typical $\epsilon$ values for imperceptibility, this results in high attack success rates, even against adversarially trained models.
-The limited effectiveness of adversarial training against decision-based attacks can be attributed to the fundamentally out-of-distribution (OOD) nature of adversarial examples, and the inherent challenges in the saddle point optimization problem in Eq.
+The limitations of adversarial training against decision-based attacks can be attributed to the out-of-distribution (OOD) nature of adversarial examples, and the saddle point optimization problem of Eq.
 (1) that make it difficult for algorithms to converge to a global solution.
-Additionally, decision-based attacks are challenging to incorporate during stochastic gradient descent (SGD) because these attacks rely on boundary navigation, and the farther the model is from convergence, the less effective the attack is.
-Scalability is another concern: while adversarial training typically involves a few steps (1-50) of white-box attacks like FGSM or PGD, decision-based attacks can take orders of magnitude more steps (queries) to produce an adversarial example.
+Furthermore, incorporate decision-based attacks in training is not scalable as it can take orders of magnitude more steps (queries) to produce an adversarial example, than white-box attacks which take a few steps (1-50 in e.g.
+PGD [4]).
 Decision-based attacks search for the optimal parameters $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\theta^{*}$ of the generation policy (2), those that given $x^{i}_{c}$, with $i$ denoting the i-th adversarial episode, minimize Eq.
@@
 where $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}x^{i}_{b}$ is the best adversarial example generated by policy $\pi_{\theta}^{\mathcal{A}}$ during episode $i$.
-Given the dimensionality, it can be intractable to learn a policy that modifies the inputs at the feature space directly [36]; CIFAR-10, for instance, has more than 3K features to perturb.
-In AI-enabled systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can therefore assume that $F_{0}=F_{1}=...=F_{t},\forall t$.
-While this is representative of real-world settings, it is also what enables adversaries to discover adversarial examples that were not identified beforehand.
-The fact however that adversaries follow a candidate generation policy, introduces a querying behavior which can be observed and exploited by a defensive methodology.
-Consequently, while model-hardening approaches like adversarial training are necessary, they can be insufficient in defending from decision-based attacks.
+Given the dimensionality of the input, it can be intractable to learn a policy that modifies the feature space directly [37]; CIFAR-10, for instance, has more than 3K features to perturb.
+In AI-based systems, the best practice is to freeze the model after validation so that no novel issues are introduced by retraining: for all queries $x_{t}$ submitted during an attack session, we can therefore assume that $F_{0}=F_{1}=...=F_{t},\forall t$.
+While this is representative of real-world settings, it also enables adversaries to discover adversarial examples that were not identified beforehand.
+Consequently, while model-hardening through adversarial training is necessary, it can also be insufficient against decision-based attacks like HSJA.
 Proposition III.1.
-Let $F_{c}$ be the discriminant function of the adversarially trained model $\mathcal{M}$.
-Then in order for $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]\geq\epsilon$ in HSJA, two capabilities are necessary: a) a decision function $D^{\prime}\neq\operatorname*{arg\,max}$, and b) additional context $\tau$ s.t for adversarial query $x_{t}$, $C(x_{t})=D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$.
+Let $F_{c}$ denote the discriminant function of an adversarially trained model $\mathcal{M}$, and let $C(x)=D(F_{c}(x))$ denote its classifier.
+Then in HSJA, to satisfy $\mathbb{E}[\sum_{i=1}^{N}d(x_{b}^{i},x_{c}^{i})]\geq\epsilon$ it is necessary that: (a) $D\neq\operatorname*{arg\,max}$, and (b) context $\tau$ exists s.t.
+for some query $x_{t}$, $D(F_{c}(x_{t}))\neq D^{\prime}(\tau,F_{c}(x_{t}))$, where $D^{\prime}$ is a stateful extension of $D$.
 Intuitively, HSJA operates in 3 stages which repeat: a binary search that puts $x_{t}$ on the decision boundary, a gradient estimation step, and projection step along the estimated gradient.
 If the model always responds truthfully, the adversary will be able to accurately perform all these steps and converge to the optimal adversarial; without loss of generality, we can extend this intuition to other decision-based attacks which navigate the boundary.
-Secondly, the model should be able to differentiate between two, otherwise identical, queries when one is part of an attack and the other is not, and this is possible through a stateful representation; see Appendix A for the proof.
+Secondly, the model should be able to distinguish between two, otherwise identical, queries, when one is part of an attack and the other is not, a capability achievable through statefulness; see Appendix A for the proof.
 III-B Defenses
-Proposition III.1 indicates that alternative classification policies are necessary in the presence of decision-based attacks, for instance classification with rejection or intentional misdirection.
-In related work, rejection has been realized in the form of conformal prediction where model predictions are sets of classes including the empty one, or learning with rejection [11, 37]; while misdirection has emerged as a technique in adversarial RL and cybersecurity domains [10, 12].
-While adversarially training the discriminant function $F$ empirically shows the capacity to resist decision-based attacks, the manner in which the model responds has a complementary potential.
-This gap between the empirical and theoretically possible robustness to decision-based attacks is the locus where a distinct from model hardening, active and adaptive defense can emerge.
+Proposition III.1 suggests that alternative classification policies are necessary in the presence of decision-based attacks, e.g.
+classification with rejection or intentional misdirection.
+Rejection has been realized in the form of conformal prediction, where model predictions are sets of classes including the empty one, or learning with rejection [11, 38]; while misdirection has emerged as a technique in adversarial RL and cybersecurity domains [10, 12].
+While adversarially training the discriminant function $F$ empirically shows some degree of robustness to decision-based attacks, the manner in which the model responds has a complementary potential.
+The gap between the empirical and theoretically achievable robustness is the source for an active defense distinct from model hardening.
 Active defenses have direct implications on attacks themselves however.
@@
 $\pi_{\phi}^{\mathcal{D}}=P(\alpha_{t}|x_{t},s^{\mathcal{D}}_{t}),\>\alpha\in\{0,1\}$ (7)
-with $x_{t}$ the query, $s^{\mathcal{D}}_{t}$ the state for the defense as created by past queries, and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\alpha$ the binary decision: when the query is deemed adversarial, it is rejected by returning $\alpha=1$.
-If this policy is stationary, the environment dynamics in turn become stationary too, and thus together with the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved.
+with $x_{t}$ the query, $s^{\mathcal{D}}_{t}$ the state for the defense as created by past queries, and $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\boldsymbol{\alpha}$ the binary decision: for queries deemed adversarial, $\boldsymbol{\alpha}=1$, otherwise $\boldsymbol{\alpha}=0$.
+When this policy is stationary, the environment dynamics become stationary in turn, thus besides the adversarial task itself, bypassing the defense can also be formulated as an MDP to be solved (Figure 2).
 In two-player, zero-sum games, the moment an agent follows a stationary policy, it becomes exploitable through the reward obtained by an adversary [39].
-Active defenses, a consequence of decision-based attacks, entail therefore adaptive adversaries.
+Active defenses, as consequence of decision-based attacks, entail therefore adaptive adversaries.
 Proposition III.2.
@@
 However, there is a second reason to consider adaptive attacks even in the absence of active defenses, as attack policies are often suboptimal with their default, empirically defined parameters.
-Adapting attack policies is essentially the optimization of these parameters, and as an approach has proven very effective in other black-box or expensive-to-evaluate contexts, like Neural Architecture Search and Data Augmentation [39, 40, 41].
-Our results in Section V further indicate the correspondence between adaptive and self-optimizing, showing that adaptive attacks consistently outperform non-adaptive, particularly against active defenses.
-Consider now an active defense that is based on a similarity or a conformal metric.
-In the twofold meaning we introduced in Section II, adaptive attack implies the capability to bypass a similarity based defense; adaptive control implies optimization instead, the active tuning of all the available tools to evade the defense and minimize the perturbation (cf.
-Fig.
-1).
-The revised adversarial objective then is to find the optimal policy that also evades detection, and the way to achieve this is by adapting the candidate generation policy (2) itself.
+Adapting attack policies is essentially the optimization of these parameters, and as an approach has proven very effective in other black-box or expensive-to-evaluate domains, like Neural Architecture Search and Data Augmentation [40, 41, 42].
+Our results in section V further indicate the correspondence between adaptive and self-optimizing, showing that adaptive consistently outperform non-adaptive attacks, particularly against active defenses.
+Consider now an active defense that is based on a similarity or conformal metric.
+In the twofold meaning we introduced in subsection II-C, adaptive attack implies the capability to bypass a similarity based defense; adaptive control implies optimization instead, the active tuning of all the available tools to evade the defense and minimize the perturbation (cf.
+Figure 1).
+The updated adversarial objective then is to find the optimal policy that also evades detection, and the way to achieve this is by adapting the candidate generation policy (2) itself.
 Notably, and despite the black-box and discontinuous nature of the task, this optimization can be fully gradient-based.
@@
 The proof is included in Appendix A.
-So far we have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) adaptive attacks can become optimal in terms of both evasion and efficiency by observing and adapting to the discrete model decisions.
+We thus have established that, a) in the presence of decision-based attacks, active defenses are necessary, yet conditional on adversarial agency they are insufficient and, b) adaptive attacks can become optimal in terms of both evasion and efficiency by observing and adapting to the discrete model decisions.
 To complete the puzzle, the last piece is turning active defenses also adaptive.
@@
 maximizing expectation $\mathbb{E}[\sum_{x_{t}\in\tau_{i}}P(\alpha_{t}|x_{t},s^{\mathcal{D}}_{t})]$, by adapting its policy against the optimal evasive policy ${\pi^{\mathcal{A}}_{\theta^{*}}}$.
+Proof.
+Since the game is zero-sum, we may define the defensive policy reward $\rho$ on any trajectory $\tau_{i}=(x_{1},\dots,x_{T})$ as $\rho(\tau_{i})=\sum_{x_{t}\in\tau_{i}}\alpha_{t}$.
+Treating the adversary’s policy $\pi_{\theta^{*}}^{\mathcal{A}}$ as fixed, we perform gradient ascent on the expected reward $J(\phi)\;=\;\mathbb{E}_{\tau_{i}\sim(\pi_{\theta^{*}}^{\mathcal{A}},\,\pi_{\phi}^{\mathcal{D}})}\bigl{[}\rho(\tau_{i})\bigr{]}$.
+Under standard smoothness assumptions, this converges to $\phi^{*}=\arg\max_{\phi}J(\phi)$, which is precisely the defender’s best response to $\pi_{\theta^{*}}^{\mathcal{A}}$.
+∎
+III-C Adversarial Markov Games
+By reasoning on both offensive and defensive capabilities, we highlight why one cannot consider them independently.
+As adaptive attacks and defenses are logical consequences of each other, their composition forms a turn-taking competitive game.
+A precise game-theoretic formulation requires full knowledge of the environment: its models, players and their utility functions, as well as the permitted interactions and the transition dynamics, something typically intractable in this and other cybersecurity settings.
+Model-free methods however can learn optimal (offensive and defensive) responses directly through interaction with their environment [12, 43], avoiding the need for explicit modeling or solving the NP-hard bi-level optimization problem of Eq.
+(1[44].
+To that end, Turn-Taking Partially-Observable Markov Games (TT-POMGs), introduced by Greenwald et al.
+[45], generalize Extensive-Form Games (EFGs) that model non-cooperative, sequential decision-making games of imperfect and/or incomplete information.
+TT-POMG is a suitable formalism for decision-based attacks and defenses, with the added benefit that it can be transformed into an equivalent belief state MDP, significantly simplifying its solution.
+Prior work has explored the competition underlying adversarial example generation in no-box and white-box settings [29, 46].
+We instead focus on decision-based, interactive environments, with unknown but stationary dynamics: all other agents present are considered part of the environment and therefore fixed in their behavior.
+By folding the strategies of other agents into the transition probabilities and the initial probability distribution of the game, an optimal policy computed in the resulting MDP will correspond to the best-response strategy in the original TT-POMG.
+The congruence between TT-POMGs and MDPs has both theoretical and practical value for securing AI-based systems: once adversarial agents and their capabilities are identified through rigorous threat modeling, the best-response strategy in the simulated environment yields the optimal defense.
 Figure 2: Schematic model of an AMG environment.
-Due to the inherent uncertainty of behavior at either side of the interface, it is a partially observable RL environment mirrored for each agent where one’s decisions become the other’s observations.
+Due to the inherent uncertainty of behavior at either side of the interface, it is a partially observable MDP, mirrored for each agent where one’s decisions become the other’s observations.
 (I) denotes an adaptive attacker (cf.
@@
 1), (II) model hardening (passive defense), and (III) an active defense.
-Proof.
-As offensive and defensive policies are strictly competitive, we can define a reward $\rho$ of the defensive policy as $\rho(\tau_{i})=\sum_{x_{t}\in\tau_{i}}\alpha_{t}$.
-Then by fixing the optimal evasive policy $\pi_{\theta^{*}}^{\mathcal{A}}$, and updating $\pi_{\phi}^{\mathcal{D}}$ with expected reward $\mathbb{E}_{\pi^{\mathcal{D}}_{\phi}}[\rho(\tau_{i})]$ gradients, at convergence the resulting ${\pi^{\mathcal{A}}_{\theta^{*}}}$ will be the optimal defensive policy.
-∎
-III-C Adversarial Markov Games
-By reasoning on both offensive and defensive capabilities, we highlighted why we cannot consider them independently.
-As adaptive attacks and defenses are logical consequences of each other, their composition results in a turn-taking competitive game.
-A precise game-theoretic formulation requires however the exact analytical description of the environment: the model, the players and their utility functions, as well as the permitted interactions and the transition dynamics, something exceedingly intractable in this as well as most other cybersecurity contexts.
-Model-free methods however can learn optimal offensive and defensive responses directly through interaction with their environment [12, 42], obviating the need to learn a model of it or to find exact solutions to the bi-level optimization problem in Eq.
-(1) that is NP-hard [43].
-To that end, Turn-Taking Partially-Observable Markov Games (TT-POMGs) introduced by Greenwald et al.
-[44] is a generalization of Extensive-Form Games (EFGs), widely used representations for non-cooperative, sequential decision-making games of imperfect and/or incomplete information.
-TT-POMGs are an apt formalism for decision-based attacks and defenses, with the additional useful property that they can be transformed to equivalent belief state MDPs, significantly simplifying their solution.
-The competition underlying adversarial example generation has been explored in no-box and white-box settings [27, 45].
-We instead focus on decision-based, interactive environments, and assume that they have unknown but stationary dynamics: any agency present is considered part of the environment and therefore fixed in its behavior.
-By folding the strategies of other agents into the transition probabilities and the initial probability distribution of the game, an optimal policy computed in the resulting MDP will correspond to the best-response strategy in the original TT-POMG.
-The congruence between TT-POMGs and MDPs is useful not only from a theoretical perspective, but also for its practical implications in the security of ML-based systems: provided that adversarial agents and their capabilities can be identified through rigorous threat modeling, computing the best response strategy in the simulated environment will correspond to the optimal defense.
-The environment that encompasses adversarial attacks, adversarial defenses, and benign queries, we describe it as an Adversarial Markov Game (AMG) – a special case of TT-POMG – and is depicted in Figure 2.
+We describe the environment that encompasses adversarial attacks, adversarial defenses, and benign queries, as an Adversarial Markov Game (AMG) – a special case of TT-POMG – depicted in Figure 2.
 Formally, we represent AMG as a tuple $\langle i,S,O,A,\tau,r,\gamma\rangle$
@@
 As with plausible assumptions we cannot assume access to the exact state of the other agent, the states $O^{\mathcal{D}}$, $O^{\mathcal{A}}$ are partial observations of the complete state $\mathcal{S}$ of the full game.
-When the competing agents that hold beliefs about each other are human, they engage in recursive reasoning expressed as [I believe that [my opponent believes [that I believe…]]].
-In the study of opponent modeling, considering other agent policies as stationary and part of the environment is equivalent to 0th level recursive reasoning: the agent models how the opponent behaves based on the observed history, but not how the opponent would behave based on how the agent behaves [46, 47].
+For instance, when the competing agents (holding beliefs about each other) are human, they engage in recursive reasoning expressed as [I believe that [my opponent believes [that I believe…]]].
+In the study of opponent modeling, considering other agent policies as a stationary part of the environment is equivalent to 0th level recursive reasoning: the agent models how the opponent behaves based on the observed history, but not how the opponent would act based on how the agent behaves [47, 48].
 In this work we consider more involved recursive reasoning out of scope, as AMGs can be solved by single-agent RL algorithms, and perform the empirical evaluation without building explicit models of opponent behavior.
-IV Environment Specification
+IV Threat Model
 The empirical study we conduct in Section V reflects diverse instantiations of the general theoretical framework introduced in Section III.
@@
 An assumption that influences the effectiveness of stateful detection is that queries can be attributed to UIDs, e.g., an IP address or a user account.
-However, adversaries can collude, create multiple accounts, use VPNs, or in fact accounts and IP addresses might not even be required to query the model.
+However, adversaries can collude, create multiple accounts, use VPNs, or in fact accounts and IP addresses might not even be necessary to query the model.
 To address this, we treat queries irrespective to their source.
 This is a strictly more challenging setting for stateful defenses, where we operate solely on the content of queries and not on any other metadata, similar to [19].
-Unlike Blacklight however, instead of rejecting queries, something that by itself can provide more information to the adversary and thus facilitate evasion (cf.
+Unlike Blacklight however, instead of rejecting queries, something that in itself provides more information to the adversary and thus facilitates evasion (cf.
 OARS [18], LABEL:tab:comparison), we misdirect by returning the second highest probability class.
-Furthermore, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution and so that is not trivial for a defense to tell adversarial noise apart.
+Furthermore, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution, so that is not trivial for a defense to tell adversarial noise apart.
 In summary, the black-box threat model we consider is delineated as follows:
@@
   • Adversary Knowledge: The model $\mathcal{M}$ is known as the black-box function that transforms inputs $x\in[0,1]^{d}$ to outputs $c\in[m]$, $m$ being the number of classes.
-The weights ${w}$ and the closed-form expression of $\mathcal{M}$ are unknown, as unknown is if an active defense $\pi_{\phi}^{\mathcal{D}}$ exists or not.
+The weights ${w}$ and the closed-form expression of $\mathcal{M}$ are unknown, as unknown is if an active defense $\pi_{\phi}^{\mathcal{D}}$ is present or not.
   • Defender Knowledge: The defender observes only the content of incoming queries, without knowing if they come from a benign user or the adversary.
-  • Adversary Capabilities: Adapt the parameters of the attack and optionally any evasive transformations; in essence, optimize the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$.
-  • Defender Capabilities For each query $x$, decide between answering truthfully with the true prediction $C(x)=c_{T}$ or misdirect with the second highest probability class $c_{S}$.
-Figure 3: Misdirection in a two-dimensional decision boundary.
-The adaptive defense controls a single parameter, the hypersphere radius around $k_{0}$ (the last known adversarial); for queries $x_{t}$ that fall within this hypersphere the model responds with a non-adversarial decision.
-$x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
+  • Adversary Capabilities: Adapt the parameters of the attack and of any evasive transformations; in essence, optimize the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$.
+  • Defender Capabilities: For each query $x$, decide between answering truthfully with the actual prediction $C(x)=c_{T}$, or misdirect with the second highest probability class $c_{S}$.
 [Similarity].
 Decision-based attacks typically follow a policy that generates successive queries: these exhibit degrees of similarity which can be quantified by an appropriate $L_{p}$ norm.
-If that norm is computed on the original inputs however, an adversary can adapt (see Figure 1) by employing evasive transformations the model is invariant to and bypass the similarity detection.
+If that norm is computed on the original inputs however, an adversary can adapt by employing evasive transformations the model is invariant to and bypass the similarity detection (cf.
+Figure 1).
 To account for this capability, we train a Siamese network with contrastive loss in order to learn a latent space $\mathcal{L}(\cdot)$ where similar inputs are mapped close together, unaffected by added noise or transformations on the inputs.
-For the stateful characterization of queries, we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign or undetected ones.
+For the stateful characterization of queries, we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign and undetected ones.
 […]
@@
 Depending on magnitude and composition of transformations $T$, the identity $F(T(x_{t}))=F(x_{t})$ might not always hold.
-As we also demonstrate in Section V, $T$ interferes with the perturbations the adversarial policy generates itself: the performance and evasiveness of an attack are thus in a natural trade-off.
-At this point we should inquire what is the correct composition of transformations $T$ to apply.
+As we also demonstrate in Section V, $T$ interferes with the perturbations of the adversarial policy: the performance and evasiveness of an attack are thus in a natural trade-off.
+At this point one should inquire what is the correct composition of transformations $T$ to apply.
 When shall $T$ be applied, and how does it affect the attack fundamentals?
@@
 Their trade-off illustrates why the twofold definition of adaptive is necessary in AML evaluations: first to impart the tools to accomplish to the task through the definition of what can be controlled, and then to find the precise optimal configuration and strategy of the attack.
+Figure 3: Misdirection in a hypothetical 2D decision boundary.
+The adaptive defense controls a single parameter, the hypersphere radius around $k_{0}$ (the last known adversarial); for queries $x_{t}$ that fall within this hypersphere the model responds with a non-adversarial decision.
+$x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
 [Agents & Environments].
@@
 Training is complicated further given that the experience upon which each agent learns arrives only after the opponent moves.
-We address these complications by developing custom learning environments for agents with asymmetries, with delayed experience collection, and asynchronous training, build with the OpenAI Gym and Stable-Baselines3 libraries.
+We address these challenges by developing custom learning environments (with the OpenAI Gym and Stable-Baselines3 libraries) for asymmetrical agents, with delayed experience collection, and asynchronous training.
 [States & Actions].
@@
 For evaluation, we define a range of scenarios intended to reflect all possible and realistic combinations between adversarial attacks and defenses, and their adaptive versions.
-Concretely, the research questions we want to validate are: 1) Are active defenses a necessary complement to model hardening and to what extent?
+Concretely, the research questions we want to evaluate are: 1) Are active defenses a necessary complement to model hardening and to what extent?
 2) Are attacks more threatening when adaptive, i.e., do they outperform their vanilla versions and evade active detection?
-3) If yes, can active defenses recoup their performance by also turning adaptive?
+3) If yes, to what extend active defenses recoup their performance by also turning adaptive?
 [Metrics].
@@
 BAGS is a stochastic, search-based method where every query submitted is a new and potentially better adversarial example.
-Contrastively, HSJA is deterministic and composed of 3 different stages where the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means to approximate the gradient at the decision boundary.
-In training and evaluation, the adversarial game is played as follows: the adversary starts by submitting a query, then the defender responds either truthfully (the true model prediction) or by misdirecting (the second highest probability class).
+Contrastively, HSJA is deterministic and composed of 3 different stages where the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means of approximating the gradient at the decision boundary.
+In training and evaluation, the adversarial game is played as follows: the adversary starts by submitting a query, then the defender responds either truthfully (the actual model prediction) or by misdirecting (the second highest probability class).
 Then the environment decides with chance $p$ if the adversary moves next, otherwise a benign query is drawn.
-In either case, it is the defender’s turn; in testing they are also oblivious to the nature of the query and know only the content.
+In either case, it is the defender’s turn; during testing they are also oblivious to the nature of the query and know only the content.
 All experiments are performed with $p=0.5$; we also evaluate our trained defense when no attack is present ($p=0$) in Appendix C.
-The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two different datasets – CIFAR10 and MNIST – and over two models with the same architecture but with different training regimes: one with adversarial training and one without.
-The transition from single-agent to multi-agent RL hides challenges: we approach the AMG as a belief-state MDP (the requirement of knowing the exact opponent policies is relaxed) and use PPO [42] agents to discover optimal policies that will also constitute best responses for the full AMG [47].
-However, learning independently of other agency breaks the theoretical guarantees of convergence [49], eg.
-in scenarios 7 & 8 where both agents learn simultaneously.
-The complete list of scenarios is:
 TABLE II: ASR and mean $L_{2}$ perturbation for 1K, 2K, and 5K queries for CIFAR-10, against normally and adversarially trained models.
@@
 reports the accuracy on benign queries of the base model plus any defenses present; in the first two scenarios (no active defense) the baseline clean accuracy is reported.
-Yellow scenarios denote the baseline attack performance, green denote active and/or adaptive defenses, and red denote adaptive attacks.
+Yellow scenarios denote the baseline attack performance, while green and red denote defensive and offensive scenarios respectively.
 The asterisk denotes where input transformations were used for evasion.
@@
 *8: AA-AD 19.86 19.86 19.86 0% 15.71 15.35 14.30 1% 87.67 87.40
+The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two datasets – CIFAR10 and MNIST – and over two models with the same architecture but different training regimes: with and without adversarial training.
+As the transition from single to multi-agent RL introduces non-stationarity, we approach the AMG as a belief-state MDP (relaxing the requirement of knowing the exact opponent policies), and use PPO [43] agents to learn optimal policies that will also constitute best responses for the full game [48].
+Note that learning independently of other agency breaks the theoretical guarantees of convergence [50], eg.
+in scenarios 7 & 8 where both agents learn simultaneously.
+Coloring denotes the learning/evaluated agent in each scenario, with their complete list being as follows:
   1. 0.
-    VA-ND – Vanilla Attack / No Defense: Baseline performance of attacks (BAGS & HSJA) out-of-the-box, without any active defense.
+    VA-NDVanilla Attack / No Defense: Baseline performance of attacks (BAGS & HSJA) out-of-the-box, without any active defense.
   2. 1.
-    AA-ND – Adaptive Attack / No Defense: How more optimal is the adaptive version of an attack compared to the baseline.
+    AA-NDAdaptive Attack / No Defense: How much more optimal is the adaptive version compared to the baseline attack.
   3. 2.
@@
   4. 3.
-    AA-VD – Adaptive Attack / Vanilla Defense: Similar to scenario (2), but now the attack is adaptive.
+    AA-VDAdaptive Attack / Vanilla Defense: Similar to scenario (2), but now the attack is adaptive.
   5. 4.
@@
   6. 5.
-    AA-TD – Adaptive Attack / Trained Defense: After the adaptive defense is optimized, its policy is fixed and an adaptive attack is trained against it.
+    AA-TDAdaptive Attack / Trained Defense: After the adaptive defense is optimized, its policy is fixed and an adaptive attack is trained against it.
   7. 6.
@@
   8. 7.
-    AA-AD – Adaptive Attack / Adaptive Defense: The first scenario where both agents learn simultaneously, making the environment non-stationary.
+    AA-ADAdaptive Attack / Adaptive Defense: The first scenario where both agents learn simultaneously, making the environment non-stationary.
 In practice, the convergence will vary and depend on the chosen hyperparameters and rewards.
@@
   9. 8.
-    AA-AD – Adaptive Attack / Adaptive Defense: The exact setup as scenario 7, but the best-case for the defense is reported instead.
-For each successive scenario we evaluate with the most successful policy found, as this is best practice in Markov Games: the worst case opponent policy is fixed and then a counter to it is learned [31, 38].
-Fixing other policies when computing a best response is representative of learning in environments that can contain multiple agents, with the additional benefit of converting the problem to single-agent that, as detailed in Section III-C, one can solve with RL.
+    AA-ADAdaptive Attack / Adaptive Defense: The exact setup as scenario 7, but the best-case for the defense is reported instead.
+In each successive scenario, we evaluate using the most successful past policy, following standard practice in Markov Games: the worst-case opponent policy is fixed, and a best response to it is learned [32, 39].
+Fixing other policies when computing a best response stabilizes learning in multi-agent environments, as it simplifies the problem to a single-agent setting – one that, as discussed in Section III-C, can be solved with standard RL.
 Comparison to SotA.
-In Scenarios 0-8 we evaluate all possible combinations between our attack and our defense.
+In Scenarios 0-8 we evaluate all possible combinations between (adaptive) attack and defenses.
 As a baseline to compare to, we additionally evaluate our approach to the state-of-the-art stateful defenses and adaptive attacks, that is Blacklight [19] and OARS [18] respectively.
@@
 As our environments do not return a rejection signal and to make a fair comparison, for OARS rejection coincides with a non-adversarial decision.
-We define 5 further scenarios:
+We thus define 5 further scenarios:
   1. 9.
-    VA-BD – Vanilla Attack / Blacklight Defense: Baseline performance of the attacks against Blacklight.
+    VA-BDVanilla Attack / Blacklight Defense: Baseline performance of the attacks against Blacklight.
   2. 10.
-    OA-BD – OARS Attack / Blacklight Defense: OARS against Blacklight.
+    OA-BDOARS Attack / Blacklight Defense: OARS against Blacklight.
   3. 11.
-    AA-BD – Adaptive Attack / Blacklight Defense: Our adaptive attack against Blacklight.
+    AA-BDAdaptive Attack / Blacklight Defense: Our adaptive attack against Blacklight.
   4. 12.
@@
     OA-AD – OARS Attack / Adaptive Defense: Our adaptive defense retrained against OARS.
-Our experiments were run on multiple machines, however to give an idea for the time complexity of our defense, on an Intel i7-7700 CPU one forward pass in CIFAR – that is one response to one query – takes $8\pm 1.4$ ms for 700 MFLOPs.
+Our experiments were run on multiple machines, yet to give an idea for the time complexity of our defense, on an Intel i7-7700 CPU one forward pass in CIFAR – that is one response to one query – takes $8\pm 1.4$ ms for $\sim$700 MFLOPs.
 V-B Results
 For consistency and comparability between evaluations, all results are from the same 100 test episodes.
-The gap value denotes the $L_{2}$ perturbation that initially separates the starting and best adversarial example, averaged over the 100 episodes.
+The gap value denotes the $L_{2}$ perturbation that initially separates the starting and the original samples, averaged over the 100 episodes.
 By testing the trained agents on budgets higher than 5K we discovered that the trend in reducing $L_{2}$ holds; to make the agent training tractable and the evaluation wider however, we limit the maximum query budget per adversarial episode to 5K.
-Tables II & III report the results for CIFAR10, while IV reports MNIST.
-Overall, the empirical results help us extract and highlight several important insights, practical observations, and general implications for the broader AML field:
+Table II & Table III report the results for CIFAR10, while  Table IV reports MNIST.
+The closer examination of the empirical results help us extract and highlight several important insights, practical observations, and general implications for the broader AML field:
 TABLE III: ASR and mean $L_{2}$ perturbation for CIFAR-10, comparing our Adaptive Attack (AA) and Adaptive Defense (AD) to Blacklight (BD) and OARS (OA).
@@
 13: OA-AD 20.01 20.01 20.01 0% 11.31 11.12 10.97 7% 87.66 87.74
+  • The initial performance of an attack can be misleading: out-of-the-box HSJA appears to be the better attack, but it is often outperformed by adaptive BAGS, especially in CIFAR and against active defenses.
+  • The performance of both attacks deteriorates considerably against active defenses (VD), however the defenses reach their full potential only when also adaptive (AD).
+  • Our adaptive defense (AD) outperforms both Blacklight (BD) and non-adaptive (VD) defenses, also when transferred (cf.
+Sc.12).
+Compared to Blacklight, it reduces ASR by $\sim$90(in HSJA) when trained against OARS, while it offers similar protection when transferred from another attack.
+  • Overall, against the strongest attacks and in the worst case (Sc.7) for it, our defense contains the ASR in the range of $1-36\%$.
+  • Our adaptive attack (AA) outperforms OARS (Sc.10 & 11) and vanilla attacks (VA) by a wide margin, without access to rejection sampling, and irrespective of the defense it faces; the one exception is our adaptive defense (AD), against which it has very limited success.
+  • The advantage of adaptive attacks is more pronounced against active defenses, where they significantly outperform non-adaptive versions, cf. Sc.0→1, Sc.2→3.
   • When comparing the upper and lower halves of each table, we can observe that adversarial training adds a limited amount of robustness; otherwise, the practical effect of adversarial training is a tax on the attacker, forcing them to expend more queries for the same perturbation or having higher perturbation for the same query budget.
-  • Our adaptive defense (AD) outperforms both Blacklight (BD) and non-adaptive stateful (VD), also when transferred (Scen.
-12).
-In HSJA, it reduces ASR by $\sim$90when trained against OA specifically, while it offers similar protection to BD when transferred from another attack.
-  • Even against the strongest attacker and for the worst case for the defense (Scen. 7), our AD achieves an ASR $\leq 23\%$.
-  • Our adaptive attack (AA) outperforms OARS (Scen.
-11) and vanilla attacks (VA) by a wide margin, without access to rejection sampling.
-and irrespective of the defense it faces; the one exception is our adaptive defense (AD), against which it has very limited success.
-  • Evasive transformations interfere with the attack operation, exemplified by the difference between BAGS and HSJA in Scen.
-5; for the attack to reach its full potential, they have to be adaptively controlled together.
-  • The initial performance of an attack can be misleading: out-of-the-box HSJA appears to be the better attack, but it is often outperformed by adaptive BAGS, especially in CIFAR and against active defenses.
-  • In Scen.
-1-8, where agents train against the best opponent policy as previously discovered, ASR oscillates as following a fixed policy enables the learning of an optimal counter to it; and eventually plateaus, indicating an equilibrium.
-  • The advantage of AA is much more pronounced against active defenses, where they significantly outperform non-adaptive versions.
-  • The performance of both attacks deteriorates considerably against active defenses, however the defenses reach their full potential only when also adaptive.
-  • The first time active defenses effectively resist adaptive attacks is in Scen. 5 of CIFAR; we employed evasive transformations from then onward.
-  • Different attack fundamentals respond differently to defenses; the gradient estimation stage of HSJA is naturally disadvantaged against similarity detection, while the jump and binary stages are advantaged.
-  • When devising the adaptive defense for HSJA, it proved nearly impossible to engineer a state the agent can learn on by leveraging our knowledge of the attack and its geometric functioning.
-What did prove effective, however, was pure computation [^2] This reminds us of Sutton’s Bitter Lesson [50], the observation that progress in AI is often driven by gains in computation rather than problem-specific expert knowledge.: we learn an embedding with Contrastive Learning [51] from raw input to use a state, which transfers exceedingly well to other attacks like BAGS.
+  • Evasive transformations interfere with the attack policy, as illustrated by the difference between BAGS and HSJA in Sc.5.
+For an attack to reach its full potential, these two should be adaptively controlled together.
+  • In Sc.1 to 8, agents train against the best opponent policy as previously discovered, and ASR oscillates since following a fixed policy enables the learning of an optimal counter to it.
+Over successive adaptations, ASR eventually plateaus, indicating an equilibrium for each specific dataset and attack combination.
+This is illustrated in Figure 4.
+  • The first time active defenses resisted adaptive attacks was in Sc.5 of CIFAR; we employed evasive transformations from then onward.
+  • Different attack fundamentals respond differently to active defenses; the gradient estimation stage of HSJA has a disadvantage against similarity detection, while the jump and binary stages have an advantage.
+  • For HSJA, engineering a state the adaptive defense could learn on, merely by leveraging our knowledge of the attack and its geometric functioning, proved impossible.
+What did prove effective however, was pure computation [^2] This is reminiscent of Sutton’s Bitter Lesson [51], the observation that progress in AI is often driven by gains in computation rather than problem-specific expert knowledge.: we used Contrastive Learning [52] to learn an embedding from raw queries, then used as state that transfers exceedingly well to other attacks like BAGS.
 TABLE IV: ASR and mean $L_{2}$ perturbation for 1K, 2K, and 5K queries and accuracy on clean data for MNIST. The evaluation scenarios are identical to Tables II and III.
@@
 Another simplifying assumption we make is that only one attack can take place at a time; however, the queuing technique we use for incoming queries is readily extensible to handle concurrent attacks.
-While we demonstrate that our active defense does transfer between attacks, another possibility to explore is training the defense on multiple kinds of attacks.
+While we demonstrate that our active defense does transfer between attacks, another possibility to explore is training the defense on queries from different kinds of attacks.
 Finally, in our evaluation we focus on a wide range of adaptive and non-adaptive scenarios where agents learn and adapt interactively, thus limiting the number of datasets we experiment with; we back our empirical study however with an extensive theoretical analysis that supports the generality of our findings independent of context.
@@
 binaries) and instead can function directly in the problem space [37].
-Another compelling and formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing instruments to bypass defenses and thus imparting controllability to adversarial tasks.
+Another compelling and formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing tools to bypass defenses and thus imparting controllability to adversarial tasks.
 Finally, in our work we considered opponent agency as part of the environment; other domains, like malware detection, might benefit from explicit opponent modeling.
 VII Conclusion
-With adaptive, decision-based attacks becoming more pervasive in multiple domains, every AI-based decision-making process that exposes a queryable interface is inherently vulnerable.
+With adaptive, decision-based attacks becoming more pervasive in multiple domains, every AI-based system that exposes a queryable interface is inherently vulnerable.
 To aggravate matters, this vulnerability cannot be mitigated by employing model hardening approaches like adversarial training alone.
@@
 We perform a theoretical and empirical investigation of decision-based attacks and stateful defenses under a unified framework we name “Adversarial Markov Games” (AMG).
-In self-adaptive, we introduce a novel twofold definition of adaptive: both devising new methods of outmaneuvering opponents and adapting one’s operating policy with respect to other agency in the environment.
-Furthermore, our theoretical analysis highlights that any combination of adversarial goals, be it performance, stealthiness [52], or disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any modality.
-As new attacks and defenses appear and get broken regularly, our evaluation methodology is generally applicable within the outlined scope by turning any approaches in the existing arms-race self-adaptive and thus ensuring accurate and robust assessment of their performance.
+In self-adaptive, we introduce a novel twofold definition of adaptive: both inventing new techniques to outmaneuver opponents and adapting one’s operating policy with respect to other agency in the environment.
+Furthermore, through our theoretical analysis we demonstrate how any combination of adversarial goals, be it performance, stealthiness [53], or disruption, can be optimized in a gradient based manner, even in the complete black-box case and in any domain.
+As new attacks and defenses constantly emerge and are surpassed, our proposed methodology is generally applicable as it turns any such approaches in the current arms-race self-adaptive, thus ensuring accurate and robust assessment of their performance.
 The AMG framework we introduce helps us reason on and properly assess the vulnerabilities of AI-based systems, disentangling the inherently complex and non-stationary task of learning in the presence of competing agency.
 By modeling the latter as part of the environment, we can simplify this task by computing a best response to the observed behavior.
-This is a significant implication for the security of AI-based systems: as long as proper threat modeling is carried out, one can readily employ RL algorithms in order to devise optimal defenses, but only after they devised optimal attacks too.
+This has a significant consequence for the security of AI-based systems independent of modality or application: as long as proper threat modeling is carried out, one can readily employ RL agents in order to devise optimal defenses, but only after they devised optimal attacks too.
 Appendix A Proofs
@@
 Proof.
-The proof is constructed in two parts: first that an alternative decision function $D$ is required, and subsequently that a stateful representation of a query $x_{t}$ is also required.
-Let us denote by $x_{c},x_{g},x_{t}$ the original (unperturbed), the starting, and the current sample at step $t$ respectively.
-Given a target class $c_{0}\in m$ we can define a function:
+We proceed in two parts, corresponding to conditions (a) and (b).
+[Part 1] Let us denote by $x_{c},x_{g},x_{t}$ the original (unperturbed), the starting, and the current sample at step $t$ respectively. Given a target class $c_{0}\in m$ we define a function:
 $S_{x_{c}}(x_{t})=F_{c_{0}}(x_{t})-\underset{c\neq c_{0}}{\max}(F_{c}(x_{t}))$ (8)
-HSJA operates in 3 stages that alternate until convergence or when the maximum query budget is reached: 1) A binary search between $x_{g}$ and $x_{c}$ that places $x_{t}$ on the decision boundary between the classes.
-2) A gradient estimation stage that approximates $\nabla S(x_{t})$.
-3) A “jump” step along the direction of the gradient $\nabla S(x_{t})$.
-For practicality we reiterate Eq.
-9 of [7] which denotes that the gradient direction is approximated via the following Monte Carlo estimate:
+HSJA operates in 3 stages that alternate until convergence: (1) binary search between $x_{g}$ and $x_{c}$ that places $x_{t}$ on the decision boundary, (2) gradient estimation approximating $\nabla S(x_{t})$, (3) a step along the direction of the gradient $\nabla S(x_{t})$.
+We repeat Eq.
+9 of [7], denoting the gradient direction as the Monte Carlo estimate:
 $\widetilde{\nabla S_{x_{c}}}(x_{t},\delta)=\frac{1}{B}\sum_{b=1}^{B}\phi_{x_{c}}(x_{t}+\delta u_{b})u_{b}$ (9)
 where $\{u_{b}\}_{b=1}^{B}$ are i.i.d.
-draws from the uniform distribution over the $d$-dimensional sphere, $\delta$ is a small positive parameter, and $\phi_{x_{c}}$ is the Boolean-valued function that all steps of HSJA rely on:
-$\phi_{x_{c}}(x_{t})=\text{sign}(S_{x_{c}}(x_{t}))=\begin{cases}1&\text{if}\quad S(x_{t})>0,\\ -1&\text{if}\quad S(x_{t})\leq 0.\end{cases}$ (10)
-Given an initial sample $x_{c}$, in search of adversarial examples HSJA applies the following update function that sequentially updates $x_{c}$:
+draws from the uniform distribution over the $d$-dimensional sphere, $\delta$ is a small positive parameter, and $\phi_{x_{c}}$ is the Boolean-valued function that all stages rely on:
+$\phi_{x_{c}}(x_{t})=\text{sign}(S_{x_{c}}(x_{t}))=\begin{cases}+1&\text{if}\quad S(x_{t})>0,\\ -1&\text{if}\quad S(x_{t})\leq 0.\end{cases}$ (10)
+Given $x_{c}$, in search of adversarial examples HSJA iteratively applies the following update function:
 $x_{t+1}=a_{t}x_{c}+(1-a_{t})\Biggl{\{}x_{t}+\xi_{t}\frac{\nabla S_{x_{c}}(x_{t})}{\|\nabla S_{x_{c}}(x_{t})\|_{2}}\Biggl{\}}$ (11)
@@
 Let us define the function $\mathcal{I}$ of two variables:
-$\mathcal{I}(a,b):=\begin{cases}1&\text{if}\quad a=b,\\ -1&\text{if}\quad a\neq b.\end{cases}$ (13)
-We can rewrite Eq. 10 through and Eq. 13 and inequalities 12 as follows:
-$\phi(x_{t})=\mathcal{I}(C(x_{t}),c_{0})$ (14)
-Provided that the gradient estimation happens at the decision boundary where $S(x_{t})=0$, Theorem 2 of [7] guarantees that the estimation of stage 2 is an asymptotically unbiased direction of the true gradient:
+$\mathcal{I}(a,b):=\begin{cases}+1&\text{if}\quad a=b,\\ -1&\text{if}\quad a\neq b.\end{cases}$ (13)
+From 12 and 13 we can rewrite Eq. 10 as follows:
+$\phi_{x_{c}}(x_{t})=\mathcal{I}(C(x_{t}),c_{0})$ (14)
+Provided that the gradient estimation happens at the decision boundary where $S(x_{t})=0$, Theorem 2 of [7] guarantees that the gradient estimation is an asymptotically unbiased direction of the true gradient:
 $\widetilde{\nabla S_{x_{c}}}(x_{t},\delta)\approx\nabla S_{x_{c}}(x_{t}),\delta\rightarrow 0$ (15)
-For $b_{t}=1-a_{t}$ and by plugging Eqs 14 & 15 in Eq. 9, and the result in 11, we get:
+For $b_{t}=1-a_{t}$ and by plugging 14 & 15 in Eq. 9, and the result in 11, we get:
 $x_{t+1}=a_{t}x_{c}+b_{t}\Biggl{\{}x_{t}+\xi_{t}\frac{\nabla S_{x_{c}}(x_{t})}{\|\nabla S_{x_{c}}(x_{t})\|_{2}}\Biggl{\}}$
@@
 (16), the iterates $x_{t}$ are guaranteed by Theorem 1 of HSJA [7] to converge to a stationary point $x_{b}$ of Eq.
-(5), that is $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$, for $\epsilon$ a standard imperceptibility threshold and $N$ adversarial episodes.
-We thus reach a contradiction, and as in Eq.
-(16) the only model-dependent term is $C(\cdot)$, an alternative classifier $C^{\prime}$ is required.
+(5), that is $\mathbb{E}[\sum_{i=1}^{N}d(x^{i}_{b},x^{i}_{c})]<\epsilon$, for $\epsilon$ a standard imperceptibility threshold and $N$ adversarial episodes, which contradicts the requirement $\mathbb{E}[\sum_{i}d(x_{b}^{i},x_{c}^{i})]\geq\epsilon$.
+Since the only model-dependent term in Eq.
+(16) is the classifier $C(\cdot)$, the contradiction can be avoided only with an alternative classifier $C^{\prime}$.
 With $C(x)=D(F_{c}(x))$, and the discriminant function $F_{c}$ unable to change without retraining, it follows that $D^{\prime}\neq\operatorname*{arg\,max}$, so for adv.
@@
 $\therefore C(x_{t})=\hat{c},\>\hat{c}=\{c_{0},m\setminus c_{0},\varnothing\}$
-where the empty decision $\varnothing$ denotes rejection and $\{m\setminus c_{0}\}$ denotes misdirection, i.e. intentional misclassification.
-Decision-based attacks initiate from examples $x_{t}$ that belong to the target class $c_{0}$; while for $t=0$, $x_{t}$ is not adversarial yet, it still is part of an ongoing adversarial attack.
-To deter the attack, a perfect defender would have to misclassify/reject this example; yet if an identical but benign example $x_{n}$ was submitted, classifier $C$ should preserve its capacity to classify it correctly.
-For the identical examples $x_{n}=x_{t}$ we want $C(x_{n})\neq C(x_{t})$, something possible only with context.
-This context $\tau$, based on the sequence of incoming queries $x\in\{x_{t},t=0,1,2,...,N\}\cup x_{n}$, can be provided as an additional argument to the decision function $D(\tau,F_{c}(x))$ s.t.
-while $x_{n}=x_{t}\implies D(\tau,F_{c}(x_{n}))\neq D(\tau,F_{c}(x_{t}))$.
+where $\varnothing$ denotes rejection and $\{m\setminus c_{0}\}$ denotes misdirection, i.e. intentional misclassification.
+[Part 2] Let us assume that at timestep $t$, $x_{t}$ is not yet adversarial, it is however still part of an ongoing adversarial attack.
+To deter the attack, a perfect defense would have to misclassify/reject this example; yet if an identical but benign example $x_{n}$ was submitted, classifier $C$ should preserve its capacity to classify it correctly.
+Since any memoryless classifier must assign the same label whenever $x_{n}=x_{t}$, we require a richer decision rule $C^{\prime}(x,\tau)=D\bigl{(}\tau,\;F_{c}(x)\bigr{)}$ that takes as auxiliary input context $\tau$, e.g.
+the history of queries $\{x_{0},\dots,x_{t}\}\cup x_{n}$.
+By construction, even if $x_{n}=x_{t}$, differing contexts $\tau_{n}\neq\tau_{t}$ can force $C^{\prime}(x_{n},\tau_{n})\;\neq\;C^{\prime}(x_{t},\tau_{t})$, thereby separating benign from adversarial queries.
 ∎
@@
 This attack is in effect a gradual interpolation from $x_{g}$ towards $x_{c}$, by first taking orthogonal steps $x_{s}$ on the hypersphere around $x_{c}$ and then source steps towards $x_{c}$ in order to minimize $d(x_{c}-x_{b})$, where $x_{b}$ is the best adversarial example found so far.
-The source step parameter $\epsilon=(1.3-\min(\lambda_{n},1))\cdot c$ — with $\lambda_{n}$ the ratio of the $n$ last queries $x_{t}$ that are adversarial and $c$ a positive constant — controls the projection towards $x_{c}$:
+The source step parameter $\epsilon=(1.3-\min(\lambda_{n},1))\cdot c$ – with $\lambda_{n}$ the ratio of the $n$ last queries $x_{t}$ that are adversarial and $c$ a positive constant – controls the projection towards $x_{c}$:
 $x_{t}=x_{s}+\epsilon\cdot(x_{c}-x_{s})$ (18)
@@
 [States].
-By merit of operating in black-box environments, the optimization process for both offensive and defensive agents can be described as a POMDP.
-To handle this partial observability, the states we engineer incorporate past information.
+To handle the partial observability, we engineer states that incorporate past information.
 For BAGS, the adversary uses an 8-dimensional state representation with the following information normalized in the range $[0,1]$ : current amount of queries $i$, average queries that are adversarial $a$, the initial gap $g$, the current gap $d$, the location $l=\frac{d}{g}$, the slope $s=m-l$ where $m$ is a moving average of the location, the frequency of improvement $f$, and $r$ which is a moving average of the perturbation reduction $n$, In HSJA the state representation is slightly different: $r=\frac{n}{g}$, and $f=\frac{1}{j}$ with $j$ number of jump steps in last iteration.
@@
 Policies are parameterized by a two fully-connected layer NN; the hyperparameter search space is shown in Table VII.
+Figure 4: Progression of ASR over successive adaptations. Red and green values in ASR denote offensive and defensive scenarios respectively.
 Blacklight & OARS.
@@
 Additionally, as our defense is not rejection based, we replace the rejection decision with a non-adversarial one.
-TABLE VII: Hyperparameter space for the PPO agent.
+TABLE VII: Hyperparameter ranges during PPO training.
 Hyperparameter BAGS HSJA
@@
 total steps 1e5 – 4e5 2e4 – 2e5
+total queries 25K – 100K 5K – 50K
 batch size 32 – 128 32 – 64
-buffer 2048 – 2048 64 – 1024
+buffer size 2048 – 2048 64 – 1024
 epochs 20 – 20 20 – 20